The CLEO III Detector

Design and Physics Goals

California Institute of Technology University of California at San Diego

University of California at Santa Barbara

University of Colorado Cornell University University of Florida Harvard University

University of Illinois at Urbana-Champaign

Ithaca College University of Kansas University of Minnesota

State University of New York at Albany Ohio State University University of Oklahoma

Purdue University University of Rochester

Southern Methodist University Syracuse University Vanderbilt University

Virginia Polytechnic Institute and State University

February 18, 1994

Contents

1 Introduction 1 1.1 Lessons from the CLEO II detector . . . . . . . . . . . . . . . . . . . . . . . 2 1.2 Design philosophy for CLEO III . . . . . . . . . . . . . . . . . . . . . . . . . 5

1.2.1 Tracking . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 1.2.2 Particle identification . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

1.3 Physics overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 1.3.1 B physics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 1.3.2 Charm physics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 1.3.3 Tau physics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

2 Physics at a Symmetric 10^{33} e^{+}e^{−} Collider 11

2.1 B decay studies . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 2.1.1 Goals of B decay studies . . . . . . . . . . . . . . . . . . . . . . . . . 11 2.1.2 V_{cb} and form-factors in semileptonic decays . . . . . . . . . . . . . . . 13 2.1.3 Rare B decays . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 2.1.4 CP violation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22

2.2 Charm decay physics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 2.2.1 Semileptonic decays . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 2.2.2 Hadron spectroscopy . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 2.2.3 Rare charm decays . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 2.2.4 Measurement of f_{D_{s}} . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 2.2.5 Charm mixing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33

2.3 Tau decay physics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 2.3.1 Tau neutrino mass . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 2.3.2 Decay modes and branching ratios . . . . . . . . . . . . . . . . . . . 36 2.3.3 Structure of the τ − ν_{τ} − W vertex . . . . . . . . . . . . . . . . . . . 41 2.3.4 Tau lifetime . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 2.4 Two photon physics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 2.5 Υ physics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 2.6 Conclusions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46

3 Simulation of CLEO II and CLEO III 47 3.1 Parametric model studies . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 3.2 Material effects on the CsI calorimeter . . . . . . . . . . . . . . . . . . . . . 51 3.3 Simulation results . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51

3.3.1 Single track simulations . . . . . . . . . . . . . . . . . . . . . . . . . 52 3.3.2 D^{∗} − D mass difference . . . . . . . . . . . . . . . . . . . . . . . . . . 54 3.3.3 B decay simulations . . . . . . . . . . . . . . . . . . . . . . . . . . . 54

3.4 Summary . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56

4 Detector Design 60 4.1 Beampipe . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 4.2 Silicon . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62

4.2.1 General detector parameters . . . . . . . . . . . . . . . . . . . . . . . 62 4.2.2 Resolution . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 4.2.3 Capacitance . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 4.2.4 Readout . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 4.2.5 Signal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 4.2.6 Noise . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 4.2.7 Design options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 4.2.8 R&D . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77

4.3 Drift chamber . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 4.3.1 Mechanical overview . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 4.3.2 Drift chamber endplate . . . . . . . . . . . . . . . . . . . . . . . . . . 81 4.3.3 Gas, wire choice, and cell design . . . . . . . . . . . . . . . . . . . . . 82 4.3.4 Wire holding . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 4.3.5 Cathodes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 4.3.6 R&D . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89

4.4 Particle identification at high luminosities . . . . . . . . . . . . . . . . . . . 90 4.4.1 Ring imaging Cherenkov counters . . . . . . . . . . . . . . . . . . . . 90 4.4.2 Aerogel . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 4.4.3 Cherenkov correlated timing . . . . . . . . . . . . . . . . . . . . . . . 94 4.4.4 High pressure Cherenkov tubes . . . . . . . . . . . . . . . . . . . . . 97 4.4.5 R&D . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99

4.5 CsI endcap calorimeter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 4.5.1 Mechanical modifications . . . . . . . . . . . . . . . . . . . . . . . . . 101 4.5.2 R&D . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102

4.6 IR mounting and assembly sequence . . . . . . . . . . . . . . . . . . . . . . . 103 4.6.1 Mounting system . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 4.6.2 Assembly sequence . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105

4.7 Detector backgrounds . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 5 Front-end electronics, triggering, and data acquisition 111

5.1 CLEO II system . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 5.1.1 Trigger . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 5.1.2 Gating and event flow control . . . . . . . . . . . . . . . . . . . . . . 112 5.1.3 Front-end electronics . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 5.1.4 Data acquisition . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 5.1.5 System performance . . . . . . . . . . . . . . . . . . . . . . . . . . . 113

5.2 CLEO III . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 5.2.1 Trigger . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115 5.2.2 Gating and event flow control . . . . . . . . . . . . . . . . . . . . . . 120 5.2.3 Front-end electronics . . . . . . . . . . . . . . . . . . . . . . . . . . . 122 5.2.4 Data acquisition . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 126 5.2.5 Slow controls . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 128

5.2.6 R&D . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 129

6 IR design 133 6.1 Overview of CESR luminosity upgrade . . . . . . . . . . . . . . . . . . . . . 133 6.2 Accelerator components . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 134

6.2.1 IR magnets . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 134 6.2.2 Parasitic beam crossings . . . . . . . . . . . . . . . . . . . . . . . . . 134 6.2.3 Optics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 134 6.2.4 Aperture . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 134 6.2.5 Mechanical requirements . . . . . . . . . . . . . . . . . . . . . . . . . 135 6.2.6 Magnet layout options . . . . . . . . . . . . . . . . . . . . . . . . . . 135 6.2.7 Mounting and support . . . . . . . . . . . . . . . . . . . . . . . . . . 138 6.2.8 Vacuum system . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 138

7 Overall summary 139

1 Introduction

The CLEO collaboration has made major contributions to many aspects of b, c, τ, and γγ physics with precision unmatched anywhere in the world. This has been made possible by the CLEO II detector, with its combination of excellent tracking and CsI electromagnetic calorimetry, and the large integrated luminosity provided by the CESR accelerator. CESR is embarking on an upgrade program over the next five years after which integrated luminosities up to 15 fb^{−1}/year—10 fb^{−1} on the Υ(4S) resonance and 5 fb^{−1} below the Υ(4S)—will be provided. CLEO data sets can be increased by up to an order of magnitude which will allow further, significant strides forward in our studies of heavy quarks and third generation leptons.

A central feature of the CESR upgrade, called Phase 3, is larger focussing quadrupoles that intrude into space now occupied by the CLEO II tracking detectors (see Fig. 1), necessitating replacement of these aging devices. It should be noted that the VD wire chamber is now 10 years old and the main drift chamber is 6 years old. These elements need replacing regardless of the accelerator upgrade and in addition the drift chamber endplates and mounting structure compromise the performance of the CsI calorimeter. The other sub-detectors—CsI calorimeter, magnet coil, and muon system—are compatible with the machine upgrade and our future physics goals with only small modifications, although higher data rates will require more sophisticated readout electronics. The new tracking system design compatible with the upgraded IR design will also provide the space for a high-quality particle identification system (see Fig. 2) that will allow 4σ π/K separation up to 2.8 GeV/c. This upgrade program will produce the CLEO III detector and will benefit all areas of CLEO physics and will be particularly important for the study of B decays, especially rare decays as a window to observe physics beyond the Standard Model.

The most important physics goals for CLEO remain those for B physics. While additional integrated luminosity would itself add to our knowledge, greater detection and reconstruction efficiency from pion and kaon identification up to momenta of 2.8 GeV/c will bring faster rewards. This is analogous to the introduction of excellent photon detection in the CLEO II detector which increased the number of reconstructed B’s per unit luminosity. Other important topics which may become accessible include the measurement of Cabibbo-suppressed “penguin” decays such as ργ. A measurement of the ratio of ργ to K^{∗}γ, the latter recently measured by CLEO II [1], is an excellent way to determine V_{td}/V_{ts}. High momentum particle identification coupled with the large CLEO III data sample will make it possible to attempt the observation of CP violation in B decays by, for example, measuring the rate asymmetries in the decays K^{−}π^{+} and K^{+}π^{−} [2]. In addition, direct measurement of one of the CP violating angles can be attempted by measuring the branching ratios for B decays to D^{0} where the D^{0} decays into a CP eigenstate.

CLEO III will also be able to address many new topics in the areas of charm and tau physics. For example, with high momentum particle identification, the Cabibbo-suppressed semileptonic charm decays such as ρ^{−}ℓ^{+}ν can be measured even though there is a much larger rate from K^{∗−}ℓ^{+}ν. For both charm physics and the study of fragmentation, the momentum range for which such particle separation is useful extends up to near the kinematic limit of approximately 5 GeV/c. Also, the tau neutrino mass can best be limited from studies of the decay τ^{−}→(K^{∗}K)^{−}ν̄ due to the large minimum hadronic mass in this channel (the present acceptance is <1%). Again, particle identification is crucial.

1.1 Lessons from the CLEO II detector

The CLEO II detector was designed with the goal of reconstructing thousands of B mesons. The detector philosophy was to create a device with excellent charged particle momentum resolution and excellent electromagnetic (electron and photon) energy resolution. This has been achieved. As implemented, the CLEO II detector has attained a charged particle momentum resolution of

(σ_{p}/p)^{2} = (0.0015p_{t})^{2} + (0.0056)^{2} for p_{t} > 0.75 GeV/c  (1)

and an electromagnetic (e and γ) energy resolution of

σ_{e}/E = 0.0035/E^{0.75} + 0.019 − 0.001E Barrel  (2)
σ_{e}/E = 0.0026/E + 0.025  Endcap  (3)

where momentum is in GeV/c and energy is in GeV. In addition, particle identification is provided by a measurement of specific ionization (dE/dx) in the central tracking chambers— 2σ π/K separation level for momenta below 750 MeV/c and 1.6σ at 2.3 GeV/c—and by a scintillator-based time-of flight system in front of the calorimeter—2σ below 1 GeV/c.

Although the CLEO II detector has performed admirably, there are weaknesses and design limitations which affect its performance at high luminosity:

1. The mechanical structure of the inner tracking system—Precision Tracking Layer (PTL) detector, Vertex Detector (VD), and the Drift Chamber (DR 2)— is incompatible with the CESR Phase 3 interaction region. The severe magnetic field requirements for a high luminosity interaction region imply a stepped permanent quadrupole (17 cm radius) along a 300 mrad line from the interaction region followed by a superconducting quadrupole (20 cm radius). The 300 mrad line is shown superimposed on the CLEO II detector in Fig. 1, indicating that the CLEO II tracking system is not compatible with the high luminosity collider.

2. The drift chamber mounting and stabilizing ring is a solid aluminum structure which shadows the last 6 rows of crystals of the barrel calorimeter (shaded in Fig. 1) and causes inferior energy and angular resolution and lost efficiency.

Figure 1: A quarter-section of the CLEO II detector showing the drift chambers (PTL,VD, and DR 2), the CsI calorimeters (BCC and ECC), the time-of-flight (BTF and ETF), and the superconducting solenoid (COIL). The muon counters are beyond the coil and the endcap. The angled line originating at the interaction point (IP) is 300 mrad from the beam line. At about 35 cm from the IP, the territory between this line and the beam line will be occupied by the CESR Phase 3 upgrade.

Figure 2: A half-section view of the CLEO III detector inside the superconducting solenoid and the CESR Phase 3 interaction region components. The 20 cm clear space around the drift chamber is reserved for the particle identification system—the detector technology is yet to be chosen.

3. The momentum range for π/K separation is limited. The present detector uses a combination of dE/dx and time-of-flight below 750 MeV/c, time-of-flight alone up to 1 GeV/c, and dE/dx alone above 2.3 GeV/c.

4. The endcap calorimeter, which represents 18% of the solid angle, has not been used in most analyses due to its low efficiency and poor resolution caused primarily by the large amount of material in front of it and its non-projective geometry.

5. The present electronics will not support the large data rates associated with 10^{33} cm^{−2}sec^{−1} operation.

1.2 Design philosophy for CLEO III

There are several constraints placed on the design of CLEO III. They are:

1. The requirements of the new CESR IR clearly determine that the inner detector must be replaced and that the maximum attainable acceptance is determined by the 330 mrad line (93% of 4π) in Fig. 2. This leaves adequate clearance between CESR and CLEO III for a shared space for cabling and plumbing.

2. The new particle identification system must fit inside the existing CsI calorimeter volume.

3. The new tracking system must allow space for the particle identification system without compromising the tracking performance or the total acceptance.

4. The barrel CsI calorimeter will remain intact but we may rework the endcap calorimeter if it can be demonstrated to greatly enhance the quality of the data.

1.2.1 Tracking

The performance of a tracking system may be quantified in terms of its resolution in momentum, angle, impact parameter, and vertexing. The present CLEO II tracking system has demonstrated that it meets all the requirements in these variables to study the physics we anticipate at high luminosity. We have therefore defined this as the minimum goal for the CLEO III tracking system. This is challenging since the tracking length in CLEO III will be shorter than in CLEO II. Consider the charged particle momentum resolution which is given by:

(\frac{dp}{p})^{2} = (\sqrt{\frac{720}{N + 5}} · \frac{σ}{0.3Bℓ^{2}} · p_{t})^{2} + (\frac{0.016}{0.3Bℓβ} · \sqrt{\frac{ℓ/sinθ}{X_{0}}} )^{2}  (4)

where B is the magnetic field strength (Tesla), N is the number of measurements along the track, ℓ is the projected track length onto the bending plane (m), X_{0} is the radiation length of the scattering medium (m), σ is the average space point resolution (m), β is v/c, and θ is the dip angle. The first term in the resolution formula is due to tracking variables, namely resolution, track length and number of measurements. The second term is the contribution due to multiple scattering. Thus, to maintain a fixed momentum resolution as ℓ is reduced requires that X_{0} must be increased correspondingly. Similar arguments hold for other parameters. Therefore, the general philosophy of the CLEO III design is to reduce the tracking length of the drift chamber to provide space for particle identification while maintaining the performance by simultaneously reducing the amount of material in that tracking length.

1.2.2 Particle identification

A major goal of the physics program is to cleanly measure rare decays to test the Standard Model and be sensitive to possible new physics. This requires excellent charged hadron identification with 4σ separation of pions and kaons up to 2.8 GeV/c. Even though a dE/dx capability will be retained in the readout of the new drift chamber, the momentum range over which we have good identification needs to be extended upward in order to access many of the most important physics issues. The new system must also be no thicker than 12% of a radiation length—the present time-of-flight scintillator thickness—in order not to compromise the performance of the CsI calorimeter.

The required performance of a particle identification system depends on the process being studied. The number of standard deviations of separation, n_{σ}, is defined as

n_{σ} = \frac{(M_{K} − M_{π})}{\sqrt{(\frac{1}{2}) (σ_{K}^{2} + σ_{π}^{2})}},  (5)

where M_{K}, M_{π}, σ_{K}, σ_{π} are the means and standard deviations of the K and π distributions respectively. Excellent work on all B decay physics can be accomplished if one has 4σ separation at 2.8 GeV/c, the maximum momentum reached in B decays. As an example, consider the problem of separating B̄^{0}→K^{−}π^{+} from π^{+}π^{−}. The separation needed depends on the relative populations in the two channels, and the signal purity desired. In Fig. 3, the plot of separation versus fraction of the signal which is π^{+}π^{−} is shown. If the π^{+}π^{−} fraction is only 10%, the rejection required is at the 3–4 σ level for reasonable (> 90%) purities (ignoring other sources of background). For the B → D^{0}K^{−} modes, the background just from D^{0}π^{−} is about 20 times the signal without particle identification and thus 4 σ rejection is essential. The separation at even higher momentum is also useful for other physics such as light quark fragmentation, charm decays, and τ decays. Although we have specifically mentioned π/K separation, another important consideration is the ability to separate protons from kaons.

Figure 3: The number of σ separation between kaons and pions required to separate K^{−}π^{+} from π^{+}π^{−} as a function of the fraction of π^{+}π^{−} and the required purity in in the K^{−}π^{+} sample.

1.3 Physics overview

In this section we briefly describe some of the physics reach of the CLEO III detector. Much more detail is provided in the next section.

1.3.1 B physics

Two topics at the frontier of B physics are b → u decays and penguin decays, both of which occur in the decays of B mesons to two light mesons. The decay B^{0}→π^{+} π^{−} is an example of a two-body b → u decay, and it is most likely the mode in which CP violation in such decays will eventually be seen. The decay B^{0}→K^{+} π^{−} is an example of a hadronic penguin decay which should be observed in the near future. For each of these modes, the other one is a serious background, depending on the relative branching ratios and the ability to separate pions from kaons. Measuring both decays will also give some estimate of the effect of penguins in the B^{0}→π^{+} π^{−} decay. A statistically significant signal has been observed in CLEO II data in the sum of these two modes and limits on the individual branching ratios are found to be less than 3×10^{−5}. At the present level of statistics the separation of the two channels is not possible. With the increased luminosity of the upgrade, these two modes will both be measurable. Unfortunately as the expected branching ratios decrease, the background suppression becomes more difficult. The existing analyses show that good momentum resolution is critical. It will be necessary to separate the two modes cleanly, especially if one turns out to be much larger than the other. This requires π/K identification at momenta of 2.8 GeV/c. Furthermore, an attempt to measure CP violation would involve telling K^{+}π^{−} from K^{−}π^{+}. In this case there is no help from kinematic constraints and the particle identification system must provide full discrimination.

One of the centerpieces of the physics program at the upgraded CESR will be measuring the CKM matrix elements as precisely as possible. These are important not only as fundamental parameters of the Standard Model, but also as the information needed to predict the CP asymmetries to be measured later. V_{ub} will be measured using the exclusive decay channels B → ρ(ω)ℓ^{−}ν̄_{ℓ} because they are the likely dominant modes beyond the b → c endpoint in the lepton energy spectrum, and because the form factors will be reasonably well understood. Although the recent CLEO analysis sets upper limits on the branching ratios for these decay modes, the size of the b → u signal in the inclusive spectrum implies that they will be observed very soon. The analysis requires a large background suppression, which relies on almost every aspect of the CLEO detector. In the present analysis, the dominant background is due to continuum leptons combined with other pions in the event, which happen to pass the strict kinematic cuts. Information from the silicon vertex detector will provide an extra handle to remove this background. Real leptons coming from charm decays will tend to have a three-dimensional impact parameter removed from the event vertex, while those from B decay will not. Light quark events which include fake leptons will be consistent with having only one vertex, while Υ(4S) events will have multiple vertices. Thus suppressing these backgrounds will give a good measure of performance for the measurement of vertices.

The familiar decay chain B̄^{0}→D^{∗+}π^{−}, D^{∗+}→π^{+}D^{0}, although not as interesting a physics mode in the future, serves as a useful measure of performance for many other B decays, such as B̄^{0}→D^{∗+}ℓ^{−}ν̄_{ℓ}. The comparison of efficiency and background suppression with the present detector tests many performance characteristics. In particular, the acceptance for the slow pion and the resolution in mass difference must be studied carefully. We already know from CLEO II how important, but difficult, it is to measure this pion’s parameters well. The D^{∗+} mass difference is one case in which the angle measurement dominates the resolution, so it tests the measurement of tan θ particularly well. It has already been shown that the event vertex constraint is needed to get full use of the silicon detector in measuring this angle.

We have studied in some detail the practicality of the suggestion of Gronau and Wyler[3] to measure CP violation with self-tagging modes. Although it is unclear if such an effect could be seen with the luminosity available with the CESR upgrade, the B decay modes involved are interesting in their own right, and provide a good test of the proposed detector. It also makes it possible to look for the CP violation, in case the relevant branching ratios turn out to be larger than expected. In particular, we have examined the decay mode B^{−}→D̄^{0}K^{−}. This is a b→u decay and color suppressed, so that the branching ratio is expected to be in the range of 10^{−6}. The decay B^{−}→D^{0}K^{−}, which is a Cabibbo-suppressed b→c decay, has a branching ratio in the range of 10^{−4} and represents a more accessible mode with the same experimental requirements. Even measuring this mode is well beyond the capability of CLEO II.

Studying these modes will give a good measure of the background suppression for a wide range of rare B decays. We will need high efficiency for detecting D^{0}’s with small background.

We will use many decay modes, most of which have higher multiplicities than the familiar K^{−}π^{+} decays. The experimental requirements are large acceptance for tracks, including those at small angles and low momentum, and good mass resolution. To separate the D K^{−} decays from D π^{−} decays will require good hadron identification at momenta around 2.3 GeV/c. Finally, it will be necessary to observe these decays with the D^{0} decaying to CP eigenstates such as K^{+}K^{−} and K_{S}η before one can attempt the CP measurement discussed above.

1.3.2 Charm physics

Some of the most important charm physics will be the study of semileptonic decays, especially Cabibbo-suppressed decays. Already the decays D→Kℓν and D→K^{∗}ℓν are providing important information on the form factors for a heavy quark decaying into a light quark. This serves to improve the picture of the form factors in exclusive B→ρℓν decay. Even better would be to make good measurements of the decays D→ρℓν and D→ωℓν.

At present, the CLEO II sample for D^{0}→K^{−}ℓ^{+}ν_{ℓ}, obtained by using D^{∗+}→π^{+}D^{0}, is the largest in the world. The other Cabibbo-favored decays also use the D^{∗} tag technique to remove backgrounds, but in some D^{+} modes, the background is still fairly large. The ability to separate charm vertices will help somewhat, especially in the case of D^{+} decay, because of the long lifetime. CLEO has seen a good D^{+}→π^{0}ℓ^{+}ν_{ℓ} signal, but the other Cabibbosuppressed decays are obscured by their Cabibbo-favored counterparts. The decay modes D^{+}→ωℓ^{+}ν_{ℓ} and D^{∗+}→π^{0}D^{+}, D^{+}→K^{∗−}ℓ^{+}ν_{ℓ} require good detection of soft π^{0}’s and charm vertex separation to improve the measurements.

A rare charm decay mode of special importance is the pure leptonic decay D_{s}^{+}→μ^{+}ν_{μ}, which measures the important decay constant f_{D_{s}}. This decay has been observed in CLEO II but will be measured with much higher precision in the upgraded CLEO experiment using the D_{s}^{∗}→γD_{s} decay. It will require excellent measurement of the soft photon parameters, and the ability to reject photons from π^{0} decay. It will also require good reconstruction of the missing neutrino using the technique of hermetic measurement of missing momentum.

1.3.3 Tau physics

At present, CLEO II is providing competitive information on almost every aspect of τ physics including decays, lifetime, and mass. With the improved CESR luminosity, CLEO will continue to have the largest τ sample in the world. The addition of particle identification will increase the number of modes that will be measured well. One measurement which should continue to improve is the limit on the ν_{τ} mass. In particular, the mode τ^{−}→(K^{∗}K)^{−}ν_{τ} will provide a good limit because of the large hadronic mass and the reasonable rate; the mode τ^{−}→(5π)^{−}ν_{τ} is another good possibility. Each utilizes different aspects of the detector performance.

Much of the advantage that CLEO has in measuring τ or D branching ratios is its good efficiency and resolution for events with one or more neutrals. We have ensured that in optimizing other, newer aspects of the detector, we lose no performance in this area. A good test of this is the decay mode τ^{−}→ηπ^{0}ν_{τ}, which requires a final state with four photons. In all cases the CLEO III detector meets or exceeds the performance of CLEO II.

2 Physics at a Symmetric 10^{33} e^{+}e^{−} Collider

In this chapter we discuss the physics goals using the CESR Phase 3 symmetric e^{+}e^{−} collider, capable of producing a peak luminosity of 1−−2×10^{33}cm^{−2}sec^{−1} at the Υ(4S). At this luminosity, the program will cover all aspects of heavy quark and tau lepton physics with high precision; the CLEO III experiment will continue to be at the very forefront of testing and understanding the Standard Model and beyond. Measurements of rare decays and form factors, precision measurements of the CKM parameters V_{cb} and V_{ub}, and the possibility that CP violation may be observed in the B system are just some of the crucial and exciting physics prospects.

To quantitatively evaluate the physics possibilities, we first make some assumptions about the future performance of CESR. We assume that CESR Phase 3 will have provided integrated luminosities of 20 fb^{−1} on the Υ(4S) resonance and 10 fb^{−1} on nearby continuum within a few years after commissioning. This comes after our projection that CLEO II will have accumulated 5 fb^{−1} of integrated luminosity and half that amount off the peak during CESR Phase 2. The physics program of CLEO III will have the advantage of an improved detector with excellent charged hadron identification and precision vertex information. The CLEO III detector thus provides the new tools necessary to probe rare processes and extract most of the important new results we foresee. The following sections outline the topics in which precision can be improved or new discoveries can be made.

2.1 B decay studies

2.1.1 Goals of B decay studies

Does the Standard Model explain quark mixing?

There are presently two fundamental questions in the forefront of high energy physics. One concerns the origin of mass which revolves around the scalar sector. This is being pursued in Higgs particle searches at super-high energies at Fermilab and CERN. The other equally deep and probably related question concerns the origin of quark mixing and, thus, CP violation. CLEO III with its unique ability to measure rare processes is in a position to pursue this question by performing redundant measurements of the unitary triangle shown in Fig. 4. These measurements couple our studies with the important question of baryonantibaryon asymmetry in the universe and thus with the very reasons that we may exist.

The unitary triangle is depicted in a two dimensional space where the Wolfenstein[4] parameters ρ and η form the horizontal and vertical axes. The sides and angles of the triangle are directly related to physical measurements. One side is proportional to the absolute value of the ratio of CKM elements V_{ub}/V_{cb}. This can be measured by observing the relative yields of b → u versus b → c in specific processes, such as semileptonic decay.

Figure 4: The unitary triangle shown in ρ vs. iη space.

Figure 5: Radiative penguin diagrams.

Another side is proportional to the ratio V_{td}/V_{ts} which can be inferred by measuring the ratio of mixing rates for B_{s}^{0} and B_{d}^{0} or by measuring the ratio of widths Γ(B → ργ)/Γ(B → K^{∗}γ). The three angles of the triangle can be found by measuring CP violation in three different types of B decay. Typical modes are shown in Fig. 4. Whereas the angles ϕ_{c} and ϕ_{u} can be determined at the Υ(4S) with an asymmetric collider, the angle ϕ_{s} can be discovered with a symmetric machine using charged B decays. Additional information constraining the unitary triangle is available from other measurements including ϵ and ϵ′/ϵ in kaon decay and B_{d} mixing, but these typically have two problems associated with their interpretation: they are functions of the unknown top quark mass and the hadronic matrix elements that relate the CKM parameters to the measurements need to be calculated theoretically. The measurements of the sides of the triangle have the latter problem, although to a lesser extent.

Physics beyond the Standard Model

B mesons can decay via loop diagrams which are sensitive to new physics, as well as physics predicted by the Standard Model. The existence of one such class of diagrams, called “penguins,” was predicted to explain the Δ I = 1/2 rule in kaon decays[5].

In Fig. 5 we show the Feynman diagram for the penguin process. Such decays have recently been established in CLEO II with evidence for the exclusive channel B → K^{∗}γ[1]. The study of radiative penguin processes, in particular measurements of the inclusive radiative penguin rate, b → sγ, are important for setting constraints on models beyond the Standard Model. As the accumulated luminosity increases, the study of the exclusive penguin process will become viable as will the search for other non-standard processes such as B → Kμe.

Table 1: Form-factor ratios at q_{max}^{2} from CLEO. The q^{2} dependence used in the fits of the data has been assumed to follow the models indicated.

A_{2}/A_{1}  V/A_{1}  |V_{cb}|A_{1}
Data fit for ISGW 1.02 ± 0.24 1.07 ± 0.57 0.035 ± 0.003 ± 0.002
Data fit for HQET* 0.79 ± 0.28 1.32 ± 0.62 0.037 ± 0.003 ± 0.002
ISGW[7]  1.14  1.27
KS[8]  1.39  1.54
WSB[9]  1.06  1.14
HQET  1.26  1.26
HQET*  1.14  1.74

QCD physics

Although the study of weak decays of b-quarks is the primary motivation for CLEO III, the b-quark is probably the only laboratory to study Quantum Chromodynamics (QCD) in heavy quark systems. The charm quark is probably too light to show the asymptotic behavior expected in Heavy Quark Effective Theory[6] (HQET) and the top quark is too heavy to form mesons prior to its decay. Moreover, a thorough understanding of QCD effects is necessary to extract the fundamental CKM parameters V_{cb} and V_{ub} with precision.

2.1.2 V_{cb} and form-factors in semileptonic decays

In purely hadronic B decays the virtual W from the b → c (or b → u) transition materializes as a quark-antiquark pair. In addition, the spectator-antiquark participates in the decay process. The result is a 4-quark system that can interact in a myriad of ways via the exchange of gluons. In such a system, predictions of branching fractions or even ratios of branching fractions are difficult. In semileptonic decays the virtual W materializes as ℓ^{−}ν̄_{ℓ} and thus only the c-quark (or u-quark) spectator-antiquark system needs to be modeled. Thus, semileptonic B decays provide the best opportunities for measurement of the CKM parameters V_{cb} and V_{ub}. For V_{cb} CLEO has shown that an excellent procedure is to use the channels B̄^{0} → D^{∗+}ℓ^{−}ν̄_{ℓ} and B^{−} → D^{∗0}ℓ^{−}ν̄_{ℓ}. These exclusive channels are selected by calculating the missing mass squared recoiling against the observed D^{∗} and ℓ^{−}. Missing mass distributions from the CLEO II data for these decays are shown in Fig. 6. Thus far, form-factor ratios have been inferred from the data and they are given in Table 1.

It is important to measure both the q^{2} dependence and the values of the form-factors at q_{max}^{2} in order to compare with HQET and to extract values of V_{cb}. HQET predicts one universal form-factor function, dependent on q^{2} and absolutely normalized at q^{2} = q_{max}^{2}, which describes all B → D^{∗}(D)ℓν̄ decays[10]. One can take advantage of HQET to measure

Figure 6: Signals for D^{∗+}ℓν (top) and D^{∗0}ℓν (bottom) from CLEO II. All combinatoric backgrounds have been subtracted. The histogram is the data, the dashed line is the D^{∗}ℓν signal and the dotted line D^{∗∗}ℓν signal from a fit to the data.

Figure 7: (τ_{B}/1.49ps)^{1/2}|V_{cb}|ξ(y) versus y extracted from CLEO II data for B̄^{0} → D^{∗+}ℓ^{−}ν̄_{ℓ}.

V_{cb} by extrapolating the q^{2} dependence of \frac{dΓ}{dy}(B → D^{∗}ℓν) to q_{max}^{2} as shown in Fig. 7. The intercept at q_{max}^{2} yields V_{cb}. Note, y = (m_{B}^{2} + m_{D^{∗}}^{2} + −q^{2})/2m_{B}m_{D^{∗}} and y(q_{max}^{2}) = 1.

One problem with this analysis is the background from the D^{∗∗} (or D^{∗}π) final states. This background can be eliminated by using only events with negative missing mass squared. This, however, results in a loss of half the signal. Improvements in background rejection are clearly necessary. Good charged particle identification will improve the D^{∗} signal; vertex detection can be used to reduce backgrounds by requiring that the D and the ℓ^{−} form an acceptable B vertex.

2.1.3 Rare B decays

Precision measurement of V_{ub}

Thus far the only evidence for the b → u transition has been in the observation of leptons with momenta above that allowed for b → c transitions. Recent measurements from CLEO II show somewhat smaller signals than reported originally[11]. Figure 8 shows the CLEO II signal and Table 2 provides a summary of recent measurements. The extracted values of V_{ub}/V_{cb} are based on the Altarelli model[12], with the results of different models differing by more than ± 50%.

CLEO has searched for evidence of the exclusive semileptonic reactions B → ρ(ω)ℓ^{−}ν̄_{ℓ}, for which ARGUS has claimed an observation[15]. CLEO does not observe these final states and quotes an upper limit B(B^{−} → V^{0}ℓ^{−}ν̄_{ℓ}) < (1.6−−2.7)×10^{−4} at the 90% confidence level[16].

Figure 8: Lepton endpoint spectrum from CLEO II after a) a strict and b) an R_{2} < 0.3 continuumsuppression cut has been applied. The data represent the sum of electrons and muons. The filled points are data on the Υ(4S) while the open points are scaled continuum data. The dashed curves are fits to the off-resonance data and the solid curves show the predicted b → cℓν contribution added to the fits of the off-resonance data.

Table 2: Summary of V_{ub}/V_{cb} measurements in Altarelli model

Experiment  V_{ub}/V_{cb}
CLEO 90[13] 0.100±0.015
ARGUS 90[14] 0.10 ±0.01
CLEO 92  0.075±0.008

The range of values reflects differences in theoretical models (see Fig. 9). With twenty times the statistics expected from the CESR upgrade and the addition of background reduction supplied by particle identification and vertex detection, we should be able to measure the branching ratios, but probably not the form-factors.

Theoretical input on these crucial B decay parameters may be obtained from applying HQET to the corresponding charm decays D → ρ(ω)ℓ^{+}ν_{ℓ}, which CLEO III can measure, and possibly from studies of b → sγ.

Measurement of “penguin” processes

Radiative b decays can occur only via the “penguin” diagrams shown in Fig. 5. The simplest such decay is B → K^{∗}γ, since the decay Kγ is forbidden by angular momentum conservation. CLEO has recently made the first observation of this process[1]. The data are shown in Fig. 10 and the measured branching ratio is (4.5 ± 1.5 ± 0.9) × 10^{−5}.

There are also hadronic penguins. The diagrams for these processes can be obtained by replacing the photon with a gluon. Examples of such processes are B̄^{0} → K^{−}π^{+} and the Cabibbo-suppressed decay B̄^{0} → π^{+}π^{−}[17]. Interference between the rare “penguin” processes and the b → u transitions may provide a means of seeing CP violation in B decays. CLEO has recently shown evidence that these processes do exist[18].

The Standard Model requires penguins to explain K decay and predicts B decays that occur only via penguins to have branching ratios in the 10^{−5} range. The inclusive branching ratio B(b → sγ) is predicted to be ∼ 4×10^{−4}[19]. The shape of the inclusive hadronic mass spectrum[20] (see Fig. 11) and exclusive branching ratios[21] (see Table 3) have also been calculated. Measuring the branching ratios for explicit channels determines values of the form-factors at q^{2}=0. These are intimately related to the form-factors in b → u exclusive semileptonic decays and thus may be useful in pinning them down.

Another exciting possibility is to determine the length of one side of the CKM triangle by measuring the penguin process B → ργ. This, when combined with B(B → K^{∗}γ), could

Figure 9: Invariant mass spectra for searches for (a) ω and (b) ρ in exclusive semileptonic decays from CLEO.

Figure 10: The B candidate mass spectrum for K^{∗}γ final states. The final states are coded as follows: shaded, K_{S}π^{−}; solid, K^{−}π^{+}; and open, K^{−}π^{0}.

Figure 11: Predicted mass spectra of b → sγ from Ali and Greub[20].

Table 3: Exclusive radiative penguin rates

Name  J^{P} B(B → K^{∗}γ)/B(B → sγ) B × 10^{5}
K  0^{−}  forbidden  forbidden
K^{∗}(890) 1^{−}  (3.5–12.2)%  1.4–4.9
K^{∗}(1410) 1^{−}  (7.2–10.6)%  2.9–4.2
K^{∗}(1680) 1^{−}  (1.1–1.5)%  0.4–0.6
K_{2}(1580) 2^{−}  (4.5–6.4)%  1.8–2.6
K^{∗}(1430) 0^{+}  forbidden  forbidden
K^{∗}(1270) 1^{+}  (4.5–10.1)%  1.8–4.0
K_{1}(1400) 1^{+}  (6.0–13.0)%  2.4–5.2
K_{2}^{∗}(1430) 2^{+}  (17.3–37.1)%  6.9–14.8
Sum  (44.1–90.9)%  17.6–36.6

be used to determine the ratio V_{td}/V_{ts}. Clearly, charged hadron identification is essential to distinguish the ρ from the K^{∗} and it also very useful in reducing backgrounds in all of the rare decay channels.

As an illustration of the sensitivity of the penguin process to high mass objects beyond the Standard Model we show in Fig. 12 the implications for the two Higgs doublet model. In such a model a charged Higgs boson H^{±} can substitute for the W^{±} in the loop. The figure shows the region in the space of mass m_{H} versus ratio of vacuum expectation values tan β that is consistent with the CLEO II limits on b → sγ branching ratio. Masses less than 250 GeV/c^{2} are excluded for this model.

Another example is the search for a supersymmetric squark. Figure 13 shows the region in the squark mass versus Higgs mass space[26] excluded by the CLEO II b → sγ limits. Only the range m̃/m_{H} > 1 is excluded.

Higher luminosity will increase the sensitivity to more loop decays and will allow a more precise comparison with Standard Model predictions, thus improving our sensitivity to high mass extensions or violations of the Model. This opens a window to new physics that is competitive with the much more expensive approach of building higher energy colliders.

The recent CLEO II observation of the sum of B̄^{0} → K^{−}π^{+} and B̄^{0} → π^{−}π^{+} mentioned above [2] is consistent with a Kπ loop decay branching ratio, (1.1 ± 0.7) × 10^{−5}. With the CESR/CLEO upgrade this measurement will improve as a result of the increased luminosity and the new high momentum particle identification system. Also we can expect many other b → sg loop decay modes to show up. Table 4 shows some of the theoretical predictions.

Decays to baryon-antibaryon pairs

The B can decay to baryons and, therefore, offers a unique opportunity to study the Table 4: Predictions for some exclusive hadronic loop decays

Mode  B(10^{−5})[22] B(10^{−5})[23]
B̄^{0} → K^{−}π^{+} 1.8 2.5
K̄^{0}π^{0}  0.7
K̄^{0}ϕ  1.5
K^{∗−}π^{+}  1.8  2.2
K^{∗−}ρ^{+}  1.6  2.7
K̄^{∗0}ρ^{0}  0.8
K̄^{∗0}ϕ  4.2
B^{−} → K^{−}π^{0} 1.2 1.2
K̄^{0}π^{−}  1.4
K^{−}ϕ 1.5 1.5
K^{∗−}π^{0}  0.9
K̄^{∗0}π^{−}  0.8
K^{∗−}ρ^{0} 0.8 1.4
K̄^{∗0}ρ^{−}  1.6
K^{∗−}ω  1.3
K^{∗−}ϕ  4.2  2.1

Figure 12: Predictions of the branching ratio B for b → sγ in the two Higgs doublet model[24], for various values of Higgs mass and the ratio tan β of vacuum expectation values, with the top mass assumed to be 150 GeV/c^{2}. CLEO II measurements[25, 1] allow 0.5 × 10^{−4} < B < 5.4 × 10^{−4}.

Figure 13: Predicted branching ratio for b → sγ (the labels on the curves) assuming a supersymmetric squark of mass m̃ and a Higgs boson of mass m_{H} [26] for a top quark mass of 150 GeV/C^{2}. The CLEO II data[1, 25] allow 0.5 × 10^{−4} < B < 5.4 × 10^{−4}.

dynamics of baryon pair creation in a reaction where at least the initial state and the initial coupling are well defined. Even at the most fundamental level we are completely ignorant of the production mechanism. For example, does it proceed through diquark-antidiquark pairs? There are data from CLEO and ARGUS on inclusive proton, Λ, and Λ_{c} yields in B decay [27] but because of the low efficiencies for reconstructing exclusive decays of B mesons and charmed baryons, a definitive study must await the larger data samples from the upgraded CESR.

Purely leptonic decays

In the Standard Model the purely leptonic decay of the B into a lepton and a neutrino should proceed as in pion decay, through the annihilation of the b and ū quarks into a W^{−}, materializing as ℓ^{−} + ν̄. The rate is proportional to |V_{ub}|^{2} and to the bū bound-state overlap factor f_{B}^{2}. Once V_{ub} is known, a measurement can give us an experimental value for f_{B}, which comes into calculations of mixing, CP violation, and other weak processes involving the interaction between the quarks in the B meson. Theoretical predictions for f_{B} from QCD sum rules[28] and lattice calculations[29] cover the range from 105 to 187 MeV.

The leptonic decay rate is also sensitive to new physics; for example, the mass m_{tπ} and coupling λ of a possible technipion can be limited [30] by

\frac{m_{tπ}}{λ} = 200 GeV (\frac{10^{−6}}{B(B → τν̄)})^{1/4} .

Since the rate should be proportional to the square of the lepton mass, the most copious mode is B → τν̄, which however is identifiable only in tagged BB̄ events, because of the missing neutrinos. The much smaller branching ratio for B → μν̄, about 10^{−7}, might be measurable without tagging, since the muons will be nearly monoenergetic near the BB̄ threshold. Very high luminosities may be required, however, since a B branching ratio of 10^{−7} yields only about one event per 10 fb^{−1} at the Υ(4S). Non-Standard Model enhancements, however, could give a much larger event rate.

Forbidden dilepton decays

Certain B decays are strictly forbidden in the Standard Model, for example, those that violate an accepted conservation law. Still other decays may be allowed but at an undetectably low branching ratio. In the cases in which the particular final state is easily recognizable experimentally and not obscured by backgrounds, the search for such modes can expose weaknesses in the model. Typical examples are B → μe and B → τμ, which are forbidden by the absence of transitions between lepton families in the Standard Model. If there were a “horizontal gauge boson” H^{0} that would allow bd̄ → H^{0} → μe, the branching ratio for B → μe would be[31]

130 (\frac{m_{W}}{M_{H}} )^{4} (\frac{f_{B}}{100 MeV} )^{2} .

The present CLEO[32] upper limit of 5.9 × 10^{−6}, limited only by the size of the available event sample, implies a lower limit on the mass of the horizontal gauge boson of 8 TeV. The mass sensitivity increases as the fourth root of the number of B decays searched.

The decays B → e^{+}e^{−} and B → μ^{+}μ^{−} violate the Standard Model GIM suppression of flavor changing neutral currents, but are allowed through loop diagrams at the level of 10^{−12} and 10^{−8} respectively[33]. The CLEO upper limits, 5.9 × 10^{−6} for both[32], can be improved simply by searching larger data samples.

2.1.4 CP violation

Possibilities for observing CP violation and measuring ϕ_{s}

Gronau and Wyler[3] have proposed a method of measuring ϕ_{s} using only measured branching fractions in charged B decay[34]. The method is best illustrated by considering the three related B^{−} decay modes, and the corresponding modes for B^{+}, B^{−} → D^{0}K^{−}, B^{−} → D̄^{0}K^{−}, and B^{−} → D_{CP} K^{−}, where D_{CP} indicates that the D^{0} decays into a CP eigenstate. The weak decay diagrams[35] for the first two modes are shown in Fig. 14. The first decay is a Cabibbo-suppressed version of B^{−} → D^{0}π^{−}, while the second is a color suppressed b → u transition where the virtual W^{−} transforms itself into a c̄s pair.

Interference is possible between the two decay modes if the D^{0} decays into a nonstrangeness specific mode. Examples of such final states include K^{+}K^{−}, K_{S}π^{0}, and K_{S}π^{+}π^{−}. To simplify the discussion, we only use such states which are in specific angular momentum configurations so that their CP is defined as +1 or -1; these states are usually denoted as

Figure 14: Quark level decay diagrams for B → DK^{−}.

D_{1}^{0} and D_{2}^{0}. We have that

D_{1}^{0} = \frac{1}{\sqrt{2}} [D^{0} + D̄^{0}] , and D_{2}^{0} = \frac{1}{\sqrt{2}} [D^{0} − D̄^{0}] .

To further simply the discussion we will consider only D_{1}^{0}’s. The amplitudes for the relevant B^{±} decay modes are related by

\sqrt{2}A_{1}^{+}(B^{+} → D_{1}^{0}K^{+}) = A(B^{+} → D̄^{0}K^{+}) + A(B^{+} → D^{0}K^{+})  (6)
\sqrt{2}A_{1}^{−}(B^{−} → D_{1}^{0}K^{−}) = A(B^{−} → D̄^{0}K^{−}) + A(B^{−} → D^{0}K^{−}).  (7)

The branching fractions for the decays B^{+} → D_{1}^{0}K^{+} and B^{−} → D_{1}^{0}K^{−} need not be equal and a difference between them would be a manifest demonstration of CP violation in charged B meson decay. Two amplitudes are equal in magnitude,

|A| ≡ |A(B^{+} → D^{0}K^{+})| = |A(B^{−} → D̄^{0}K^{−})|  (8)
|Ā| ≡ |A(B^{+} → D̄^{0}K^{+})| = |A(B^{−} → D^{0}K^{−})|.  (9)

Then denoting the relative hadronic phase between the amplitudes A and Ā by δ[37] we have

\sqrt{2}A_{1}^{+} = |A|e^{i(+ϕ_{s}+δ)} + |Ā|  (10)
\sqrt{2}A_{1}^{−} = |A|e^{i(−ϕ_{s}+δ)} + |Ā|.  (11)

To determine ϕ_{s} we need to measure the branching fractions for the modes in Eqs. 6 and 7. Two triangles are constructed from the square roots of the branching fractions as illustrated in Fig. 15. The angles ϕ_{s} and δ are then determined.

This method leads to a four-fold ambiguity in the value of sin ϕ_{s}. Half the ambiguity is removed if one uses the measurement of ϵ in K^{0} decay which leads to the boundary condition sin ϕ_{s} > 0. The other ambiguity arises between the determination of ϕ_{s} and δ. This ambiguity can be removed by measuring other related decay modes, such as B^{−} → D^{∗0}K^{−}, which are likely to have a different strong phase shift δ but must have the same CP violating phase. It

Figure 15: The two triangles constructed out of the four different measured rates and the phase relationships between the two triangles.

is also important to realize that this method can find a value for ϕ_{s} even if δ is zero. In this case the two triangles for B^{−} and B^{+} are congruent.

Observation of CP asymmetry is expedited by having comparable amplitudes for the two interfering decays. At the CKM level (ignoring color suppression), the magnitude of the amplitude Ā is V_{cb}V_{us} = aλ^{3}, while the magnitude of the amplitude A is |V_{ub}|V_{cs} ≈ (|V_{ub}/V_{cb}|)V_{cb} ≈ 0.08aλ^{2} where a and λ are the parameters defined by Wolfenstein[4, 38, 39]. Therefore at the CKM level the ratio of the amplitudes is |A/Ā| ≈ 0.08/λ ≈ 1/3. However, Ā occurs through a color-allowed diagram while the diagram for A is color-suppressed. However, there is a simple way to predict the branching fraction corresponding to the amplitude A. Consider the measured decay mode, B^{−} → ψK^{−}, which is very similar to the mode B^{−} → D̄^{∗0}K^{−}. The decay diagrams are shown in Fig. 16. The difference between the two diagrams is that the ψ decay occurs when there is a b → c transition, while the D̄^{∗0} occurs when there is a b → u transition. We choose to discuss D̄^{∗0} here rather than D̄^{0} since D̄^{∗0} is a vector meson as is the ψ. Correcting for phase space effects we have

B(B^{−} → D̄^{∗0}K^{−}) = B(B^{−} → ψK^{−}) |\frac{V_{ub}}{V_{cb}} |^{2} (\frac{f_{D^{∗}}}{f_{ψ}} )^{2} · P,  (12)

where P is a phase space factor equal to 1.3, and the f’s are form-factors. These form-factors are not known; theoretical predictions give (f_{D^{∗}}/f_{ψ})^{2} ≈ 50%[45]. Using B(B^{−} → ψK^{−}) ≈ 1.1×10^{−3}[41], the prediction is that B(B^{−} → D̄^{∗0}K^{−}) = (4.8±2.7)×10^{−6}. The error comes from the uncertainty on the ψK^{−} branching ratio and the uncertainty in V_{ub}/V_{cb}. A similar prediction can be made for the D̄^{∗0}K^{∗−} final state and is listed in Table 5.

The final states with a scalar D^{0} rather than a vector D^{∗0} are more difficult to predict because the analogous b → c mode, η_{c}K, has not been observed, nor are there any upper limits. There are two theoretical predictions that inclusive η_{c} modes are half the ψ modes[40]. The estimates given in Table 5 use the assumption that is also true for the exclusive modes. For the sum of branching ratios we use a value of 18 × 10^{−6} for subsequent estimates.

To estimate the branching fraction B(B^{−} → D^{0}K^{−}) with amplitude of type Ā, we take the new CLEO measurement[41] B(B^{−} → D^{0}π^{−}) ≈ 0.47% and multiply it by the Cabibbo suppression factor λ^{2} and the enhancement factor (f_{K}/f_{π})^{2} ≈ 1.5[42]. The result is B(B^{−} → D^{0}K^{−}) ∼ 3.3 × 10^{−4}.

In a 20 fb^{−1} data sample, the expected number of events is given by the product of

Figure 16: Decay diagrams for (a) B^{−} → ψK^{−} and (b) B^{−} → D̄^{∗0}K^{−}.

Table 5: Branching ratio estimates for B^{−} → D̄K^{−} modes (with amplitudes of type A).

B^{−} Mode  B
D̄^{0}K^{−}  2.4 × 10^{−6} ^{†}
D̄^{∗0}K^{−}  (4.8 ± 2.7) × 10^{−6}
D̄^{0}K^{∗−}  3.5 × 10^{−6} ^{†}
D̄^{∗0}K^{∗−}  (7.0 ± 4.3) × 10^{−6}
Sum  ∼ 18 × 10^{−6}

† This estimate is half the corresponding rate with a D̄^{∗0}.

Table 6: CP eigenstate modes of D^{0} decay.

mode B(%) ϵ  B × ϵ(%)
K^{+}K^{−} 0.5 0.74  0.37
K_{S}ω  1.9 0.09  0.16
K_{S}π^{0}  1.4 0.15  0.21
K_{S}η′  1.0 0.04  0.04
K_{S}ϕ  0.4 0.18  0.07
K_{S}η  0.4 0.07  0.03
K^{∗0}π^{0†} 0.8 0.10  0.08
π^{+}π^{−} 0.2 0.74 0.15
SUM  1.11

† The branching ratio is only for the K_{S}π^{0} decay.

several factors, including the integrated luminosity, the Υ(4S) cross section (1 nb), the D̄^{0} detection efficiency (0.09), the efficiency for finding a single charged pion (0.9), and other cuts used to enhance signal to background. The D̄^{0} detection efficiency is calculated using the three modes K^{−}π^{+}, K^{−}π^{+}π^{0}, and K^{−}π^{+}π^{+}π^{−}. The cuts used to enhance the signal to background include a cut on the momentum direction of the B candidate compared with the thrust axis of the rest of the event. These have an efficiency of 80%. This gives a total of 400 such events summed over B^{−} and B^{+} decays[43]. We can add other decay modes, although the analysis must be carried out for them individually because of the ambiguity between ϕ_{s} and δ in each mode. Using more than one mode provides a way of resolving this ambiguity. The other obvious final states are D̄^{∗0}K^{−}, D̄^{0}K^{∗−}, and D̄^{∗0}K^{∗−}. Summing all of these modes leads to approximately 1600 events. The number of type A decays, based on the summed branching ratio of 18 × 10^{−6}, is 20. These rate estimates may be conservative in that they use the current CLEO II efficiencies.

The CP eigenstate decay modes are more prolific. Table 6 shows that the branching ratio times efficiency for these decays is approximately 10% of that of the D^{0} decays used above.

In order to evaluate how accurate a measurement of sinϕ_{s} can be made with 20 fb^{−1}, we use the expression given by Gronau and Wyler[3] for determining sin ϕ_{s} from |A|, |Ā|, and |A_{1}^{±}|. The numbers of events going through the channels Ā and A have been taken to be 800 and 20. (The use of 800 rather than 1600 events is based on lower measurements of type Ā than are currently available; this change is not expected to significantly affect the accuracy of the measurement.) Choosing values for ϕ_{s} and δ defines the number of events in the A_{1}^{+} and A_{1}^{−} channels. These latter values are calculated applying the law of cosines to the triangles in Fig. 15. For example for ϕ_{s} = δ = 90^{0}, the number of events in the A_{1}^{+} and A_{1}^{−} channels is 53 and 29, respectively, after allowing for the 10% relative detection efficiency of D_{CP} to D^{0}.

The procedure used is to perform many “experiments” starting with initial central values of 800 and 20 events for the A and Ā channel for specific values of ϕ_{s} and δ. Then we allow all measured event numbers to fluctuate statistically around their central values. The distribution in sinϕ_{s} for several choices of ϕ_{s} and δ is shown in Fig. 17. Taking r.m.s. values of σ(sin ϕ_{s}) < 0.16 as necessary for the observation of CP violation, the potential “discovery” region is shown in Fig. 18.

To be able to discover CP violation using this method requires a detector which has similar (but hopefully better) photon and charged particle reconstruction capabilities as CLEO II, excellent charged particle identification, and excellent vertex detection. The former two are important to increase the discovery potential or guard against smaller branching ratios. The latter two are necessary to reject backgrounds in these rather small signal samples.

It is not clear that particle identification and vertex detection will be sufficient to reduce the background to the required level. Examination of current CLEO data reveals that the equivalent branching ratio of the background in the D^{0}π^{−} and D^{∗0}π^{−} signal regions is ∼10^{−3}; this is large compared to the ∼ 18 × 10^{−6} branching ratio we are attempting to measure in the A channel. However, the particle identification of the current CLEO II detector is minimal for these final states and there is no vertex information on the D^{0} or the B. On the other hand, low background levels are possible, being almost non-existent in the ψK^{−} final state. If the background from fake D^{0} and non-K^{−} combinations could be eliminated, we would still be faced with the problem of the background from real D̄^{0}K^{−} combinations in continuum events.

Rate asymmetries

Any final state which can be produced with two different amplitudes can show CP violating effects. For example, it is possible for a “penguin” amplitude to interfere with a “tree” level amplitude. The measurement consists of finding a different number of events of B^{−} versus B^{+} events. For example, one could search for an asymmetry, A, in the conjugate decay modes K^{−}π^{0} and K^{+}π^{0} where

A = \frac{N(K^{−}π^{0}) − N(K^{+}π^{0})}{N(K^{−}π^{0}) + N(K^{+}π^{0})}

In B^{−} decays the final states could be K^{0}π^{−}, K^{−}π^{0}, etc..

The sample size for any particular mode is N = N_{BB̄}Bϵ, where B is the branching ratio and ϵ is the detection efficiency. The total number of BB̄ events required to get an s standard deviation effect in one mode is N_{BB̄} = s^{2}/BA^{2}ϵ. Provided that there is at least one mode with Bϵ > 2 × 10^{−5} and A > 20%, an integrated luminosity of 20 fb^{−1} would be sufficient to establish CP violation to 4 standard deviations.

Angular correlations in vector-vector decay modes

When there is a decay into two vector mesons, and if two amplitudes contribute, then it is possible for CP violation to be observed in differences between the angular distribu-28

Figure 17: Values of sin ϕ_{s} found for 1000 experiments and different input values as specified.

Figure 18: Region in ϕ_{s} − δ plane of possible discovery of CP violation in 20 fb^{−1}.

tions of the B and B̄. One way of looking for these asymmetries has been suggested by Valencia[44]. Defining the decay in terms of the momentum and polarization vectors as B → V_{1}(k, ϵ_{1})V_{2}(q, ϵ_{2}), one can look directly at several asymmetries for both the B and the B̄. For example

A_{B} = \frac{N(k · ϵ_{1} × ϵ_{2} > 0) − N(k · ϵ_{1} × ϵ_{2} < 0)}{N_{T otal}},

is defined for the B decay and a similar quantity, AB̄ defined for the B̄ decay. A difference between A_{B} and AB̄ would be evidence of CP violation.

Predictions of angular asymmetries in many channels such as K^{∗−}ω have been given by Kramer and Palmer[46]. The predictions are quite sensitive to hadronic phase shifts. Therefore, it is difficult to make quantitative predictions, but such effects should be explored.

CP-odd energy asymmetries

It has been suggested by Burdman and Donoghue that CP violation might show up in three body decays modes, accessible from both B^{0} and B̄^{0} decays, in Dalitz plot asymmetries[47]. One possible mode is B → K_{S}π^{+}π^{−}. A distribution of interest is the shape of the summed energy spectrum of the π^{+} and π^{−}. The average energy distribution can be written as ⟨E_{π}⟩ = a + b(E_{π^{+}} − E_{π^{−}}). Finding that b ≠ 0 is a manifest observation of CP violation. Again, there are no quantitative predictions.

2.2 Charm decay physics

The opportunities for performing charm physics studies at a high luminosity e^{+}e^{−} collider at a center-of-mass energy around 10 GeV are enormous especially when photons or π^{0}’s are involved. Just below B meson threshold, cc̄ events account for ∼ 40% of the total cross section. Thus an integrated luminosity of 10 fb^{−1} will yield ∼ 2 × 10^{7} charmed particles. On the Υ(4S) which decays to B mesons, most of the B’s decay to charmed particles increasing the yield further. This plus the excellent efficiency and resolution of the CsI calorimeter allows CLEO II and, therefore, CLEO III, to compete successfully with Fermilab fixed target experiments in situations where photon detection is important.

2.2.1 Semileptonic decays

Cabibbo-suppressed semileptonic decays provide an exciting laboratory for studying charm decays. These decays provide information on the ratio of CKM elements V_{cd}/V_{cs}. For decays to the vector final states, ρ(ω)ℓ^{+}ν, measurement of the form-factor ratios can be related directly to the same final states in B decays which can allow extraction of V_{ub} from b → u decays.

Already, the decays D → K(K^{∗})ℓν and D^{+} → π^{0}ℓ^{+}ν have been observed in CLEO II[48] and are providing important information on the form factors in charm decay. The approach is to use the decay chain D^{∗+} → π^{+}D^{0} or D^{∗+} → π^{0}D^{+} as a means to tag the D^{0} or D^{+} and to reduce background. Presently, the D^{0} → K^{−}ℓ^{+}ν_{ℓ} data sample is the largest in the world and a clean signal D^{+} → π^{0}ℓ^{+}ν_{ℓ} has been observed. Other Cabibbo-suppressed decays are obscured by their Cabibbo-favored counterparts. The π^{0}ℓ^{+}ν_{ℓ} signal is shown in Fig. 19 and is compared with the similar final state K^{0}ℓ^{+}ν. From these two signals CLEO finds V_{cd}/V_{cs}=0.21±0.02±0.02, where the second error results from the theoretical uncertainty in the value of the form-factor ratio at zero momentum transfer. Extending the semileptonic decay studies to the rarer ρ and ω final states requires excellent π^{0} tagging efficiency and high luminosity, both of which will be available to CLEO III.

2.2.2 Hadron spectroscopy

HQET provides a wealth of predictions concerning the spins and decay patterns of charmed mesons and charmed baryons. Both will be copiously produced in CLEO III. Precision measurements of mass differences[49] may be used to determine the decay constant f_{D}[50]. Presently the D^{∗} − D mass difference errors are of the order of 0.1 MeV/c^{2}. The CLEO III data sample of such decays should be more than an order of magnitude larger than that shown in Fig. 20 which comes from CLEO II.

CLEO III will thus provide the opportunity to measure the branching fractions, identify the excited states, and unravel the complexities of the weak decays of the D_{s}, Λ_{c}, Ξ_{c}^{0}, Ξ_{c}^{+}, and Ω_{c} mesons. This program for the D^{0} and D^{+}, which took approximately 15 years, is only now nearing completion.

The study of excited vector states (e.g., D^{∗∗}) usually means fitting decay angular distri-

Figure 19: The D^{∗+}-D^{+} mass difference distributions, ΔM, for the candidates (a) D^{+} → π_{fst}^{0}ℓ^{+}ν and (b) D^{+} → K̄^{0}ℓ^{+}ν. The dashed curves are fits to the background and the solid curves are fits to the data including signal and background.

butions, and hence requires an extremely large data sample. Such studies will be possible in CLEO III providing an opportunity to make stringent tests of QCD models of the spectroscopy of mesons containing one heavy quark.

Spectroscopy has historically yielded surprises and the high statistics data sample of CLEO III will make it an ideal place to search for qg bound states, glueballs, and other exotica.

2.2.3 Rare charm decays

The separated vertices for charm mesons together with precise vertex location, excellent particle identification, and excellent π^{0} reconstruction make it possible to reduce backgrounds and study rare decay modes not accessible in other experiments. One example is the doubly Cabibbo-suppressed decay D^{0} → K^{+}π^{−}. Without good (4σ) particle identification this process is not easily separated from the background process D^{0} → K^{−}π^{+}.

Other rare decays provide a unique opportunity for looking beyond the Standard Model. Flavor-changing neutral currents, manifested by the decay D → Xℓ^{+}ℓ^{−}, should exist and may have branching fractions of the order of 10^{−6}. Present limits by CLEO II are of the order of 10^{−4}.

Figure 20: The D^{∗} − D mass difference distributions (ΔM) for (a) D^{∗0} − D^{0} with D^{0} → K^{−}π^{+}, the histogram shows the data cut on the D^{0} mass, while the solid points show the sidebands. In (b) the sideband has been subtracted and the data fit to the signal plus plus background shapes described in the text; (c) and (d) are for D^{0} → K^{−}π^{+}π^{0} and (e) and (f) show the D^{∗+} − D^{+} mass difference with D^{+} → K^{−}π^{+}π^{+}.

2.2.4 Measurement of f_{D_{s}}

Of special importance is the purely leptonic decay D_{s}^{+} → μ^{+}ν. Measuring the width of this reaction determines the decay constant f_{D_{s}}. This decay constant can be calculated theoretically as can the decay constant f_{B}. In fact, lattice gauge calculations have made estimates of both numbers. Knowledge of f_{B} is crucial in testing Standard Model parameters. It is, in fact, very difficult to measure f_{B} directly since the purely leptonic decays needed to measure it directly proceed through the b → u channel and have extremely small branching ratios. It is possible, however, to measure f_{D_{s}} through its dominant c → s decay channel. CLEO II has made the first measurement of the ratio Γ(D_{s}^{+} → μ^{+}ν)/Γ(D_{s}^{+} → ϕπ^{+}=0.25 ± 0.03 ± 0.07, using D_{s}’s tagged by the decay D_{s}^{∗} → γD_{s}[51]. In order to extract a value for f_{D_{s}}, the value of the ϕπ^{+} branching ratio must be known. Using an estimate relating D_{s} semileptonic decays to D^{+} semileptonic decays, CLEO II finds f_{D_{s}} = (344±37±52±42) MeV. The last error is due to the uncertainty on the ϕπ^{+} branching ratio, (3.7± 0.9)%, which is consistent with more recent estimates[52]. CLEO III will be able to substantially reduce the errors on both the partial width ratio and the absolute branching ratio.

2.2.5 Charm mixing

Expectations for the D^{0}D̄^{0} mixing parameter range from ∼ 10^{−4} to 10^{−7} in the Standard Model[53]. Based on our present experience with CLEO II, we expect that sensitivities below 10^{−5} will be achieved with CLEO III. The most obvious method to detect such mixing is by observing the decay chain D^{∗+} → D^{0}π^{+}, D^{0} → Kπ. The sign of the π from the D^{∗+} decay tags the decay as either a D^{0} or D̄^{0}; then the sign of the K indicates whether the D^{0} has mixed before decaying. The main background comes from doubly Cabibbo-suppressed decays of the D^{0} with a rate proportional to tan^{4}θ_{c}=0.003, which is measured to be much larger than the mixing signal[54]. Nevertheless, a good measurement of the time dependence should easily separate the two processes. The conventional doubly Cabibbo-suppressed decays should follow a e^{−Γt} while for mixed events the distribution is

e^{−Γt} (1 − 2cos(Δmt)e^{ΔΓt/2} + e^{ΔΓt})  (13)

This measurement relies heavily on the vertexing capabilities of the silicon detector, on the angular resolution of the slow π, and on good π/K separation. A large value of the mixing parameter would indicate physics beyond the Standard Model.

2.3 Tau decay physics

Results from SLC and LEP indicate that there are only three generations of leptons[55] within the standard gauge theory of electromagnetic and weak interactions[56]. The heaviest of these leptons, the τ, serves as an excellent laboratory for studying many aspects of the

Table 7: Data sample for selected τ decay modes assuming an integrated luminosity of 30 fb^{−1}.

Decay Mode  Number of Events ×10^{6}
All τ’s  54
τ^{−} → e^{−}ν_{τ} ν̄_{e}  10.8
τ^{−} → μ^{−}ν_{τ} ν̄_{μ}  10.6
τ^{−} → π^{−}ν_{τ}  7
τ^{−} → K^{−}ν_{τ}  0.4
τ^{−} → K^{∗−}ν_{τ}  0.9
τ^{−} → ηX^{−}ν_{τ}  > 0.1
τ^{−} → ωX^{−}ν_{τ}  1
τ^{−} → K^{−}K^{+}π^{−}ν_{τ}  ∼ 0.13
τ^{−} → 3 charged tracks  8
τ^{−} → 5 charged tracks  0.06

Standard Model[57]. Of all the known leptons only the τ is massive enough to decay into both purely leptonic states and semileptonic states with accompanying hadrons. The purely leptonic decays allow the study of lepton universality and the τ − ν_{τ} − W vertex, while the semileptonic decays permit the investigation of the weak hadronic current and strong interaction effects. The decay of the τ allows the measuring of the ν_{τ} mass and the search for Standard Model forbidden decays involving three generations of leptons with or without accompanying hadrons.

To date all meaningful studies of the properties of the τ have been carried out at e^{+}e^{−} colliders. At these facilities the cross section for τ production is relatively large while the associated backgrounds are both calculable and small. These advantages suggest that e^{+}e^{−} colliders will be the place to measure the decay characteristics of the τ for many years to come. For CLEO III we assume an integrated luminosity of 30 fb^{−1} for which there will be about 27 million τ pairs produced. In Table 7 we list the data samples which will be available for various τ decays. In this section, we discuss a selected sample of physics topics that can be studied at an upgraded CESR/CLEO facility. We will assume that the one-charged-particle decay discrepancy has been resolved and thus will not be discussed here.

2.3.1 Tau neutrino mass

The Standard Model has no predictive power on the neutrino masses. However, a non-zero neutrino mass is predicted in many of the fashionable extensions of Standard Model such as compositeness, Grand Unified Theory, and the left-right symmetry model[58]. A sufficiently massive neutrino is also a popular solution to the solar neutrino and dark matter problems. Since there are only three generations of leptons, ν_{τ} is the only potential candidate to solve these problems. The upper limit on the mass of ν_{τ} is 31 MeV/c^{2} at 95% confidence level[59]; for comparison, the mass limits on ν_{e} and ν_{μ} are 17 eV/c^{2} and 270 KeV/c^{2}, respectively[42]. In some models, an assumption is made for the mass hierarchy

m_{ν_{τ}} = m_{ν_{e}} × (\frac{m_{τ}}{m_{e}} )^{2} .

Given m_{ν_{e}} = 17 eV/c^{2}, m_{ν_{τ}} is expected to be 206 MeV/c^{2}. Therefore, the limit on m_{ν_{τ}} already exceeds the sensitivity of ν_{e} and ν_{μ} limits in these models. A sensitivity to m_{ν_{τ}} of 3 MeV/c^{2} corresponds to a sensitivity of 0.25 eV/c^{2} on m_{ν_{e}}. Measuring the τ neutrino mass, therefore, can be a sensitive probe of the phenomenon of neutrino mass.

The best probe of τ neutrino mass is the shape of the end-point mass spectrum of the hadronic decay products. One of the decay modes with sensitivity to m_{ν_{τ}} is τ^{−} → π^{−}π^{+}π^{−}π^{+}π^{−}ν_{τ}. In general, the mass spectrum of the hadronic system depends on the unknown intermediate resonance structure. Fortunately, the end-point mass spectrum[60, 61] in this decay is dominated by the weak decay matrix element M and the phase space factor Φ,

\frac{dΓ}{dm} ∝ mM^{2}Φ ,

where m is the mass of the hadronic system,

M^{2} = (m_{τ}^{2} − m^{2})(m_{τ}^{2} + 2m^{2}) − m_{ν_{τ}}^{2} (2m_{τ}^{2} − m^{2} − m_{ν_{τ}}^{2} ) ,

and

Φ^{2} = [m_{τ}^{2} − (m + m_{ν_{τ}})^{2}] × [m_{τ}^{2} − (m − m_{ν_{τ}})^{2}] .

The candidates for τ^{−} → π^{−}π^{+}π^{−}π^{+}π^{−}ν_{τ} can be selected with an e or μ tag with no photon above a certain threshold energy. Based on the experience with CLEO II, we expect a detection efficiency of ∼ 15%. For a data sample of 3 × 10^{7} τ^{+}τ^{−}, this corresponds to at least 2000 events in the final sample which allows the detailed examination and fitting of the full mass spectrum. This is much more satisfactory than previous methods which are dominated by the highest mass event (which may actually be background). With the expected mass resolution of ∼ 10 MeV/c^{2}, the sensitivity on m_{ν_{τ}} is a few MeV/c^{2}.

While τ^{−} → π^{−}π^{+}π^{−}π^{+}π^{−}ν_{τ} has historically been the favored decay mode to probe m_{ν_{τ}}, there are several other candidate reactions which may play an important role in this area that are accessible with the detector upgrades proposed here. In Table 8 we list several of these modes. In particular, the mode τ^{−} → K^{−}K^{+}π^{−}ν_{τ} is attractive due to the large minimum hadronic mass in the decay. At the present time the CLEO experiment is unable to use this mode as it lacks the necessary particle identification. A phase space Monte Carlo of τ^{−} → K^{−}K^{+}π^{−}ν_{τ} shows that only ∼1% of the decays have both kaons in the momentum interval p < 0.8 GeV/c where kaons can be unambiguously identified in CLEO II. Raising the momentum cutoff to 3 GeV/c in CLEO III improves the acceptance to 90%. Thus with the proposed CLEO III improvements in high momentum kaon identification, the mode τ^{−} → K^{−}K^{+}π^{−}ν_{τ} will compete with τ^{−} → π^{−}π^{+}π^{−}π^{+}π^{−}ν_{τ} as the standard channel to use for measuring m_{ν_{τ}}.

Table 8: Candidate τ decay modes for measuring m_{ν_{τ}}.

Decay Mode  Branching Ratio ×10^{−3} Minimum hadronic mass (GeV/c^{2})
π^{−}π^{+}π^{−}π^{+}π^{−}  0.6  0.7 or 1.68^{ab}
π^{−}π^{+}π^{−}π^{+}π^{−}n(π^{0})  1.1  0.7
π^{−}π^{+}π^{−}π^{0}π^{0}  ∼4.8  0.7 or 1.07^{c}
K^{−}K^{+}π^{−}  ∼2.2  1.13
K^{∗}(890)K^{−}n(π^{0})  3.2  1.39^{b}

a assuming ρ^{0}ρ^{0}π^{−} dominates.

b must take into account resonance width.

c for ωπ^{−}π^{0} submode.

2.3.2 Decay modes and branching ratios

The partial widths for most of the τ decay modes can be calculated with the present understanding of the Standard Model. The purely leptonic decays can be calculated rigorously, including radiative corrections. Most of the major semileptonic decays with accompanying hadrons can be calculated reliably using experimental data such as π and K decay constants and the cross section of e^{+}e^{−} annihilation into hadrons. Since some of the semileptonic decays cannot be calculated accurately, the total decay width is unknown. It is therefore customary to express the branching ratios normalized to the electron branching ratio B_{e}. Alternatively, the dependence on B_{e} can be eliminated by taking the ratio of two branching ratios. This is actually a preferred method of comparing the experimental measurement with the theoretical expectation because many systematic errors cancel in the ratio. In fact, some of the current experiments may be systematics limited after collecting data for a few more years. Since we do not expect any dramatic improvement in the systematic errors in the large data sample expected with the detector under discussion here, comparing the ratio is the only reliable method for the precision test of the prediction. In this section, we discuss the expected precision of the ratios. Also included is a discussion on the sensitivity on the highly suppressed second-class-current decay and the forbidden lepton-number-violating decays.

Tau decays to e and μ

In the Standard Model, the purely leptonic decays of the τ can be calculated rigorously, including the electroweak radiative corrections. This allows a precise test of lepton universality. The branching ratios for the two decays differ by a small phase space factor[60, 62],

\frac{B_{μ}}{B_{e}} = 1 − 8y + 8y^{3} − y^{4} − 12y^{2} ln y = 0.973 ,

where y = (m_{μ}/m_{τ})^{2}. The current ratio is 0.988 ± 0.022[42].

At the present time the most precise measurement of the electronic branching fraction B_{e} is from CLEO II using a data sample in which both τ’s decay into electrons[63]. This ee analysis exploits the power of the CsI calorimeter and charged particle tracking system to reduce the backgrounds from radiative Bhabha and μ pair events to below 1%. A measurement of B_{μ} using a similar technique is now in progress. Since the momentum of a μ candidate is required to be greater than ∼1 GeV/c for proper identification, the same momentum cut must be imposed on the e candidates for cancellation of the systematic uncertainties from tracking in B_{μ}/B_{e}. Based on the CLEO II experience, we expect to obtain some 6 × 10^{4} ee events and a similar number of μμ events from 3 × 10^{7} τ pairs; this corresponds to a statistical precision of 0.28% on the ratio B_{μ}/B_{e}. The background from other τ decay modes is expected to be less than a few percent and hadronic backgrounds are expected to be less than 1%, contributing equally to both samples. The overall systematic error in B_{μ}/B_{e} from the background corrections should be less than 0.1%. The other main source of systematic error is the detection efficiency which has two components: trigger and particle identification. The uncertainties in the e and μ identification efficiency do not cancel in the ratio because of the very different techniques involved. Detection efficiency uncertainty in the ratio could be known to 0.7% and this will dominate the overall precision.

An alternative method would use ρ vs. μ and ρ vs. e events to obtain B_{μ}/B_{e}. This would likely reduce the systematic error due to more complete cancellation of various contributions, especially from trigger effects. We estimate our efficiency for such decay modes to be 15%, giving us some 4 × 10^{5} of each type. The statistical uncertainty on B_{μ}/B_{e} in that case would be 0.22%. With either of these analyses, lepton universality, therefore, can be tested with much improved precision over current measurements.

Tau decays to π and K

The partial widths for the two pseudo-scalar decays, τ^{−} → π^{−}ν_{τ} and τ^{−} → K^{−}ν_{τ}, can be calculated using the experimental measurements of the π and K decay constants, f_{π} and f_{K}. The ratio of the two partial widths is related by the Cabibbo angle θ_{C} with a phase space correction factor [60, 62],

\frac{B_{K}}{B_{π}} = \frac{Γ_{K}}{Γ_{π}} = tan^{2} θ_{C} \frac{f_{K}^{2}(1 − m_{K}^{2}/m_{τ}^{2})^{2}}{f_{π}^{2} (1 − m_{π}^{2}/m_{τ}^{2})^{2}} .

Therefore, a precise measurement of the ratio of the branching ratios allows an accurate determination of the Cabibbo angle in τ decay with high momentum transfer. At present, the ratio is known only to ∼ 30%[42].

The CLEO II detector is not well suited for the measurement of this ratio. An accurate measurement of B_{K}/B_{π} requires good π/K identification at high momenta since the momentum spectrum of the K or π is relatively flat up to 5 GeV/c and B_{π} is about 17× B_{K}. A preliminary study of B(τ^{−} → K^{−}ν_{τ}) using one-prong tau decays (1 vs. K) finds only ∼ 43 events/fb^{−1} yielding a detection efficiency of 0.45% in CLEO II. The detection efficiency is reduced greatly because a momentum cut of p_{K} < 0.7 GeV/c is required to obtain a relatively background free sample of τ → Kν decays. Increasing the momentum acceptance of the kaon to 3 GeV/c in CLEO III would increase the efficiency by roughly an order of magnitude.

For a data sample of 3×10^{7} τ^{+}τ^{−} and improved particle identification at high momentum we expect to identify 1.2×10^{4} K candidates. Based on our preliminary analysis, the expected background level in the sample is ∼ 1%. Assuming a similar detection efficiency for τ^{−} → π^{−}ν_{τ}, we expect 2 × 10^{5} events. Thus, B_{K}/B_{π} could be measured with a statistical precision of ∼ 1%, a dramatic improvement over the current world average measurement.

Tau decays to ρ and K^{∗}

The decays τ^{−} → ρ^{−}ν_{τ} and τ^{−} → K^{∗−}ν_{τ} involve the coupling of the weak vector current to ρ(770) and K^{∗}(890), respectively. Measurements of the branching ratios allow studies of the hadronic weak current. B_{ρ}/B_{e} = 1.32 ± 0.05. The Cabibbo-suppressed K^{∗} decay is related[60] to the Cabibbo-favored ρ decay by

B_{K^{∗}}/B_{ρ} = tan^{2} θ_{C} · f · \frac{g_{K^{∗}}^{2}}{g_{ρ}^{2}} .

where the factor f corrects for the differences in the available phase space,

f = \frac{m_{ρ}^{2}}{m_{K^{∗}}^{2}} \frac{(1 − m_{K^{∗}}^{2}/m_{τ}^{2})^{2}(1 + 2m_{K^{∗}}^{2}/m_{τ}^{2})}{(1 − m_{ρ}^{2} /m_{τ}^{2})^{2}(1 + 2m_{ρ}^{2} /m_{τ}^{2})} ,

g_{ρ} and g_{K^{∗}} are the coupling strengths of the ρ and K^{∗} to the vector current. The relationship between the two couplings depends on whether the SU(3) symmetry is exact or broken. If SU(3) were exact, then g_{K^{∗}} = g_{ρ} giving

B_{K^{∗}}/B_{ρ} = 0.038 .

On the other hand, for broken SU(3) the Das-Mathur-Okubo[64] sum rules give g_{K^{∗}}/m_{K^{∗}} = g_{ρ}/m_{ρ} and the prediction becomes

B_{K^{∗}}/B_{ρ} = 0.052 .

The current world average measurement of the ratio is 0.061 ± 0.009, which excludes the exact SU(3) symmetry at ∼ 2.5σ level[42].

For the measurement of the ratio, B_{K^{∗}}/B_{ρ}, the ρ and K^{∗} candidates can be selected using events where one τ decays via τ → eνν or τ → μνν. For a data sample of 3×10^{7} τ^{+}τ^{−}, we expect to collect 6×10^{5} ρ candidates for a detection efficiency of 15%. Traditionally, the K^{∗} decay is identified with a detached vertex through the decay chain, K^{∗−} → π^{−}K_{s}^{0} → π^{−}π^{+}π^{−}. Including the branching ratio for the decay chain, the detection efficiency is expected to be ∼5%. This yields 1.7 × 10^{4} K^{∗} candidates, corresponding to a statistical precision of 0.8% on the ratio of branching ratios. The dominant background in the ρ sample is the 3% feed down from the decay τ^{−} → π^{−}2π^{0}ν_{τ}. In the K^{∗} sample, we expect a contamination of 5% from the three-charged-pion decay. The background from the misidentification of π as a lepton tag has equal contributions to both samples. We expect an overall systematic error of 0.4% from the uncertainty in the background subtraction. Since the topology of the K^{∗} decay is quite different from that of the ρ, the systematic errors in identification efficiencies do not cancel in the ratio. We expect a systematic uncertainty of 1.6% on the ratio of these efficiencies. This yields an overall precision on the ratio of 1.8%, allowing a sensitive test of the Das-Mathur-Okubo sum rules.

An independent method of measuring the ratio is to use the other decay channel K^{∗−} → K^{−}π^{0}. A preliminary analysis using CLEO II data where the K^{−} is required to have p_{K} < 0.7 GeV/c yields 18 τ → K^{−}π^{0}ν candidates per fb^{−1} with about 20% background. The overall detection efficiency for this process is only ∼0.23% and is dominated by the limited momentum range of particle identification. Extending the π/K identification up to 3 GeV/c would increase the detection efficiency for this mode by a factor of 15. A similar gain in the purity of the sample is expected with improved particle identification. Thus, for a 30 fb^{−1} data set, we expect 4500 events providing a measurement of B_{K^{∗}}/B_{ρ} with a statistical precision of 1.5% and with much lower systematics than the previous method.

Tau decays with η mesons

CLEO has recently observed the first example of τ decays with an η meson in the final state, τ → π^{−}ηπ^{0}ν_{τ}[65]. The branching ratio for the decay has been calculated using the measured cross section for e^{+}e^{−} → ηπ^{+}π^{−} together with the CVC hypothesis. Using the cross sections measured at Novosibirsk and DCI, Gilman predicts[66] the branching ratio to be B_{πηπ^{0}} = 0.15%. This is in good agreement with the CLEO II measurement of (0.17 ± 0.02 ± 0.02)%.

There is no firm theoretical prediction on the branching ratio for the decay τ^{−} → π^{−}η2π^{0}ν_{τ}. However, the branching ratio is related to that for the decay τ → π^{−}ηπ^{+}π^{−}ν_{τ} by isospin invariance: B_{πη2π^{0}} ≤ B_{3πη}.

The decay τ^{−} → K^{−}ην_{τ} is allowed in the Standard Model but is suppressed. The branching ratio has been estimated using a Chiral Effective Lagrangian by A. Pich[67],

B_{Kη} ∼ 1.2 × 10^{−4} .

There are other decay modes such as τ^{−} → π^{−}2ην_{τ} and τ^{−} → π^{−}2ηπ^{0}ν_{τ} which are predicted to have very small branching ratios.

To study the above decays requires large data samples and these can only be observed and studied at a facility with CESR Phase 3 levels of luminosity and a detector comparable to CLEO III.

Second-class-current decay

The decay τ^{−} → π^{−}ην_{τ} is of particular interest in the Standard Model of electroweak interactions as it is an example of a decay via a second class current. Weak currents are classified according to their G parity as shown in Table 9.

Table 9: Classification of weak currents according to their G parity

Current Class Vector Axial Vector
First  G = +1  G = −1
Second  G = −1  G = +1

Since the final state contains two pseudoscalars, the decay must proceed by a vector current, but the G parity of the πη is G = −1, hence this proceeds through a second class current. Second class currents are strongly suppressed due to isospin symmetry in the Standard Model and a branching ratio of order 10^{−5} is expected. Observation of a sizable branching fraction could indicate the existence of second class currents or some other new interaction. The current upper limit on this branching ratio is 3.4 × 10^{−4} at the 95% confidence level[65] and could approach 3 × 10^{−5} for CLEO II, but will likely be limited by background from τ → Kην_{τ} given the poor CLEO II π/K separation.

The simplicity of the decay process τ^{−} → π^{−}ην_{τ} → π^{−}γγν_{τ} provides a clean laboratory for the search[68]. The candidates for the decay can be selected with a lepton tag. A π tag with no accompanying photon may also be possible. The η meson is required to decay into the γγ final state. The other η decay channel, η → π^{+}π^{−}π^{0}, is not usable because of the large background from the unsuppressed decay τ^{−} → π^{−}π^{+}π^{−}π^{0}ν_{τ}. The dominant backgrounds are expected to be due to the following decay modes:

• τ^{−} → π^{−}π^{0}ν_{τ} , where one of the photons from the π^{0} decay is lost and the other photon is combined with initial state radiation or a fake photon created by the crystal clustering algorithm due to shower fluctuation or noise.

• τ^{−} → π^{−}2π^{0}ν_{τ} , where one photon from each π^{0} decay is lost because the photon is very soft or merged with another photon.

• τ^{−} → Kη̄ν_{τ} , with the K misidentified as a π.

A special clustering algorithm optimized to reduce fake or merged photons may be required. The ability to detect (or veto) low energy photons will be crucial for suppressing these backgrounds. The last background should be manageable with the improved π/K separation of CLEO III. We expect a sensitivity to the branching ratio of order 10^{−5} for a data sample of 3 × 10^{7} τ^{+}τ^{−}.

Forbidden decays

The search for decays forbidden by the Standard Model has long been a tradition of particle physics. In recent years the immense interest in models that go beyond the Standard Model, such as compositeness, technicolor, lepto-quark, and new horizontal gauge bosons, has intensified interest in these kinds of searches[69].

The traditional hunting ground, such as μ or K decay, has a limited number of decay channels because of the limited phase space. The τ lepton is an ideal laboratory for the search of the decay of a third generation lepton into first or second generation leptons or hadrons. The large lepton mass allows the decay into a number of purely leptonic final states such as τ^{−} → e^{−}e^{+}e^{−} and τ^{−} → μ^{−}e^{+}e^{−} or semileptonic final states such as τ^{−} → e^{−}K^{0} and τ^{−} → μ^{−}ρ^{0}. In several interesting models the rate for a lepton number violating decay involving a τ is many orders of magnitude larger than a similar decay involving a muon due to the large mass of the τ. For example, in a superstring model τ decays are enhanced by (m_{τ}/m_{μ})^{5} over muon decays leading to the prediction that B(τ → γμ) ∼ 2 × 10^{5}B(μ → γe). In a model with a charged Higgs, τ decays are enhanced by (m_{τ}/m_{e})^{2} leading to the prediction that B(τ → μμμ) ∼ 2 × 10^{7}B(μ → eee).

To date the most sensitive search for lepton number violating τ decays has been carried out by the CLEO II experiment. Here a 95% confidence limit of 4 × 10^{−6} has been set for the decay τ → γμ[70]. Other experiments have set limits on similar decays which are in the range of 10^{−3} to 10^{−5}[42]. Since almost all of these searches have been limited by statistics (e.g., the CLEO II result on τ → γμ), we expect to reach the sensitivity of 10^{−7} for most channels with a sample of 30 fb^{−1}.

2.3.3 Structure of the τ − ν_{τ} − W vertex

The large data sample expected permits a detailed investigation of the structure of the τ − ν_{τ} − W vertex. The Michel parameter ρ has been measured with a precision of 7% using the decay τ^{−} → e^{−}ν_{τ} ν̄_{e} and 8% using the decay τ^{−} → μ^{−}ν_{τ} ν̄_{μ}[71]. This is in sharp contrast to the precision of 0.3% in μ decay[42]. Current τ measurements of this quantity are limited by statistics. A large sample of τ events can be selected using a ρ tag where the momentum of the lepton must be as low as practical for particle identification. We expect to achieve a precision comparable to that obtained in the μ decay. Unlike the μ decay, however, the two leptonic decay modes of the τ permit a test of lepton universality.

The measurement of the ρ parameter consistent with 0.75 could still allow some admixture of V + A, scalar, vector, or tensor interactions. The most general Hamiltonian for the four-fermion point interaction has ten complex coupling constants, corresponding to 19 independent parameters and an arbitrary phase. In contrast to the τ lepton where only the ρ parameter has been measured, the weak couplings of the μ have been completely determined. It is not necessary to measure all 19 observables in order to determine the interaction. Fetscher, Gerber, and Johnson[72] have shown only five measurements are needed to describe the interaction:

• lifetime

• decay asymmetry parameters ξ and δ

• polarization parameter ξ′ • inverse τ decay cross section S for ν_{τ}e^{−} → τ^{−}ν_{e} with ν_{τ} of known helicity.

The lifetime measurement will be discussed in the next section. The decay asymmetry parameters ξ and δ of the daughter lepton with respect to the spin direction of the parent lepton can be measured using the fact that the spin of the two τ′s produced in e^{+}e^{−} annihilation are strongly correlated[60]. The spin of the τ can be determined from the two-body decay modes. For example, if we assume a pure V − A interaction in the decay τ^{−} → π^{−}ν_{τ}, the π^{−} is emitted preferentially parallel to the spin of the τ^{−}. The spin of the τ^{−} can be selected by requiring the momentum of the π^{−} to be in the region corresponding to the π^{−} being produced forward or backward in the center-of-mass of the τ^{−}. The polarization parameter measurement requires the analysis of the polarization of the daughter lepton which is not yet practical at any e^{+}e^{−} collider. The cross section for ν_{τ}e^{−} scattering is even less likely to be measured in the near future since direct ν_{τ} scattering is yet to be observed. Fetscher[73] finds that ρ, ξ, δ, and even η, the “low energy parameter,” can be measured to a statistical precision of a few percent with the data sample expected from CLEO III.

2.3.4 Tau lifetime

The τ lifetime provides a direct measure of the coupling strength of the τ to the charged weak current. In the Standard Model, the decay τ^{−} → e^{−}ν_{τ} ν̄_{e} proceeds in perfect analogy to the μ decay μ^{−} → e^{−}ν̄_{e}ν_{τ}. Assuming μ − τ universality of the weak coupling and that the τ neutrino is massless, the τ lifetime is related to the μ lifetime[60]. Neglecting corrections of O(10^{−4}), we have:

τ_{τ} = (\frac{m_{μ}}{m_{τ}} )^{5} τ_{μ}B_{e} .

The statistical precision of the lifetime measurement for a sample of N events is given by

P = \frac{1}{\sqrt{N}} [A + B (\frac{σ_{v}}{βγcτ} )^{2} + C (\frac{σ_{b}}{βγcτ} )^{2}]^{1/2} ,

where σ_{v} is the error in a single decay length measurement, σ_{b} is the beam spread, and βγcτ (∼ 250μm at \sqrt{s}=10.5 GeV) is the average distance a τ travels before decaying. The constants A, B, and C depend on the event topology. For events with the 1 vs. 3 topology, A = B = C = 1 while 3 vs. 3 events have A = 0.5, B = 2, and C = 0. The first term in the above expression is due to the exponential nature of the decay process. In the second term, with the silicon vertex detector, we expect σ_{v} to be ∼ 150μm. The third term in the above expression takes into account the finite extent of the beams. To minimize this term, we can either select the events in the vertical plane or use events where both τ’s decay into three charged tracks (3 vs. 3). The measurement using 3 vs. 3 events is independent of the beam size (C = 0) as it is the distance between the τ’s that is measured.

To date, the τ lifetime has been measured with a statistical precision of 2% by individual experiments[42]. With the CLEO III τ sample it should be possible to make several independent measurements of the lifetime each with statistical uncertainty below 1%. For example, using events where one τ decays into an electron or muon we expect a precision of ∼0.3% assuming a detection efficiency for this mode of 15%. From a similar analysis using 3 vs. 3 decays we expect a lifetime measurement with precision ∼0.5% assuming a 15% detection efficiency for this mode.

Systematic errors will most likely dominate the measurement of the τ lifetime. The systematic error in this measurement has several components including charged particle tracking, and uncertainty in the size, composition, and lifetime of the non-τ events in the sample. With the large data sample and improved tracking due to the silicon vertex detector, we expect that a systematic error below 1% is feasible.

2.4 Two photon physics

The high luminosity running of CESR Phase 3 and the superb vertexing, particle identification, and calorimetry of the proposed CLEO III detector provide a unique opportunity for the study of two photon reactions. Studies of collisions at very high energies will only be possible at LEP or a linear collider, but the two photon physics program of CLEO will be more than competitive in the medium energy regime. In addition, this program will complement future investigations at low energy electron-positron colliders and studies of low energy proton-antiprotron annihilation.

The current two photon physics program at CLEO may be divided into four main categories, all of which will benefit from the improved luminosity.

• Two photon production of meson-antimeson and baryon-antibaryon pairs.

• Two photon production of mesonic resonances.

• Two photon production of charmonia.

• Two photon production of gluonia (ggg or gggg) or exotic quark bound states (qq̄qq̄).

Over the past two years CLEO has obtained results for the two photon production of proton-antiproton pairs[74]. Measurements of this reaction, when extended to sufficiently high two photon center of mass energies W, provide a test of perturbative QCD. The recent CLEO results extend previous measurements to values of W well beyond 3 GeV. The results show a measured cross section that is larger than the perturbative QCD prediction by approximately a factor of five at low W. The data approach the QCD prediction at large W where the perturbative calculation is expected to yield more accurate predictions. A calculation assuming the excitation of “diquark” intermediate states, on the other hand, is much more successful at predicting the reaction rate at lower W.

More recently CLEO has identified charged pion and kaon pairs at values of W up to 5 GeV[75]. These results also provide a test of QCD, but because of the smaller number of quarks involved and the larger values of W, the prediction is less model dependent. With more luminosity the range of these QCD tests will be extended even further so that the predictions may be challenged more accurately.

The “diquark” calculation also predicts the rates for the production of strange baryon pairs. The large data sets proposed should allow for successful searches for the production of Λ Λ̄, Ξ Ξ̄, and Σ Σ̄ pairs. Measurements of the rates for these reactions should yield considerable information on the “diquark” production mechanism.

The two photon widths of charmonium states provide tests of the charmonium model and, when measured precisely, also test the QCD correction to the two photon widths. In the search for charmonium we have found evidence for the two photon production of the χ_{c2} through its decay into a photon and a J/ψ, and have measured its two photon width[76]. The two photon production of the χ_{c2} has also been detected through its decay into four charged pions. Although these processes have been observed by earlier experiments, the two photon width of the χ_{c2}, nevertheless, remains somewhat controversial as all present studies are severely limited by the available statistics. It will be possible with a much larger data set to settle the question of the two photon width of the χ_{c2}.

Two photon production of the η_{c} has been measured through its decay into K_{S}K^{±}π^{∓}, and K^{+}K^{−} π^{+}π^{−}. The two photon production of the η_{c} will prove to be an important laboratory for the study of this resonance. In particular, the relative branching ratios of the η_{c} will be measured to a precision of about 5%, far better than at present.

The search for gluonium has started by looking for the production of two photon produced all-neutral final states. In this search good use is made of the high resolution CsI electromagnetic calorimeter.

The three tensor resonances f_{2}(1270), a_{2}^{0}(1320), and f′_{2}(1515) are observed at present in both neutral and charged final states. With increased statistics it will be possible to measure the resonance parameters to high precision, but more interestingly, the mixing angle in the tensor nonet will be extracted with a precision of a few percent.

The superb resolution of the calorimeter has enabled the reconstruction of several “all photon” final states including f_{2}(1270) → π^{0}π^{0} and a_{2}(1320) → π^{0}η. With increasing luminosity, the unusual line shape of the a_{0}(980) will be resolved well through the decay a_{0}(980) → π^{0}η. In addition, searches for other scalar states will result in very restrictive upper limits on the two photon widths.

Previous experiments have observed associated production of vector mesons in two photon collisions. The production rates have evidenced many puzzling features which have led to speculation that exotic intermediate states are involved. However, except for the case of ρρ production, all of the previous studies have been limited by the available statistics. CLEO III will be able to measure the production rates and decay spectra very well and set stringent upper limits on the previously unobserved final states ωϕ and ϕϕ.

The greatest challenge for the upgraded CLEO III detector will be maintaining an efficient two photon trigger. With the capabilities of the vertex detector for rejecting cosmic ray, beam-wall and beam-gas backgrounds, we have the promise of a high luminosity source of two photon events to probe QCD processes and study rare or exotic states.

2.5 Υ physics

It is almost certain that CESR will remain the premier laboratory for studies of bb̄ bound states for the next decade. Potential models describing the quarkonium interaction have met enormous success in predicting mass splittings in the heavy Υ system. Results from CESR on the masses and widths of the vector Υ systems and the masses of the orbitally excited χ_{b} states have been able to provide direct verification of such model predictions.

However, the study of Υ spectroscopy is far from complete. Remaining physics objectives include:

1. Measurement of the bb̄ D-state masses.

2. Measurement of the dipion transitions among the χ_{b} states.

3. Measurement of the radiative widths of the χ_{b} and χ′_{b} states with J = 0.

4. Measurement of the h_{b}(^{1}P_{1}) and η_{b}(1^{1}S_{0}) masses.

5. Determination of the partial waves present in the dipion decays of the Υ(2S) and Υ(3S) states.

6. Observation of two-body radiative decays of the Υ, giving information on possible glueball states.

Some of these analyses should yield positive results using resonant data taken before the upgrade to CLEO III. Those that do will almost certainly await confirmation from future data sets. Also, at the higher luminosities expected for the future, we do not exclude the possibility that a larger data sample may provide some as yet unanticipated results. In any case, the future Υ spectroscopy program at CESR promises to refine our understanding of the basic quark-antiquark interaction.

2.6 Conclusions

Important τ, charm and B physics results are guaranteed with 30 fb^{−1} and a significantly upgraded detector. We can make precision measurements of the CKM elements V_{cb}, V_{ub} and V_{cd}. We can measure the form-factors in the semileptonic decays B → D^{∗}ℓ^{−}ν̄_{ℓ} and D → ωℓ^{+}ν over the entire q^{2} range and it may be possible to find B → ργ and thus determine another side of the CP triangle. There is also the exciting possibility that we can see evidence of CP violation.

3 Simulation of CLEO II and CLEO III

The initial comparison of various concepts for the inner detector was accomplished using programs which contain parametric models of the detector performance for single charged tracks. We examined at the resolutions for tracking variables as a function of momentum and angle for each candidate detector and compared them with a similar model of the existing CLEO II detector. These studies gave the quick and reliable feedback necessary to optimize the conceptual design. In addition, this study identified potential problem areas to be addressed in a more complete simulation.

Once the rough design of the detector was specified by the track parameter studies, a full Monte Carlo simulation of the tracking system was performed. The simulation used the CLEO track reconstruction code DUET. It deposited ionization in each device, smeared the ionization, added noise, introduced multiple scattering from all the material traversed including the gas and discrete wires, and removed hits according to an efficiency table. The tracking code was then asked to do the pattern recognition to find the appropriate hits on each track and then reconstruct the track parameters by fitting those hits. This code was used to measure the single track parameters and to measure the resolution and reconstruction efficiency for physics quantities of simulated B decays. Using this code we were able to reproduce the present CLEO II track parameter and mass resolutions and compare them with the CLEO III model.

3.1 Parametric model studies

The CLEO III upgrade is primarily driven by the desire to study rare B decays. Operationally, this means providing space for particle identification by reducing the tracking length relative to CLEO II while maintaining the tracking performance by simultaneously reducing the amount of material. The material in the various devices of CLEO II which a charged particle under normal incidence traverses, is listed in Table 10. The CLEO III design was arrived at by studying the effects of changes to the inner detector geometry and material on the track parameters: dp/p, σ_{tan θ}, σ_{IP}, σ_{z}, and σ_{ϕ}. All designs were compared with a similar model for CLEO II. In these studies, the various tracking layers were defined by their radii and intrinsic spatial resolution, multiple scattering was added at discrete radii, and the track parameter resolutions for a given momentum and angle were derived by a matrix inversion technique. The designs studied were constructed out of modular pieces consistent with the new IR. The cell design for the tracking chambers was based on that used in the CLEO II drift chamber. The modular pieces for the parametric model were chosen among the specifications shown in Table 11.

The track parameter studies showed that only two tracking devices are necessary, an inner silicon detector followed by a large, uninterrupted, low mass drift chamber. Both the silicon and the inner section of the drift chamber will follow the stepped structure of the

Table 10: Material in the CLEO II inner detector at normal incidence.

Subsystem  Material (% X_{0})
Beam Pipe  0.44
PTL  0.40
VD  0.94

DR

- inner wall/cathode  0.77
- gas  0.43
- wires  0.32
- endplates and electronics  60–100

Silicon^{†}

- detectors  0.4/layer
- outer support  0.10

† This three layer, double-sided detector will replace the PTL in 1994.

Table 11: Definitions of devices used in the track parameter study.

Subsystem  rϕ resolution z resolution  Material
Si  15 μm  30 μm  0.40%/layer
Tube Detector  50 μm  —  0.06%/layer
VD  100 μm  —  variable
DR wires  150 μm  3–5mm (stereo) variable
DR cathodes  1.8mm

machine around a 300 mrad line to achieve an active detector region beginning at 330 mrad. The minimum achievable drift chamber radius is set to be 17.5 cm by the radius of the accelerator rare-earth, permanent quadrupole. The maximum drift chamber radius is set at 82 cm to allow 20 cm of space for the particle identification system.

The critical parameter for achieving the best momentum resolution in every design was found to be the uninterrupted length of the drift chamber; i.e., with the outer radius determined, the inner edge should extend to the smallest radius without encountering any walls or additional material. In the CLEO III design all walls and/or material between the silicon and the drift chamber have been reduced to the equivalent of one 0.5 mm beryllium layer. Furthermore, a drift chamber gas with twice the radiation length of argon-ethane (X_{0} = 166 m for Ar-ethane) is needed to maintain the tracking resolution equal to that of CLEO II. We note that under these circumstances the addition of any tracking chamber between the silicon and drift chamber degrades the performance of the detector.

The silicon detector begins immediately after the beam pipe which is chosen by background considerations to have a radius of 2.2 cm. The detector has four layers at the approximate radii of 2.5 cm, 3.75 cm, 7.5 cm, and 10.0 cm. The very high resolution of a silicon detector compared to a wire chamber leads to a separated function detector. Namely, the silicon measures the impact parameter in z and polar angle whereas the drift chamber measures curvature. Both devices contribute to measuring the impact parameter in rϕ and the azimuthal angle.

As stated earlier, the mark of a successful design is that it at least reproduce the current CLEO II tracking system performance. In Table 12 we show the five track parameters generated by the parametric studies for the CLEO II detector, the CLEO II/Si detector in which the PTL is replaced with a 3 layer silicon detector, and the proposed CLEO III detector. For every parameter and at all momenta the CLEO III design is comparable to or better than the previous detectors.

The CLEO III design described above has the added benefit that by moving the outer radial edge of the drift chamber to smaller radius, the solid angle of “good barrel” crystals increases from 70% to 77% of 4π. This has the effect of deemphasizing the importance of the endcaps.

We note that increasing the magnetic field above 1.5 Tesla would also improve the tracking as would an improvement in the spatial resolution of the tracking chambers. There is no mechanism, however, for increasing the CLEO magnetic field and there is no obvious technology that would allow a significant improvement in drift chamber spatial resolution technology beyond the present DR 2, so we have not pursued these possibilities. We also note that the CLEO III design is especially sensitive to the placement of the material: the addition of 0.5% X_{0} at r = 30 cm has the effect of deteriorating the tracking by an amount equivalent to placing 2.5% X_{0} at r = 17.5 cm. Care in placement of material is also important for the energy resolution and efficiency of the crystal calorimeter.

Table 12: Results of the track parameter studies at 90^{∘} for CLEO II, CLEO II with the 1994 silicon detector, and CLEO III.

p (GeV/c)
Detector  0.25 0.50 1.0 2.0 5.0

dp/p^{2} (%/GeV/c)

CLEO II  1.10 0.58 0.32 0.20 0.13
CLEO II/Si 1.10 0.58 0.32 0.20 0.13
CLEO III 0.98 0.54 0.31 0.20 0.13

σ_{tan θ} (10^{−3})

CLEO II  6.6 3.7 2.4 1.9 1.8
CLEO II/Si 4.8 3.7 1.7 1.2 1.0
CLEO III 4.7 2.5 1.3 0.7 0.5

σ_{IP} (μm)

CLEO II  166 102 64 42 30
CLEO II/Si 101 57 35 22 14
CLEO III 105 58 34 21 14

σ_{z} (μm)

CLEO II  944 783 733 720 716
CLEO II/Si 116 74 53 43 39
CLEO III 117 67 43 33 30

Table 13: Simulated reconstruction efficiency and mass resolution for 250 MeV/c π^{0}’s generated with | cos θ| < 0.7 where both photons are within | cos θ| < 0.7. The radiation lengths are for 90^{∘} and include all material in front of the CsI crystals. The reconstruction efficiency is relative to that for the calorimeter with no material other than its support container—the absolute reconstruction efficiency in this case is 68%.

Configuration  PID  Total Recon. eff.  σ_{m_{π^{0}}}
(% X_{0}) (% X_{0})  (%)  (MeV/c^{2})
CsI container  —  1.8  100  4.2±0.1
CLEO III w/ aerogel  5.0  13.0  88±2  4.2±0.1
CLEO III w/ RICH  8.0  16.0  85±2  4.3±0.1
CLEO III w/ high pressure tubes  9.0  17.0  84±2  4.2±0.1
CLEO II w/ TOF scintillator  12.4  20.4  81±2  4.5±0.1

3.2 Material effects on the CsI calorimeter

In Table 13 we show the results of a GEANT simulation of the effects of material in front of the barrel crystal calorimeter for CLEO II and for various CLEO III configurations. For this simulation the π^{0}’s were generated uniformly in |cosθ| < 0.7 (θ between 45 and 135^{∘}) with a momentum of 250 MeV/c. Both photons were required to be detected in the good barrel region. The material of the particle identification systems was distributed as might occur in an actual implementation. The radiation lengths are representative and might differ slightly in a constructed detector. The results of this simulation indicate a linear loss of efficiency even for small amounts of material and an indication of loss of π^{0} mass resolution when more than 0.2 X_{0} is traversed. In Fig. 21 we show the π^{0} mass resolution attained in the CLEO II detector when one photon is constrained to the good barrel region and the other passes through varying material radiation thicknesses. A striking degradation of the mass resolution from 5 to 12 MeV/c^{2} can be observed. This work indicates that in the central region the total amount of material in front of the calorimeter should be kept below 20% X_{0} in order to maintain the calorimeter performance relative to CLEO II. It also indicates a need to reduce the amount of material in the barrel/endcap transition region and in front of the endcaps.

3.3 Simulation results

The parametric tracking study described above provided a means for rapidly prototyping various CLEO III tracking configurations and comparing them with the corresponding parametric model of CLEO II. Once an optimum was achieved, the configuration was fully simulated using standard CLEO Monte Carlo and reconstruction codes. The results presented below are for Monte Carlo generated events reconstructed with the CLEO II track

Figure 21: The CLEO II measured mass resolution for π^{0}’s as a function of material traversed by one photon when the second photon is required to be in the good barrel region, | cos θ| < 0.7. The points around 1 X_{0} are in the endcap region and the point around 7 X_{0} occurs at eight discrete ϕ locations at the outer edge of the drift chamber endplates where the support system for the CESR quadrupoles is attached.

finding code, DUET, and with a version modified to include silicon for the CLEO III tracking system. The algorithms used in the track finding code were identical for the CLEO II and CLEO III detectors. This may not be optimum for CLEO III track reconstruction but was used as a basis for comparison. The position of the drift chamber layers for CLEO III were chosen to be identical to CLEO II. The smaller outer radius of the drift chamber was simulated by removing the appropriate number of CLEO II layers. Once again this choice may not be optimum for CLEO III but was used as a reference.

3.3.1 Single track simulations

In order to study the effects of track reconstruction on the various track parameters, ten thousand single charged pion events were generated isotropically over the solid angle for each of several momenta. The performances versus momentum for CLEO II and CLEO III are shown in Figs. 22, 23, and 24 for momentum resolution, z resolution, and tan θ resolution, respectively. These Monte Carlo results reproduce those obtained in CLEO II data and confirm what was indicated in Table 12, namely that the new detector is comparable to CLEO II everywhere except at low momentum where it is noticeably better, primarily since there is less material in the tracking path.

Figure 22: The momentum resolution determined from the Monte Carlo simulation for CLEO II (solid squares) and CLEO III (open squares).

Figure 23: Comparison of z resolution determined from the Monte Carlo simulation for CLEO II (solid squares) and CLEO III (open squares).

Figure 24: Comparison of tan θ resolution determined from the Monte Carlo simulation for CLEO II (solid squares) and CLEO III (open squares).

3.3.2 D^{∗} − D mass difference

As a more stringent test we have studied physics processes involving mass reconstruction using algorithms and code that are presently used in CLEO II. Ten thousand D^{∗} events with the decay D^{∗} → D^{0}π followed by D^{0} → Kπ were generated isotropically in cos θ. In Fig. 25 we show the D^{∗} − D^{0} mass difference, Δm = m_{Kππ} − m_{Kπ}, as a function of D^{∗} momentum. A cut on this mass difference is a common method used to tag D^{0}’s in the CLEO II detector. Over the entire momentum range the new detector is comparable or better than CLEO II. This can be attributed to two effects, the superior momentum resolution and the superior angular resolution of CLEO III. We note that the CLEO III detector reaches σ_{Δm} = 0.7 MeV/c^{2} at p_{D^{∗}} = 2 GeV/c^{2} and by more appropriate use of the silicon this might be reduced to 0.6 MeV/c^{2}.

3.3.3 B decay simulations

Before finalizing the design, more detailed simulations of the detector capabilities were undertaken. In particular, the performance of the detector for specific physics modes was studied in detail. These processes described below reflect some of the important physics goals of CLEO III and were chosen in such a way as to test the performance of the detector for a wide range of physics. For each process, e^{+}e^{−} → BB̄ was simulated with one B forced to decay in a particular fashion while the other decayed according to the B branching ratios. Thus all track overlap, lost track, multiplicity, and combinatoric problems were present.

Presently CLEO II has a large program to reconstruct charged and neutral B decays

Figure 25: The D^{∗} − D^{0} mass difference resolution as a function of D^{∗} momentum determined from the Monte Carlo simulation for CLEO II (solid squares) and CLEO III (open squares).

via the mode D^{∗}π. In this analysis, two variables are plotted: the D^{∗} − D^{0} mass difference and the D^{0} mass. After cutting on these variables, a beam energy constraint is applied to determine the candidate B mass: m_{B} = \sqrt{E_{beam}^{2} − (Σp_{i})^{2}}. In Fig. 26 we show the D^{∗} − D^{0} mass difference for D^{∗}’s from the B decay B^{0} → D^{∗}π. The resolution is comparable in CLEO II and CLEO III. This is also true for the D^{0} mass shown in Fig. 26 where the D^{0} has been reconstructed through its Kπ mode.

As another example, we consider the rare B decay B^{+} → D^{0}K^{+} where the D^{0} then decays to K^{+}K^{−}. This is an interesting decay since it may be used to observe CP violation in a symmetric machine environment. In Fig. 27 we show the reconstructed D^{0} mass and in Fig. 27 we show ΔE, the energy difference between the beam energy and the sum of the D^{0} and K energies. Again, comparable performance for CLEO III relative to CLEO II has been achieved.

As yet another example we consider the rare B decay B^{0} → π^{+}π^{−}. In Fig. 28 we show the reconstructed ππ mass resolution for CLEO II and CLEO III. In Fig. 28 we also show the difference between the beam energy and the total energy of the two pions and the B mass for a constrained fit using the beam energy and the reconstructed π momenta. Again, comparable performance to CLEO II has been maintained.

In Table 14 we present a summary of the results obtained in the physics simulations as shown in Figs. 26-28. Over a wide range of physics modes, the CLEO III detector maintains the excellent performance of CLEO II. As stated earlier, additional gains may be made by using CLEO III specific tracking reconstruction code and by more appropriate placement of the tracking layers. This work in presently underway.

Table 14: A summary of the comparison of tracking simulations for CLEO II and CLEO III for the benchmark processes B → D^{∗}π, B → D^{0}K, and B → π^{+}π^{−}. The resolutions are the standard deviations of fits to the individual plots in Figs. 26-28.

CLEO II CLEO III
Process Measurement resolution resolution
B → D^{∗}π
D^{∗} → D^{0}π m_{D^{∗}} − m_{D} (MeV/c^{2}) 0.48±0.02 0.42±0.02
D^{0} → Kπ m_{Kπ} (MeV/c^{2})  6.31±0.28 6.07±0.25
B → DK
D → K^{+}K^{−} E_{beam} − E_{B} (MeV)  16.1±0.3 17.0±0.3
m_{K^{+}K^{−}} (MeV/c^{2})  5.59±0.13 5.62±0.12
B → π^{+}π^{−}
m_{π^{+}π^{−}} (MeV/c^{2})  20.9±0.4 21.3±0.4
m_{π^{+}π^{−}} constrained  2.40±0.03 2.38±0.03
E_{beam} − E_{B} (MeV)  20.9±0.3 22.1±0.3

3.4 Summary

The physics modes investigated test virtually every aspect of the proposed detector. It is hard to imagine that other processes will require physics capabilities not covered in one of the above modes. Our wealth of experience in using CLEO II for all aspects of b, c, and τ physics together with our detailed simulation of the CLEO III detector give us great confidence that the new detector is optimally matched to our physics goals. a) b)

c) d)

Figure 26: For the decay chain B → D^{∗}π, D^{∗} → D^{0}π, D^{0} → Kπ, the reconstructed D^{0} mass is shown for a) CLEO II and b) CLEO III and the D^{∗} − D^{0} mass difference is shown for c) CLEO II and d) CLEO III. The solid lines are fits to a gaussian and a flat background.

a) b)

c) d)

Figure 27: For the decay chain B → D^{0}K, the reconstructed D^{0} mass is shown for a) CLEO II and b) CLEO III and the difference between the beam energy and the sum of energies for the decay products is shown for c) CLEO II and d) CLEO III. The solid lines are fits to a gaussian and a flat background.

a) b)

c) d)

e) f)

Figure 28: For B → π^{+}π^{−}, the reconstructed ππ mass is shown for a) CLEO II and b) CLEO III, the difference between the beam energy and the total energy of the two π’s is show for c) CLEO II and d) CLEO III, and the ππ mass reconstructed with a beam energy constraint applied is shown for e) CLEO II and f) CLEO III. The solid lines are fits to a gaussian and a flat background.

4 Detector Design

The detector upgrade proposed for CLEO II is divided into charged particle tracking and charged particle identification. The particle identification technology is undetermined at this time although the minimum criterion is 4σ π/K for particles with momentum up to 2.8 GeV/c. The charged particle tracking system must perform a number of functions which include measurement of the charged particle momentum and angles, locating the vertex for charm and tau decays, particle identification using dE/dx, and providing information for the trigger. The tracking system must meet the following criteria under high luminosity conditions :

• It must operate with short bunch spacing, large beam currents, increased background rates, and with quadrupoles within its boundaries.

• It must allow 20 cm for particle identification.

• It must provide track resolutions, mass resolutions, and vertex resolutions which are comparable to or better than those achieved with CLEO II.

The global features of the design are based on the analytical calculations of track resolutions and full Monte Carlo simulations of pattern recognition with background noise, track finding, and track reconstruction for the physics processes discussed in the simulation section. The details of the design draw heavily on the experience of the CLEO collaboration.

A side view of the CLEO III detector is duplicated in Fig. 29. Tracking detectors fill the radial region from just outside the beampipe at r=2.2 cm to r=82 cm. The space between 82 cm and 102 cm is reserved for particle identification. The angular coverage is defined by the IR quadrupoles which restrict the available solid angle to 93% of 4π. The silicon detector consists of four layers at radii of 2.5, 3.75, 7.5, and 10.0 cm; a low mass drift chamber begins at a radius of 17.5 cm and extends to 82 cm with approximately 43 sense wire layers. The magnetic field is 1.5 Tesla.

The distinctive feature of this design is in the separation of the tracking functions between the silicon and the drift chamber. The drift chamber dominates the measurement of momentum, and the silicon dominates the measurement of cot θ and z vertex. Both devices share equally in the measurement of ϕ and the impact parameter. This separation of functions is a natural consequence of extending the radial length of the silicon and reducing the length of the drift chamber to accommodate the IR quadrupoles and particle identification. To compensate for this restriction and ensure no compromise in momentum resolution, the drift chamber inner wall is made of beryllium and its gas must have a radiation length of 332 m, twice that used in CLEO II (50% Ar, 50% C_{2}H_{6}).

In the following section the details of the various devices are described. The options and R&D plans for the various detector sub-systems are also discussed.

Figure 29: A half-section view of the CLEO III detector inside the superconducting solenoid and the CESR Phase 3 interaction region components. The 20 cm clear space around the drift chamber is reserved for the particle identification system—the detector technology is yet to be determined.

4.1 Beampipe

The beampipe for the present CLEO upgrade (to be installeded in September 1994) is designed for the same beam intensity as for the CLEO III upgrade. It consists of a 2 cm radius, double-wall beryllium cylinder with 0.5 mm gap for water coolant, and can maintain a 3^{∘}C water temperature rise for 400 W of heat deposit over its 30 cm length.

The total material amounts to 0.6 % of radiation length, half from 10 μm of gold coating which absorbs synchrotron radiation and the rest approximately equally shared by water and beryllium. The mechanical structure has been analyzed by a finite element program and the buckling pressure was found to be 14 atm of outside pressure for an infinitely long tube. Since the actual structure has both ends fixed to circles, this estimate is a conservative one, and was confirmed by a three dimensional numerical model. Extensive corrosion/erosion and flow tests have been performed with positive results.

For the CLEO III upgrade, the radius of the beampipe will change to 2.2 cm. This will make the structure slightly weaker but still well within the buckling point. Experience from the CLEO II silicon detector operation will determine if it is possible to reduce the gold thickness below 10 μm.

4.2 Silicon

4.2.1 General detector parameters

In this section we describe the silicon vertex detector. The exact parameters given will probably change as the design evolves and we try to optimize the overall design, but at the present time the design provides a working model of sufficient detail to permit us to estimate some performance characteristics.

The silicon tracker consists of four layers of double sided silicon strip detectors starting at r = 2.5 cm, ending at r = 10.00 cm, and filling 93% of the solid angle. The design summarized in Table 15 can be assembled from ∼ 500 identical detector modules, each 2.4 cm × 6.5 cm × 300 μm. In this baseline model we assume the p- and n-side strips are orthogonal, though other stereo angles may be considered. We assume the beampipe outer radius is 2.2 cm with 3 mm clearance to the first layer. The other detectors are placed at radii that are integral multiples of a common base value (1.25 cm in this case) in order to maintain simple scaling relationships among all the layers. With this approach we generate a design in which the full detector can be built entirely from identical devices, thereby reducing costs and simplifying the testing and R&D. This simplicity will also propagate outward to the external electronics and the mechanical structure.

Figures 30 and 31 show the rϕ and rz views of the proposed detector. In the rϕ view one sees that the detectors overlap each other by approximately 2 mm in order to facilitate

Table 15: Arrangement of layers and detectors

Layer Radius (cm) # Detectors in ϕ # Detectors in z
1  2.50  8  2
2  3.75  12  3
3  7.50  24  6
4  10.00  32  8

Figure 30: Transverse section through proposed silicon detector.

alignment; the rz view illustrates the lengthwise daisy chaining of detectors. The left and right ends of the detector are read out separately and the detectors are not joined in the middle. In layer 2, where an odd number of detectors (3) span the length, one readout handles one detector, and the other readout handles two detectors.

Readout of the z-side information is an important issue. This design is based on the double layer design used in the CLEO II silicon vertex detector, in which z strips, perpendicular to the ϕ strips, are read out via longitudinal traces on top of an insulating layer of SiO_{2}. The CLEO II detector design is illustrated in Fig. 32. Our final design will differ from the CLEO II design in many important respects, but some form of the double layer readout structure will have to be incorporated into the CLEO III detectors, so the CLEO II design provides a convenient starting point. For the time being we assume the z-side is the n-side, as in CLEO II, but plan to examine this choice more carefully at a later date.

Figure 31: Longitudinal section through proposed silicon detector. The length scale is compressed a factor of two relative to the radius scale.

Figure 33 shows the layout of traces and strips on the two sides of the detector. Note that all longitudinal traces and strips have a pitch of 44 μm, and on the n-side, the sense strips have a pitch of 124 μm. These choices are the result of the following considerations:

1. With the detector dimensions fixed by simple geometric considerations given above, the rϕ pitch is then chosen such that the number of readout channels is 512, to be compatible with using, for instance, four 128-channel readout chips.

2. Strips on the z-side must be mapped 1:1 onto readout traces, without ganging strips into one readout channel. Ganging leads to increased noise and in some ways complicates pattern recognition. The daisy chaining of multiple detectors already imposes an unavoidable ganging which must not be increased. The rather large pitch on the z-side, 124 μm is a result of this restriction. In order to keep all detectors identical, the pitch is constant.

3. No floating strips may be used on the z-side. While it is common practice to employ floating strips in a detector to reduce the strip pitch without increasing the number of readout channels, it is not possible to use such strips on the z-side. Capacitance to the overhead, orthogonal readout traces would trap some fraction of the signal charge, dispersing it across all channels and thereby effectively losing it. At a later point in this document we will discuss some options for reducing this capacitance, but it is unlikely to be small enough to overcome this restriction. On the rϕ-side we also dispense with floating strips because with the small (44 μm) pitch they are not necessary.

4. Granularity is not an issue. Estimated backgrounds yield layer 1 occupancies of 0.05% per strip for the expected beam currents, the proposed shielding, a 1 μs integration time in the preamplifiers, and the above choice of pitch. This is negligible and, therefore, plays no role in determining strip pitch.

Figure 32: Schematic of the CLEO II silicon detectors, including blow up of the double layer structure showing sense strips and longitudinal readout traces with contact vias.

Figure 33: Schematic of proposed detector strip layout.

We note that 44 μm is not a convenient match to existing readout chips, and are exploring alternative designs that permit use of a standard 50 μm input pitch.

4.2.2 Resolution

Resolution is intimately connected with choice of pitch, the signal to noise ratio (S/N), and the algorithm used for clustering pulse heights. It depends, in addition, on the angle of the incident track as this determines both the number of strips crossed by the track and the mean pulse height on each strip. We have simulated energy deposition and carrier transport and diffusion in a silicon detector in a stand-alone Monte Carlo. Cluster position is the pulse height weighted mean of the first and last strips in a cluster, with a minimum cut on pulse height required. Figure 34 shows the resulting resolution for the z and rϕ sides of the detector, respectively, as a function of the angle of incidence of the track, for three values of S/N = 10, 15, and 20. We see that for the higher values of S/N the resolution is dominated mainly by the choice of strip pitch and by Landau fluctuations with a less strong dependence on S/N. For lower values of S/N, however, the resolution of steeply inclined tracks degrades rapidly. For S/N=15, we average over angles and find in z, σ ∼ 24μm, and in rϕ, σ ∼ 10μm.

In light of these results, we take our target for detector and electronics design to be S/N ≥ 15 when all sources of noise and degradations of signal are taken fully into account.

4.2.3 Capacitance

Detector capacitance plays a key role in determining S/N. By “detector capacitance” we refer to the sum of all parallel capacitances that load the preamplifier front-end. These may be categorized as follows:

1. Interstrip capacitance (C_{is}) between sense strips. It can be usefully parametrized[77, 78]

a) b)

Figure 34: a) The calculated z resolution as a function of incidence angle for various S/N ratios.

b) The calculated r − ϕ resolution as a function of incidence angle for various S/N ratios.

by

C_{is} = (0.87 + 1.61\frac{w}{d})L  (14)

where w is the strip width, d the pitch, and L the length in cm. All capacitances discussed here are in pF.

2. Backplane capacitance (C_{back}) between a sense strip and the opposite side of the de-

tector. We note that

C_{back} = ϵ\frac{Ld}{t}

where ϵ = 11.9ϵ_{0} is the dielectric constant of silicon, ϵ_{0} = 8.85 pF/m, L and d are the strip length and pitch in centimeters, and t is the detector thickness in millimeters.

3. Interreadout capacitance (C_{ir}) between longitudinal readout traces on the z-side. This cannot be calculated with Eq. 14 as for interstrip capacitance because some of the field lines terminate on the sense strip electrodes, and thus contribute to the overlap capacitance (see below). In general, it is less than the interstrip capacitance.

4. Overlap capacitance (C_{ov}) between the longitudinal readout traces on the z-side and the electrodes of the z strips themselves. Each overlap contributes a capacitance of order ∼10 fF, but the exact value depends on the widths of the traces, the pitch of the z strips (and for n-type z strips, the width of the gap between the p isolation strips), and the thickness and dielectric constant of the material insulating the readout traces from the sense electrodes. For n:1 ganging of strips, and N readout channels, it is easy to show that the total overlap capacitance loading one channel is given by 2n(N − 1)C_{ov} which can easily be of order 10 pF.

5. Stray capacitance (C_{stray}) is the presence of miscellaneous capacitances due to structures not accounted for in the above list, and we allow typically 1–3 pF for this contingency.

We have estimated the capacitance of the proposed detector units, and in Table 16, compare the proposed detector with the CLEO II (layer 1) detectors. For the CLEO II detectors we have measured both the individual contributions to the total capacitance, and the total capacitance of a single strip. We find the sum of the “exclusive measurements” matches the “inclusive measurement” almost perfectly on the p-side, and there is a 10% shortfall on the n-side. That shortfall is within errors, but it may also represent some small components not taken into account (such as direct capacitance between strip and readout traces).

The key differences are the absence of ganging in the new detector, and a modest reduction in the overlap capacitance, C_{ov}, which we will achieve by narrowing the n-side effective strip width from 36 μm to 18 μm and increasing the SiO_{2} thickness from 3 μm to 5 μm. Since edge effects play a dominant role in this capacitor, we do not anticipate a large impact from these changes, but assume we can reduce C_{ov} by about 25%. Note that the effective strip width for the n-strips is not just the width of the strips, but that plus the region of accumulation layer, out to the p-isolation strip. This directly affects the size of the overlap capacitance. Use of kapton readout circuits, glued directly to the detector, could avoid the need to build the double layer readout, and may slightly reduce the associated capacitances, but appears to be practical only in cases where the aspect ratio of the detector is modest. In our case, where we have an aspect ratio of about 10:1 in the outer layer, such readout may not be so useful. Nevertheless, we intend to follow closely the experience of other groups using or planning to use this technique.

To be safe, we inflate the estimated z-side capacitance per detector to account for any possible missing capacitance. Thus for the z-side, we estimate C_{d} = (\frac{27.5}{24.6}) × 16.7 pF = 18.7 pF. Taking the most difficult case, layer 4, we see that the preamplifiers will have to deal with a load of approximately 36 pF on the p-side and 75 pF on the n-side.

4.2.4 Readout

In Table 17 we summarize performance figures of existing and planned VLSI preamps. Some of the noise figures listed here are estimates or are based on prototypes only.

Based on noise performance alone, two chips look attractive and can be considered candidates at this point. The JAMEX and the VIKING both achieve noise figures of about 150 + 15C_{d} in equivalent noise charge (ENC). Taking these numbers at face value, we conclude that for the layer 4 z-side, the total noise would be ∼150 + 15×75 = 1275 electrons. Absent any other sources of noise, barring shortfall in performance, and ignoring issues of availability and reliability, these chips would just permit S/N∼13.5 (see section on signal size below).

Table 16: Capacitances in the CLEO II and Proposed Detectors.

CLEO II  Proposed
p-side  Detector
Strip Length (cm)  6.0  6.5
Strip Pitch (μm)  112  44
Strip Width (μm)  10  8
C_{is} (pF)  6.1  7.5
C_{back} (pF)  2.4  1.0
C_{stray} (pF)  0.5  0.5
Sum of Individual Capacitances (pF)  9.0  9.0
Observed Total Capacitance (pF)  9.1  −

n-side

Strip Length (cm)  2.12  2.26
Strip Pitch (μm)  105  124
Strip Width (μm)  20+16  8+10
Ganging  3:1  1:1
C_{is} (pF)  3 × 2.7 = 8.1  2.5
C_{back} (pF)  3 × 0.75 = 2.3  1.0
C_{ov} (pF)  6 × 188 × 8fF = 9.0 2 × 511 × 6fF = 6.2
C_{ir} (pF)  3.2  4.0
C_{stray} (pF)  2.0  3.0
Sum of Individual Capacitances (pF)  24.6  16.7
Observed Total Capacitance (pF)  27.5  −

Table 17: Summary of available and planned readout chips.

Chip  Expt.  # ch Pitch  ENC  τ_{RAD}  τ_{int} Power
μm  (kRAD)  μs mW/ch
• CAMEX64 ALPH/CLII  64 100  340+31C_{d}  25  ∼1  ∼2.0
CAMEX128 (none)  128  50  never made
• JAMEX  CLEO II  64 100  150+15C_{d} ∼ 100’s  ∼1  ∼6
AMPLEX  UA2  16 200  300+35C_{d}  NA  0.8  1.1
• VIKING  DELPHI  128  50  135+12C_{d}  50+  2.0  ∼1
• MX3  DELPHI  128  50 350\sqrt{C_{d} + 3.5}  15  ?  .5–1
MX5  OPAL  128  55  325+23C_{d}  5  ?  1.3
MX7  128  hard
KEK-proto KEK  128  50 1400 @20pF  NA  ∼1  NA
KEK-Bfac  KEK  128  50 1000 @20pF  NA  ∼1  NA
• SVX Rev H3 CDF, L3  128  50  300+33C_{d}  1000  ∼1  1.8
SVX-II  D0  128  50  330+57C_{d}  1000 0.132 adj  3.0
ARC  (FNAL)  128? 50?  600+50C_{d}  NA 0.132 adj  1.3

NA = information not available •= CLEO collaborators have experience with these chips

We can expect further noise sources (see section on noise, below), which will degrade this S/N ratio. Moreover, the JAMEX pitch would force 2:1 ganging which will increase the capacitance and the noise; our experience so far with the JFET-CMOS process on which the JAMEX is based is that it is unreliable and yields are low.

Nevertheless, we find it encouraging that existing chips may already meet at least the noise performance requirements. In a later section we will discuss in more detail some explicit detector designs which are less demanding on the electronics than the baseline detector.

Other considerations besides noise will also figure into any decision about preamplifiers. The backend of the preamplifier should match the data acquisition system requirements—the chips listed above differ in the features they offer. To meet the DAQ requirements, it is desirable for the silicon front-end IC to incorporate analog to digital conversion, sparsification, and data buffering, such as are available in the SVX-II. It is not necessary that multiple events be buffered if the conversion time plus the time for data transfer to the off-detector electronics is 10–50 μs.

Since there are several hundred hybrids, depending on design, it is imperative that the number of signals routed to and from them be minimized. To reduce the bulk of the cable plant it is desirable that the data be transferred off-detector serially. Likewise, a single clock should control the data recording sequence and other control information, such as event trigger and configuration should be done serially. To further reduce the cable plant, the serial links should be provided at the hybrid level, not at the IC level. Power distribution should also be at the hybrid level and the number of voltages sent from the outside to the hybrids needs to be limited. Two voltages and a return are ideal—one voltage for the preamp/storage/back-end and one for depletion.

Although backgrounds and, hence, radiation damage are generally expected to be low in CLEO III (see backgrounds section), rad-hard CMOS processes are readily available and provide cheap insurance against the unexpected and the catastrophic. As can be seen from Table 17, such processes provide for hardness at the megarad level. We assume that all on-detector electronics will be rad-hard.

4.2.5 Signal

A minimum ionizing particle traversing 300 μm of silicon deposits charge that is approximately Landau distributed with the peak of the distribution occurring at 78 keV[79]. This corresponds to 21,500 electron-hole pairs generated.

In AC coupled detectors there are non-negligible losses due to finite charge collection efficiency. Signal charge deposited on a strip is shared among all capacitances directly coupled to that strip, including capacitance to the nearest neighboring strips, to more distant neighbors, to the opposite side of the detector, to overhead readout traces, and, most importantly, to the readout electrode. Denoting the last capacitance as C_{c}, the coupling capacitance, we see that the signal collection efficiency is given by

ϵ_{signal} = \frac{C_{c}}{C_{c} + ∑_{i=1}^{n} C_{i}}

where the C_{i} are all the capacitances listed above. Any capacitance on the preamplifier side of the coupling capacitor does not enter the sum, unless it happens also to contact the sense strip, as in the case of capacitance to overhead readout traces. For the CLEO II detector, where we have measured most of these capacitances, we estimate a signal collection efficiency from the above formula to be between 80 and 85%, allowing for uncertainties in the measurements. In practice we observe an efficiency around 80%, or a signal size of about 17,000 electrons.

Clearly any realistic estimation of S/N in an AC coupled design must take into account collection efficiency. While maximizing C_{c} will drive ϵ → 1 in the above formula, technical constraints will limit how far this game can be safely and cost-effectively played.

AC coupling is generally desirable because it provides some immunity from detector dark currents and permits stable pedestals. The integration of the coupling capacitors into the detectors, however, as in the CLEO II detectors, may not be the best choice for the future as it lowers the detector yield, raises the costs, and hampers detector testing. External coupling capacitor chips look promising, and good designs[80] include clamping diodes to protect against catastrophic damage.

DC coupling should not be ruled out. Nearly full signal recovery can be achieved with the large dynamical capacitance typical of charge integrating amplifiers, and that will help in achieving maximal S/N. In addition, the relative simplicity of a DC coupled system (no integrated capacitors, no external capacitor chips) makes it attractive—and probably cheaper than an AC coupled system.

4.2.6 Noise

Noise originates from both the detector and the preamplifier. The detector noise is dominated by three terms:

1. Shot noise. Statistical fluctuations in the arrival of charge carriers in the detector leakage current lead to a noise term contribution

ENC^{2} = 2qI_{leak}|F|^{2}τ

where τ is the peaking time of the amplifier and F is the transfer function to propagate the noise from the source point to the front of the amplifier. This source is typically small, contributing 100–200 electrons of noise, but each detector in a daisy chain adds a new and independent source, so the total noise from leakage current scales as n^{1/2} where n is the number of detectors strung together.

2. Bias resistor noise. Thermal noise in the bias and feedback resistors constitutes a parallel noise source leading to a contribution

ENC^{2} = \frac{4kT}{R_{B}} |F|^{2}τ.

In the CLEO II detectors this is significant only on the n-side where, with three 20MΩ bias resistors in parallel, the net contribution is about 200 electrons. Again, this source also scales with n^{1/2}.

3. Trace resistance noise. Finite resistance, R_{r}, in aluminum readout traces gives rise to

a series noise source with

ENC^{2} = 4kT C_{d}^{2}R_{r}|F|^{2}

for the (relevant) case in which C_{d}R_{r} is small compared to the integration time of the amplifier[81]. The detector capacitance appears in the equation because it provides the path through which the noise current loop is closed. In the CLEO II detectors, this source is negligible on the p-side but may be significant on the n-side where the combination C_{d}\sqrt{R_{r}} is about 4 times larger. In any case, we note that this contribution scales as L^{3/2}, where L is the length of the detector array. For detectors of a given length, this means the noise will scale as n^{3/2} which is a stronger dependence than for the other two terms and needs to be watched.

An important aspect of detector design thus consists of ensuring that the detector noise sources remain small compared to the amplifier noise. The basic ideas are clear from the simple scaling laws illustrated above: keep R_{B} large, R_{r} small, and, as always, minimize the detector capacitance. For daisy chained groups of detectors it is important to take into account the scaling with n^{1/2} and n^{3/2} as noted above. For the purposes of analyzing the detector proposed here, we will assume that we can achieve a detector noise comparable to the 250 electrons we presently observe on the p-side of the CLEO II detectors, and will scale with n as appropriate. We note, however, that the n-side of the CLEO II detectors exhibits more noise than is accounted for by the three sources listed above, and we are still working to understand this additional component. Such n-side effects have also been noted by others[82].

In reality we must deal with a slightly more complex case, namely, with detectors that are multi-strip objects. On the z-side in particular, the cross-coupling between strips is not entirely negligible and multi-strip effects must be taken into account. We have found it useful in the development of the CLEO II detector to model the detector in some detail to understand how noise from various sources propagates through the complicated RC network that describes the multi-strip detector. The detector model is based on a set of detector measurements[83] and standard network analysis techniques, and it reproduces the measured transfer functions of the detectors.

The final noise figures from all sources i, (i = 1...n) are calculated by summing incoherently over sources and integrating over frequency-dependent transfer functions:

ENC^{2} = \frac{1}{π} ∫_{0}^{∞} |F_{C}(ω)|^{2} ∑_{i=1}^{n} (\frac{dI_{RMS}^{2}}{dω} )_{i} |F_{i}(ω)|^{2}dω.

In this formula, F_{C}(ω) is the transfer function of the amplifier, F_{i}(ω) is the transfer function from the i^{th} source to the front of the amplifier, and (dI_{RMS}^{2}/dω)_{i} is the spectral power density for the i^{th} source. This kind of modelling will be used in designing the new detectors to minimize the detector noise contribution.

In general, amplifier noise dominates. It arises from two main sources:

1. White noise. Fluctuations in the bias current of the front-end transistor give rise to a thermal noise current in the channel just as they do in a resistor, and have a spectral noise density given by

\frac{dI_{RMS}^{2}}{dω} = \frac{1}{2π} 4kT \frac{2}{3}g_{m},

where g_{m} is the transconductance of the transistor. Referring this channel noise to the gate gives a voltage that is impressed across the sum of detector capacitance and input (gate-source) capacitance of the front transistor, and consequently the noise:

ENC^{2} = (C_{d} + C_{i})^{2} \frac{4kT}{3πg_{m}} ∫_{0}^{∞} |F_{C}(ω)|^{2}dω,

or ENC = A+B ×C_{d}. The actual noise figure that one can achieve will depend on the transistor design (through g_{m}) and the shaping (through F_{C}(ω)). Table 17 illustrates that, despite widespread effort, achieving a noise slope below 15/pF is not an easy task.

2. 1/f noise. In addition to this thermal noise source, there is also 1/f noise which arises from trapping and detrapping of charge carriers in and near the channel. Such effects can dominate the noise characteristics of CMOS devices where the charge transport occurs primarily at the defect-ridden silicon-SiO_{2} interface; in JFET’s charge transport is deeper in the bulk silicon, and such effects are less important, leading to overall better noise performance. The spectrum (referred to the gate again) is given by:

\frac{dV_{RMS}^{2}}{dω} = \frac{2πF}{W L} \frac{1}{ω} ,

where W and L are the transistor width and length, and F is an empirical factor that depends on the process. The JAMEX amplifier differs from the CAMEX in that JFET’s are used in place of MOSFET’s in the noise-sensitive parts of the amplifier, and this underlies the factor of 2 improvement in the noise performance.

To achieve S/N ≥ 15 requires that we realize ENC ≤ 1200 after summing over both detector and amplifier sources, and assuming the signal collection efficiency is comparable to what we achieve in CLEO II. Improving the signal collection by DC coupling or reducing the noise will improve S/N and help provide some margin of safety.

Figure 35: Longitudinal section of proposed detector, showing split layers and kapton readout circuits.

4.2.7 Design options

We now examine two possible designs for the CLEO III silicon detector.

Design one: Brute Force

This design follows the CLEO II design closely. Detectors are daisy chained in z in each layer, as necessary, with all preamps located outside the tracking volume. Fig. 31 illustrates this design. With the detector parameters as defined in Table 16 we find total capacitance ranges from 9 pF for the p-side of layer 1 to 75 pF for the n-side of layer 4.

The distinctive feature of this design is that it places all the burden on the preamplifier. In Layer 4, with four detectors daisy chained, we can estimate that the detector noise will be approximately \sqrt{4} times the noise of one detector. Amplifier noise must therefore be less than 1090 electrons, which implies the noise slope must be better than 11e/pF, assuming an input capacitance of 25 pF. To judge by the noise slopes shown in Table 17, this is ambitious[84]. The burden on the preamplifiers is therefore substantial, and there is little margin for error. The detectors, however, are relatively simple, and this is the main advantage of this design.

Design two: Modified Brute Force

In this design we avoid some of the large capacitance of the brute force design by using kapton flex circuits to provide low-capacitance, low-mass readout lines. We require the detectors in layers 3 and 4 to be staggered radially, as shown in Fig. 35. Then, at most two detectors are daisy chained together, and the innermost pairs of detectors are read out by kapton circuits, as shown in the figure. A simple and robust means of attaching the kapton circuits to the detectors is illustrated schematically in the inset.

Table 18: Comparison of the two options

Option  Advantage  Disadvantage
(1) Brute  •Simple detectors  •Difficult electronics?
Force  •Cheaper detectors
•Could start detectors today  •No fall back position if
•CLEO II experience is directly electronics slips

applicable

(2) Modified •Same detector advantages as •Mechanically more difficult
Brute Force above •More electronics channels
•Electronics achievable  •Backend electronics needs
•Could start today  development

•High confidence of success

Each 0.001 inch thick kapton circuit has a thickness of 0.008% X_{0}, and copper traces of 6 μm thickness and 12 μm width[85] will add another 0.01% X_{0}, for a total of about 0.02% X_{0}, or about 1/15 of a silicon detector thickness. The interstrip capacitance on the kapton circuits is estimated for a 13 cm length of kapton circuit with strip pitch 44 μm and strip width 12 μm to add about 6 pF. In layer four, therefore, the z-side will see a worst case capacitance of 2 × 19 + 6 = 44 pF. This is less than the capacitance in the layer 3 z-side of the CLEO II detector.

This design has the advantage of simple detectors, as in the previous “brute force” design, but the demands on the preamplifier are much reduced. A noise slope of 18/pF would yield a total amplifier noise of approximately 1150 electrons, and with the detector contribution the total would be under 1200. The design is also robust in allowing a fall-back position in which no detectors are daisy chained, but every detector is connected either directly, or through a kapton circuit, to a preamplifier. This increases the number of channels, but reduces the capacitance per channel and eliminates all ganging.

The principal advantage of this design is that it breaks little new ground and could be started today with high confidence of success.

Comparison of the two options

In Table 18 below we compare the advantages and disadvantages of the options proposed above.

4.2.8 R&D

We list below some of the areas where R&D should begin.

• We will study the capacitance of a large variety of 3-dimensional structures, especially crossed strips, with different geometries and dielectric media. Experimental measurements and development of code to permit calculation of capacitance between arbitrary structures in 3D will be done. Such expertise will form part of the technical underpinning of the detector design.

• We will study the properties of kapton circuits in close detail, to understand the capacitance, trace conductivities, and dielectric noise.

• High resistance p-side and n-side biasing will be studied. Polysilicon will not suffice. We will pursue an early design that we tested but did not retain for CLEO II, in which MOSFET’s are operated close to pinch-off. For such operation, the variation in the threshold voltage from one transistor to the next, across the detector, could be important and will be studied. Other alternatives include moving the bias resistor and AC coupling capacitor off the detector.

• Radiation damage in detectors, particularly if MOSFET’s are used as described above, will be tested. We will use the CHESS facility for such investigations, as we have done previously in assessing the radiation hardness of the CAMEX. Radiation damage in readout electronics will also be evaluated.

• We will investigate available IC bonding technologies to determine their applicability to our assembly. The efficiency of the bonding procedure is important and vendors will be tested with prototype production.

• Noise studies in the CLEO II detectors will be brought to a conclusion. The remaining n-side noise sources will be identified.

• Full scale detector prototypes will be constructed within the first year since our experience with CLEO II was that the full scale detectors with double layer readout exhibited features not present in small test structures. We are presently working with a variety of vendors to come up with designs they can mass produce cost-effectively.

• Engineering evaluation of the front-end readout electronics requirements, especially the noise performance and power dissipation requirements is critical and available chips will be tested.

4.3 Drift chamber

The present CLEO II tracking system, including the Precision Tracking Layer detector (PTL), Vertex Detector (VD), and Drift Chamber (DR 2), has shown that it can perform under high luminosity conditions. The addition of a silicon detector will enhance the system. Therefore, we use the attained CLEO II tracking performance as a benchmark for CLEO III. This includes excellent momentum resolution, dp/p < 0.56% at p = 1 GeV/c, excellent angular resolution, σ_{tan θ} ∼ 0.004, and good impact parameter resolution, σ_{IP} < 75 μm (< 35 μm when the silicon is added). These tracking parameters translate directly into the resolutions of physics variables. For example, the D^{∗}-D^{0} mass difference resolution is presently ∼1 MeV which provides a 40:1 signal to background for D mesons tagged by this method. The present CLEO II tracking system performance has been achieved by using a 1.5 Tesla magnetic field, ∼0.8 m long tracking length with minimal material, and good spatial resolution—50 μm in the PTL, 100 μm in the VD, and 150 μm in DR 2. The goal of the CLEO III drift chamber is to attain comparable or better tracking performance while allowing 20 cm of space for a particle identification system and space for the high luminosity interaction region quadrupoles.

4.3.1 Mechanical overview

The momentum resolution of a tracking system depends primarily on the uninterrupted tracking length, the amount of scattering material, the number of measurements, and the magnetic field. For good momentum resolution, the drift chamber must fill the available space as completely as possible. It must also allow room for the insertion of the inner detector and for a 20 cm particle identification device at the outer radius. The inner radius is limited to 17.5 cm by the support system of the silicon vertex detector and the assembly procedure for the interaction quadrupoles. This places the first sense wire layer at 18.8 cm. The physical outer radius has been limited to 82 cm, 12 cm less than the current CLEO outer drift chamber[88], providing 20 cm of radial space for particle identification in front of the CsI crystal holder. After allowing space for an outer support skin and segmented cathode readout on the outermost layer, the available active tracking length is ∼62 cm with the final sense layer radius at 79 cm. The longitudinal dimension of the detector can be the same as that of the current drift chamber, 94.0 cm inside the endplates, and is limited by the existing CLEO magnet, the endcap calorimeters, and the endcap particle identification systems. Additional space for the particle identification device in the endcap region will be provided by freeing longitudinal space currently used for the interaction quadrupole support. This has the immediate advantage that the obstructions of the barrel crystal calorimeter due to the endplates and, more importantly, the stiffening ring at the outer radius of the endplates, are moved to a larger | cos θ|.

To maximize the polar angle coverage while providing access to the mounting points for the vertex detector and accommodating the intrusion of the interaction quadrupole, we have chosen a multi-step endplate design as shown in Fig. 36. Such a design allows interleaving of the IR quadrupoles and the drift chamber endplate to obtain maximal detector solid angle. The positions of these steps will depend on the final IR quadrupole design and the exact cell size and shape employed. These issues are presently under study in our R&D program.

Figure 36: Drift chamber structural design.

Using a similar cell design and wire holding scheme to the present CLEO chamber, the polar angle can be extended to | cos θ| = 0.93 with 12 layers in the stepped sections followed by a section of flat endplate with 31 layers, providing a total of 43 sense wire layers.

Wire tension will be held at the outer radius by composite skin panels with an aluminum thickness of 1.6 mm as in the current CLEO drift chamber. We have found that these panels (total thickness of 2% X_{0}) do not degrade the electromagnetic calorimeter in back of them, and we have tested the strength of this system at 10 times the required load. An inner beryllium tube, 0.5 mm thick (0.15% X_{0}), will be used as an inner gas seal and may hold some of the wire tension. It also provides mechanical protection and RF shielding with the minimum material between the inner detector and the drift chamber measurement volume.

Reducing and redistributing the material in the drift chamber is also important for maintaining or improving the performance of the CsI calorimeter and may have substantial impact on the particle identification system. To date, most CLEO II analyses only use the “good barrel” region of the calorimeter, | cos θ| < 0.7, for candidate photons because elsewhere the efficiency has been lower and the background higher, due in large part to material in front of the calorimeter. In the proposed detector, the “good barrel” region will extend out to | cos θ| < 0.77, with the endcaps covering 0.77 < | cos θ| < 0.93. This is the result of the main drift chamber having a smaller radius and a different support scheme. In Table 19 we identify the relevant material in CLEO II and indicate a set of reasonable possibilities for CLEO III. Significant reductions in material thickness may be obtained in both the radial and endcap directions by choosing a “thin” chamber gas and a new endplate material, by removing the intermediate tracking chambers and their associated cables, electronics, and

Table 19: Materials affecting particle identification (PID) and CsI calorimeters. The “r” represents material in the radial direction | cos θ| < 0.77; the “e” in the endcap direction 0.77 < | cos θ| < 0.93.

Subsystem  CLEO II (%) CLEO III (%)
Si Detector (r)  1.40  1.70
PT Detector (r)  0.40  –
VD Detector (r)  0.94  –
VD Mounting Hardware(e)  6.90  –
DR Inner Wall (r)  0.77  0.15
DR Gas (r)  0.43  0.18
DR Gas (e)  0.58  0.29
DR Wires (r)  0.32  0.27
DR Wires (e)  0.42  0.42
DR Outer Cathode (r)  0.95  0.95
DR Outer Wall (r)  2.0  2.0
DR Endplate (e)  35.6  20.0
DR Electronics (e)  8.9  4.9
Cooling (e)  11.0  –
DR Field Ground (e)  2.8  1.4
PT,VD,DR Cables (e)  0.5  0.5
IR Quad Support (e)  0.8  –
Total (r) affecting PID  7.21  5.25
Total (e) affecting PID  67.08  27.51
PID (e,r)  12.0  12.0
CsI support (e)  7.0  7.0
Total (r) affecting CsI  19.21  15.25
Total (e) affecting CsI  86.08  44.51

cooling, and by redesigning the DR electronics and cable plant.

The drift chamber materials listed in Table 19 can have a direct, but as yet undetermined, effect on the performance of the particle identification systems that follow. For the CsI calorimeters, the drift chamber material plus that of the particle identification system affect the performance. For CLEO II, this amounts to 19% X_{0}, 12% X_{0} from the time-of-flight and 7% X_{0} from the crystal support structure. For CLEO III, the corresponding number is approximately 17% X_{0}, 10% X_{0} from the particle identification system and 7% X_{0} from the crystal support structure.

4.3.2 Drift chamber endplate

The present DR 2 endplates are manufactured from flat 3.2 cm aluminum plates. With an elastic modulus of 10^{7} psi, they deflect under the wire tension load (4.0 psi) by 0.70 cm, about 20% of the endplate thickness. A thick outer support ring is used to transfer the wire tension load to the outer skin panels and to provide a support point. The outer support ring does not provide a fixed support, as evidenced by the observed deflection in DR 2, and should be reduced in the new design.

The smaller outer radius and larger inner radius of the straight section of the CLEO III drift chamber does not necessarily allow the straight section of endplate to be half of the present thickness. Such a system would deflect by 3.6 cm if simply supported. The stresses would reach about 12 kpsi, 1/3 of the yield stress. Under these conditions it would be difficult to maintain wire position tolerances. Therefore, we are considering three alternatives for the drift chamber endplates:

1. We could use shaped aluminum endplates such as those implemented by the SLD detector. In this design the endplate radius is 1.0 m and the thickness is about 0.6 cm. The shape of the thin endplate is not stable with a simple support at the outer edges and is, therefore, maintained with a thick outer support ring. This ring would significantly degrade the photon efficiency in the region that it shadows near | cos θ|=0.80. Before we implement a similar design, it would be necessary to quantify the loss of photon efficiency in the region shadowed by the outer support ring and the gain in efficiency in the region shadowed by the thinner endplate.

2. We could retain a flat endplate design, but use magnesium which has a radiation length of 14.5 cm and an elastic modulus of 6.5×10^{6} psi. Magnesium alloys can have yield points as high as 38 kpsi. A 2.9 cm thick magnesium endplate would have the equivalent radiation thickness of 0.56 of the present chamber endplate and the deflection under the wire tension load would be 0.86 cm. The maximum stress is about 5 kpsi, 13% of the yield stress. Since this deflection is with simple support at the outer radius, there is no need for an outer support ring other than what is necessary to transfer the load to the outer skin. Magnesium is very reactive but does not appear to pose a problem for the machining required to manufacture an endplate. Reaction of magnesium with chamber gases is presently being studied as part of the R&D program.

3. We could retain the flat endplate design but use a material such as carbon fiber which has a radiation length of 24.0 cm. A 5 cm thick endplate manufactured out of two 0.5 cm carbon fiber skins enclosing a 4 cm foam would have 0.14 of the present chamber endplate number of radiation lengths. This material was used as the inner support tube for the VD. It has a unidirectional elastic modulus of 45×10^{6} psi[89]. With random fiber orientations, the elastic modulus is approximately half of this value. Quasi-isotropic, carbon fiber/epoxy materials are reported to have an elastic modulus of 7 − −15×10^{6} psi. Using structural beams made of woven fiber with half of the fibers oriented along the length of the beam, we have observed an elastic modulus of 6×10^{6} psi[90]. Our measurement using the carbon filament tube used as the inner wall of the current drift chamber is that, for a tube with half of the fibers in the longitudinal direction, the elastic modulus is 2x10^{6} psi. This wide variation indicates that the results are dependent on the details of the fabrication procedure. The strength of a composite endplate ultimately will depend on the strength of the skin material. R&D is proposed to address the suitability and use of this material.

The inner tapered section of the endplate defines and, therefore, lies entirely outside the active solid angle, | cos θ| = 0.93. Each section can then be an aluminum ring of the required thickness to provide wire positioning, mounting points for assembly, and support for the silicon vertex detector. Individual rings are joined by stainless steel hoops. With the DR 2 wire holding scheme and cell size, the stainless steel thickness is limited to less than 1.5 mm which is sufficient to hold the tension between the endplate rings.

Decisions on the above approaches are complicated by the uncertainty in the cell configuration as discussed below. The ratio of the number of field to sense wires directly affects the loads that must be supported by the endplates and the support skins.

4.3.3 Gas, wire choice, and cell design

The current CLEO drift chamber is a small cell design with 14 mm square cells, 3 field wires per sense wire, and half-cell staggering for local ambiguity resolution. The field wires are 110 μm aluminum plated with 1.6 μm of gold which contribute 0.29% X_{0} of material in the current 70 cm of chamber radius. Another 0.03% X_{0} of material comes from the 20 μm tungsten sense wires. We currently use argon-ethane gas with no other additives, which contributes 0.43% X_{0}. We are currently investigating whether to retain the rectangular cell design in the new chamber or to use hexagonal cells, also with 3 field wires per sense wire.

In order to achieve the resolution comparable to or better than the CLEO II detector in the shorter path length proposed, the scattering material in the active region of the drift chamber must be reduced to half of that in the current chamber. We therefore are considering the use of a helium based gas. With argon-ethane in a 1.5 Tesla field the present chamber shows a loss of efficiency at drift distances longer than 6 mm. Care must be made in choosing a light gas for CLEO III so as to address this problem; otherwise smaller drift cells would have to be used, adding scattering material, increasing the force on the endplates, and complicating the design. Slower drift velocity gases can also help if the problem is a question of Lorentz drift angle. Studies of different gas mixtures must be performed to demonstrate that we can achieve, and hopefully improve upon, the current spatial resolution and efficiency using a light gas in a 1.5 Tesla field.

We are considering reducing the material in the wires by using more, but thinner, field wires per sense wire. For example, in a square cell, instead of using three 110 μm field wires per sense wire, one could use seven 55 μm field wires per sense wire. These two designs have nearly the same electric field at the surface of the field wire, however, the tension to field ratio is reduced by a factor of two which may lead to electrostatic instabilities. Because the volume of the field wires in the 7 field wire design is only 58% of the volume in the 3 field wire design, it would appear that the latter design is preferable. Nevertheless, the actual improvement in momentum resolution due to multiple scattering is marginal as it comes at a significant increase in the complexity of building and maintaining the chamber. The key issues are that scattering due to field wires is only one contribution to momentum resolution, a highly non-gaussian one, and that even with 1.0 μm gold plating, as used on the recently obtained wire samples, the scattering in the field wires is due mostly to the gold. The unavoidable contributions to the multiple scattering material, those from the gas and from the sense wires, is 0.21% X_{0} for a 43 layer chamber using the cell size as in the current CLEO drift chamber and a helium based gas with twice the radiation length of argon-ethane. This material creates a spread in the momentum, dp/p, which is nearly gaussian with a width of 0.271%. On the other hand, tracks that hit field wires undergo large scattering and can have their momentum measurement substantially affected.

It is necessary to consider the frequency of such hits and the fraction of the time that the momentum measurement is affected—in the 3 field wire design, only 56% of tracks hit any field wire. It is useful to compare the fraction of tracks with momentum spread less than 0.54%, roughly 2σ for the case of no field wires. In the 55 μm seven field wire design, 87.6% of the tracks and, in the 110 μm three field wire design, 86.8% of the tracks satisfy this criterion. This represents a possible gain of 0.8% in efficiency per track for processes where the detection is dominated by momentum resolution but at the expense of greater mechanical complexity. On the other hand, a larger increase in efficiency can by realized by reducing the thickness of the gold plating. It may be possible to reduce the gold thickness to 0.5 μm. In this case, the efficiency for the 3 field wire design increases to 88.4%, an increase of 1.6% over the same wire pattern with thicker gold plating. Similarly, the efficiency for the case of the 7 field wire design increases to 90.1%, an increase of 1.7% over the efficiency of the 3 field wire design with the same gold plating thickness.

The discussion above favors the use of a design with a large ratio of field to sense wires and a thin gold plating. Possible overriding factors are the ability to construct a chamber with greater than 3 field wires per sense wire, the reliability of aluminum wire much smaller than 110 μm, the availability of high quality, thin gold plated aluminum wire, and the ability to sufficiently define a field cage for a gas in 1.5 Tesla. Additional work is presently underway to finalize the field cage design.

Scattering material in the field wires may be reduced by using a material lighter than aluminum (X_{0}=8.9 cm). Carbon fiber, with a radiation length of 24 cm, would provide a factor of two reduction in the uncoated field wire material which, by extension of the example above, would result in an increase in efficiency of about 1%. The availability of durable, gold-plated carbon fibers, or other materials, with a 110 μm diameter is being considered.

Stereo wire layers are necessary to provide longitudinal measurements in the drift chamber. The number and placement of stereo wire layers not only defines the longitudinal resolution of the detector but can influence the r − ϕ pattern recognition and the spatial resolution as well. The CLEO II drift chamber has 11 single stereo layers separating triplets of axial layers identified in Table 20. In order to save space and provide the optimum r − ϕ pattern recognition, the field cage of the stereo layers is formed by the axial field wires of the adjacent axial layers. With no separate stereo field layers, the stereo sense wires effectively vary in height within the cell resulting in a longitudinal variation of the electric field configuration and, therefore, a longitudinal variation of the time-distance relationship. This variation has been difficult to calibrate, leading to a degradation of the resolution in the stereo wires from about 150 μm to about 220 μm. Alternate wire patterns are actively being studied to minimize or eliminate this problem.

The simplest way to obtain the present performance of the CLEO II drift chamber is to use the same wire pattern for CLEO III. This has the advantage that the performance is well understood. Modifications must be made, however, to account for the stepped inner section. Taking into account the obvious physical differences leads to the design shown in Table 21. In this design the last sense wire is located at 79.0 cm. We believe that such a design is well understood, and with the reduction to the material described above, will provide better performance than the CLEO II tracking system. However, the loss of efficiency at the cell edges seen in CLEO II should be corrected and means for accomplishing this are now under study.

One interesting alternate wire pattern may eliminate the variation of the height of the wire in the field cage without the addition of a large number of extra field defining layers. The configuration is shown in Table 22. There is a natural break in the chamber, between the sections of stepped and flat endplate sections, where the only transition from axial to stereo wires occurs with the addition of one extra field defining layer. This configuration may be able to minimize the cell distortion by adjusting the stereo angles of successive layers. This provides constant hyperbolic sag with dip angle resolution equal to that of the first 43 layers of the DR 2 design. This may have the consequence of an overall improvement in spatial resolution. Such a device is being prototyped to study whether this pattern does indeed enhance the resolution of stereo wires so that such a device may meet the design criteria. A tracking chamber of this type can use the unbroken group of axial layers at low momentum together with the inner silicon detector to provide the pattern recognition for tracks that curl at low radius. At high momentum, each change in sign of the stereo angle will provide an additional longitudinal measurement. One issue being studied is whether this all stereo chamber can attain equivalent pattern recognition performance to an axial/stereo chamber.

Table 20: Present CLEO II tracking wire configuration.

Wire Layer Detector Stereo direction
1 to 6  PTL axial
7 to 16 VD axial
17 to 19  DR 2 axial
20  DR 2 stereo +1.9^{∘}
21 to 23  DR 2 axial
24  DR 2 stereo −2.2^{∘}
25 to 27  DR 2 axial
28  DR 2 stereo +2.4^{∘}
29 to 31  DR 2 axial
32  DR 2 stereo −2.4^{∘}
33 to 35  DR 2 axial
36  DR 2 stereo +2.6^{∘}
37 to 39  DR 2 axial
40  DR 2 stereo −2.9^{∘}
41 to 43  DR 2 axial
44  DR 2 stereo +2.9^{∘}
45 to 47  DR 2 axial
48  DR 2 stereo −3.1^{∘}
49 to 51  DR 2 axial
52  DR 2 stereo +3.3^{∘}
53 to 55  DR 2 axial
56  DR 2 stereo −3.3^{∘}
57 to 61  DR 2 axial
62  DR 2 stereo +3.6^{∘}
63 to 67  DR 2 axial

Table 21: One proposed DR 3 wire configuration based on the CLEO II drift chamber.

Wire Layer Endplate Section Stereo direction
1 to 12  inner(stepped)  axial
13 to 15  outer  axial
16  outer  stereo −2.4^{∘}
17 to 19  outer  axial
20  outer  stereo +2.6^{∘}
21 to 23  outer  axial
24  outer  stereo −2.9^{∘}
25 to 27  outer  axial
28  outer  stereo +2.9^{∘}
29 to 31  outer  axial
32  outer  stereo −3.1^{∘}
33 to 35  outer  axial
36  outer  stereo +3.3^{∘}
37 to 39  outer  axial
40  outer  stereo −3.3^{∘}
41 to 43  outer  axial

Table 22: One alternate wire configuration.

Wire Layer Endplate Section Stereo direction
1 to 12  inner(tapered)  axial
13 to 16  outer  stereo +1.18^{∘}
17 to 20  outer  stereo −1.26^{∘}
21 to 24  outer  stereo +1.33^{∘}
25 to 28  outer  stereo −1.41^{∘}
29 to 32  outer  stereo +1.48^{∘}
33 to 36  outer  stereo −1.54^{∘}
37 to 40  outer  stereo +1.61^{∘}
41 to 43  outer  stereo −1.66^{∘}

Table 23: A survey of various aluminum wire holding schemes. For the failure rate (FR) calculation, one failure was assumed in those cases where none has been observed.

Experiment  Holding  Number Wire Number Years of FR/wire
method  of Al  diam.  of  operation (10^{−5}/yr)
wires  (μm) failures
ALEPH ITC  Crimp/Glue  4,000  147  0  5.0  < 5.0
AMY DR  Solder  25,000  163  3  7.0  1.7
CLEO II VD Taper Pin/Glue 2,880 142&163  0  9.0  < 3.8
CLEO II DR 2  Crimp  36,240  110  46  7.0  18.0
SLD DR  Crimp  33,000  152  0  4.0  < 0.8

4.3.4 Wire holding

Extreme care must be exercised in holding aluminum field wires. Aluminum tends to creep and any wire holding scheme must take care not to damage the wire to the point that it will break immediately or while in service. In principle, if not in practice, crimping, soldering, glueing, and combinations have been used to hold wires. However, the large number of wires in the drift chamber make this choice a critical one.

In Table 23 we provide a survey of the wire holding schemes presently in use and their reported failure rate. One observation is that the DR 2 chamber has the smallest diameter wire and the highest wire breakage rate. Our R&D program includes a serious study of wire holding schemes including the construction of prototypes.

Even though a prototype of the wire holding system can be tested and modified to reduce the number of broken wires, the final device has so many wires compared to any reasonable prototype that the broken wire count, in years of service, cannot be ensured to be zero. For example, the current CLEO drift chamber suffers from an aluminum wire breakage rate of 1 wire per 6000 wire-years. This rate can not be simulated in a prototype. Therefore, it is necessary that any wire holding method adopted for a new chamber allow for the identification and removal of broken wires.

We are presently investigating alternatives to the scheme used in CLEO II. One possible method for holding wires is a modification of that used in the current CLEO drift chamber. Wires are held and positioned by a crimp pin, insulated from the endplate by a bushing as shown in Fig. 37. Since each wire is held by a separate pin/bushing there is no problem with accessibility; broken wires can be removed by removing the associated bushing. The problem with this method comes from the crimp. To maintain a gravitational sag of 60 μm, the 110 μm aluminum wires are tensioned with 170 grams, about 60% of the yield strength. The crimp pins are copper so the wire is deformed at the crimp. Thus, the local stress is increased closer to the yield strength, increasing the broken wire rate. Fluctuations in

Figure 37: Wires are held and positioned by a crimp pin, insulated from the endplate by a bushing.

the dimensions of either the crimp pin, wire, or the crimping tool can also contribute to the variations in the deformation at the crimp and therefore the local stress. There are several modifications to the present wire holding method that could be employed to reduce the possibility of breaking wires. It may be that the breakage is due to variations in the dimensions of the crimp pin so that the amount of deformation of the wire is unpredictable. A quality control program where each crimp pin is measured for both inside and outside diameter could eliminate the out of tolerance pins. This may be impractical because of the difficulty in measuring an inside diameter of 170 μm. Similar pins manufactured out of aluminum would not deform the aluminum wire as much as copper pins. This would reduce the local stress on the wire but may be more susceptible to slipping. This approach has been used in SLD.

Greatly modified bushings and pins (or taper pins) could be used to allow gluing the wires after crimping thereby using both schemes independently. Breakage would then require both schemes to fail. Such a procedure has been successfully used with aluminum wires in the present CLEO intermediate tracking chamber (VD). In constructing this device we found that care must be taken not to let the epoxy run to the edge of the taper pin. If this happens any vibration may cause the wire to break on the sharp edge of the epoxy. In the case of the VD, this problem was identified and solved by putting the epoxy on the back end of the taper pin where it could not reach the wire. The wire breakage during construction was measured to be less than one part per thousand for both the taper pin and the epoxy independently. This type of problem illustrates the care that must be taken for any scheme chosen.

Our R&D program includes a vigorous plan for testing wire holding schemes. We have presently constructed a chamber similar to DR 2 and are attempting to reconstruct the failure rate we have observed over the years. All proposed designs will be tested before implementation. Tests will include some means of reliably accelerating the failure rate to demonstrate improvements over the current wire holding method.

4.3.5 Cathodes

The current CLEO drift chamber has segmented cathode readout on the innermost and outermost anode layers to provide accurate (1.2 mm) measurements in the longitudinal direction. The proposed design does not allow for a cathode readout in the inner layer because the extra scattering material between the last silicon layer and the inner drift chamber layer would degrade the momentum resolution. Furthermore, more precise longitudinal measurements will be available from the silicon vertex detector. Segmented cathode readout, similar to those in the current chamber, may be installed at the outer radius since this is outside the momentum measuring volume. Although inferior to the silicon measurements for determining the longitudinal angle, this will provide a longitudinal position measurement, more precise than can be achieved from stereo wires, for redundancy and for extrapolation into the particle identification system.

4.3.6 R&D

In the previous sections we have identified the areas of R&D necessary to complete the drift chamber project in a timely fashion. All are tightly coupled, mandating careful oversignt of the individual R&D efforts. These we enumerate below:

• A gas of at least twice the radiation length of argon-ethane is required in CLEO III. Measurements of various mixtures will be performed for possible cell designs and must be capable of operating in a 1.5 Tesla field. Drift time, efficiency, and specific ionization measurements will be made in single tubular cells at zero field with the leading candidates then tested in a prototype in the CLEO II magnetic field.

• The optimum ratio of field wires to sense wires will be studied. Test chambers with a small number of square and hexagonal cells with 3 : 1 ratios are being tested in the CLEO II solenoid. Monte Carlo studies need to be continued to better determine the real-world performance gain of alternative field cage designs. We also need to assess the costs of larger ratios in time and money.

• The organization and type of layers—axial or stereo—can have an impact on track finding and hardware trigger algorithms. This has to be evaluated in detail before a final design is chosen.

• The radiation length of the drift chamber endplates, support system, and electronics affects the performance of the adjacent particle identification and calorimeter systems. Alternative materials—aluminum, magnesium, composites —and alternative shapes—flat, bowed—need to be evaluated for structural strength, long term-stability, manufacturability, and physics impact.

• Mounting schemes for the drift chamber will be explored to minimize the impact of this passive structure on the particle identification and calorimeter systems, particularly in the endcap region.

• The CLEO II experience with aluminum field wires indicates that the wire holding scheme needs to be improved. This may be a refinement of the present scheme or a completely new design.

• Low mass materials (relative to Al) for field wires will be surveyed and tested if deemed appropriate.

4.4 Particle identification at high luminosities

A major goal of the physics program is to cleanly measure rare decays to test the Standard Model and be sensitive to possible new physics. This requires excellent charged hadron identification with 4σ π/K separation up to 2.8 GeV/c. All techniques for possibly attaining this quality of separation are based on Cherenkov radiation with two basic approaches utilizing thresholds or measuring the opening angle of the emitted light ring. A constraint on all implementations is that they be no thicker than 12% of a radiation length—the present time-of-flight scintillator thickness—in order not to compromise the performance of the CsI calorimeter.

4.4.1 Ring imaging Cherenkov counters

Ring imaging detectors consist of a radiator to produce Cherenkov photons, a gap to let the Cherenkov cone expand, and a photosensitive device featuring good spatial resolution. Large systems called “RICH” ( or “CRID”) have been constructed for the DELPHI detector at LEP and the SLD detector at SLAC. These detectors utilize gaseous radiators (C_{5}F_{12} or C_{4}F_{10}) or liquid radiators (C_{6}F_{14}) and detect photons using heated TMAE gas. To minimize the number of electronics channels, the photo-electrons are drifted to the end of the detector, as in a TPC, and then perform the multiplication in small proportional chambers. Since the drift time is of the order of 20 μs, these detectors are not suitable for high rate environments. Furthermore, the TMAE chambers and large electric drift fields introduce complexity and large amounts of material (∼20% of a radiation length). On the other hand, there is good evidence that they work[91].

The College de France–Strasbourg group at CERN, led by J. Seguinot and T. Ypsilantis, has achieved considerable progress in the development of a simplified design called “Fast RICH” where both photon coordinates are measured at the detector surface via the pad readout of a multiwire proportional chamber with very fine pad segmentation (approximately 5×7 mm^{2}). Detector signals are processed by front-end digital electronics which have been developed by this group in collaboration with the Rutherford Laboratory[92]. A prototype

Figure 38: The detector system in use at CERN for testing RICH counters.

system has been successfully tested with cosmic rays and in hadron beams at CERN[93]. A sketch of this system is shown in Fig. 38. In this prototype the photon converter is presently TEA gas. A CsI photocathode is being fabricated and will be tested in this detector[94].

For CLEO III, we propose to use liquid radiators (C_{5}F_{12} or C_{6}F_{14}). A sketch of two possible systems, both of which are ∼9% of a radiation length, is shown in Fig. 39. For the barrel region, the proximity focus system is proposed while a spherical mirror focus seems to be best for the endcap. The latter detectors could be a good match to the proposed reconfiguration of the crystals in the endcaps. In Fig. 40 the momentum for 4σ particle separation is shown as a function of incident particle angle. The curve is generated assuming a 2.5 × 2.5 mm^{2} effective photodetector pixel size and an average quantum efficiency of 28%[95]. The minimum separation occurs at the barrel, endcap transition.

The proposed liquid radiator is being prototyped and will be tested at CERN using a CsI photocathode in early 1994. Also in 1994, a new detector with CsI photocathodes will be tested and compared with the previously used solid radiators (LiF and CaF_{2}). This test is crucial since one critical issue for RICH is whether the average CsI quantum efficiency between 6 eV and 7 eV can be maintained at the 15–28% level over a long period of time. Several groups have reported measurements of CsI quantum efficiency with a large spread in the results—approximately a factor of 2. It has been argued that this is due to poor handling of the photocathode material and variations in the conditions of the measurement. For example, the electric field and chamber gas appear to effect the efficiency, but the extent of these effects is controversial. Two measurements are shown in Fig. 41[94, 96]. Initial

Figure 39: Possible RICH configurations with estimates of material thickness: a) proximity focus system and b) system where the mirror acts as a focussing element—in the endcap region this would be a spherical mirror.

Figure 40: The momentum for 4σ π/K separation as a function of the dip angle (Θ_{P}) for a proximity focussing system in the barrel and a spherical focussing system in the endcap.

Figure 41: Measurements of CsI quantum efficiency from Seguinot et al.[94] on the left and Anderson et al.[96] on the right.

tests indicate that the operation of a liquid radiator with good optical purity over a long time not be a problem. The DELPHI experiment[97] has shown that a variety of techniques can be used to achieve this goal. In particular, the particle separation obtainable with 15% efficiency is good. As part of an R&D program, we have started work on techniques for producing and testing CsI photocathodes.

4.4.2 Aerogel

Threshold Cherenkov counters have been used successfully in particle physics experiments for over 30 years. For a particle of mass m and a radiator with an index of refraction n, the momentum threshold for light emission is given by

p_{thres} = \frac{m}{\sqrt{n^{2} − 1}}.  (15)

If the number of protons expected in the data sample can be neglected, a single threshold Cherenkov counter with appropriate choice of the index of refraction is sufficient for π/K separation. In an environment with significant proton yield two consecutive measurements with different Cherenkov thresholds are needed. For CLEO III, π/K separation to 2.8 GeV/c requires n = 1.01 to 1.02, a range achievable by silica aerogels. Several experiments and R&D projects have reported successful use of aerogel for particle identification[98]. All of them, however, reported relatively small numbers of detected photons per unit length of radiator. For CLEO III a minimum of 8 detected photons are needed to provide π/K separation at the 10^{−4} level. Although the 2–3% photon collection efficiency reported by the TASSO[99] and Caltech[100] studies may be sufficient, we plan to study ways to increase this efficiency. We have started an initial effort that uses aerogel with an index of refraction n=1.02 to study light yield, wavelength shifting, and light collection efficiency.

The R&D program on the aerogel for CLEO III addresses three independent aspects of the light collection. The first is a problem of relatively small transmittance of produced light to the photodetector. Most of the Cherenkov photons are produced with short wavelengths in the UV region while the photodetectors used in previous studies were sensitive to photons in the visible range of the spectrum. A test of the light collection efficiency with UV sensitive photodetectors may lead to significant improvement. An alternate approach is to use a wavelength shifter coated photodetector to shift the wavelength of the UV photons into the visible light range and thus improve the effective light collection efficiency.

The second problem is associated with light absorption in the aerogel. The transmission spectra show very low transmittance for photon wavelengths up to about 300 nm and then a rapid rise to about 80% at 600 nm. This large self absorption for UV photons is due to Rayleigh scattering effects since the average path length between the production point of UV photons and an external photodetector is comparable to the absorption length, hence most photons are absorbed before reaching the readout device. On the other hand, the transmittance of the aerogel for the visible light photons is high. Two paths to recover the short wavelength photons are being pursued. One intersperses a thin layer of wavelength shifting compound within the silica aerogel. This can be achieved by sandwiching thin layers of UV transparent plastic (e.g., 50 μm thick acrylic) coated with the wavelength shifters between 2 cm thick sheets of aerogel. In this arrangement the average path of the UV photon before it reaches the wavelength shifter is relatively short, much smaller than its absorption length. A large number of such wavelength shifting compounds with large Stokes shifts are available and there are several laser dyes that can shift UV photons into the 600 nm range. The other path uses plastic fibers doped with wavelength shifter in order to minimize the path of UV photons to the readout system. These fibers are dispersed within the radiator.

The third problem is related to the operation of photodetectors in a magnetic field. The response of photodetectors in a realistic arrangement is being studied. This includes a conceptual design of an aerogel detector that can fit into the available space of CLEO III. Among the candidates are the phototubes used in the CLEO II endcap time-of-flight counters, their variants with UVT windows and UV sensitive photocathodes, avalanche photodiodes (APD), and vacuum APD’s.

4.4.3 Cherenkov correlated timing

Cherenkov correlated timing (CCT) is a new particle identification technique[101] based on using precision timing measurements to determine the Cherenkov angle of photons emitted by particles passing through a quartz radiator[102]. As opposed to conventional imaging designs, the Cherenkov light is not projected onto a large position-sensitive detector surface. Instead, the correlation between the Cherenkov angle and the path-length of totally internally reflected photons is exploited and fast timing is used to deduce this path length.

Figure 42: Principle of Cherenkov correlated timing.

This very simple principle is illustrated by the 2-dimensional example shown in Fig. 42. A charged particle traverses the quartz bar. The track angle θ and the entrance point are known with sufficient accuracy from the central tracking system. If the particle is fast enough, it will radiate photons at the Cherenkov angle χ which is related to the particle’s velocity and the refractive index n of the medium by the simple equation cosχ = 1/(βn). The photons hit the quartz/air boundary at an angle α = θ + χ, and if α is greater than the critical angle sin^{−1}(1/n), the Cherenkov photon will be trapped inside the rod via total-internal reflection. The length of the photon’s path as it bounces toward the end of the rod is a function only of the distance d and the angle α. For the photons shown in Fig. 42, this length is simply given by L = d/ sin α. Since the speed of light in the rod is known, L can be deduced and the Cherenkov angle χ = α − θ can be found by measuring the flight time of the photons to reach the end of the quartz bar.

Fig. 42 also shows the exaggerated difference in flight path for light emitted by pions (dashed) and kaons (solid). Clearly, for the same values of d and θ, the light from a K will arrive later at the photon detector than the light from a π. The arrival time of the Cherenkov light at the photon detector has two components: the flight time of the particle from the interaction point to the radiator bar, and the time it takes for the Cherenkov light to propagate from the creation point to the end of the radiator. Since for a given momentum a K is slower than a π, the two components add constructively resulting in a device whose ultimate resolution is bounded from below by that of a conventional time-of-flight system, and having much better performance in most circumstances.

A CCT detector has several potential advantages:

• It is compact.

• It is fast. The latency time is determined only by the speed of light and the response time of the photon detector. Hence a CCT detector can be used in the first trigger level.

• It is simple. Measuring times accurately does not require complicated position sensitive photon detectors and large projective geometries, but relies rather on a small number of compact phototubes.

Disadvantages are:

• Two centimeters of quartz has a radiation length of 24%, about twice that of the CLEO III design requirement.

• Due to the cylindrical geometry of CLEO III, the separation quality is a function of the angle from the beams and does not meet the 4σ requirement at all angles and momenta.

• A suitable configuration for the endcap region is not obvious.

A detailed Monte Carlo simulation[101] was developed to investigate the feasibility of a CCT detector. The prototype detector used in the simulation is shown in Fig. 43 and consists of individual 2 × 4 × 350 cm^{3} quartz bars surrounding a cylindrical tracking volume with dimensions consistent with the CLEO III detector. Uncertainties in the measured track angles θ and ϕ are taken to be 1 mrad and 0.5 mrad respectively. For the photon detector, a photomultiplier tube is modeled with realistic photocathode efficiency and a single electron transit time jitter of 230 ps (FWHM)[103]. The figure of merit is the number of σ’s of π/K separation and is defined for this device as

n_{σ} = \frac{(t_{K} − t_{π})}{\sqrt{\frac{1}{2}(σ_{K}^{2} + σ_{π}^{2}) + σ_{t}^{2}}}  (16)

where t_{K}, t_{π}, σ_{K}, σ_{π} are the means and standard deviation of the K and π distributions respectively. σ_{t} (= 50 ps) is a constant component representing additional uncertainties in measuring the arrival time, mainly due to the longitudinal size of the beam bunches in the storage ring. A summary of our results is given in Table 24. These results indicate that our simple CCT model detector should deliver good π/K separation for particles with momentum up to 2 GeV/c. For a reduced range of polar angles, n_{σ} > 2 is maintained up to 3 GeV/c.

An R&D effort to experimentally assess the key parameters of the CCT technique and test findings of the Monte Carlo simulation has begun. Several bars of high quality UV transparent quartz have been purchased, along with ultra-fast phototubes and appropriate readout electronics. A scintillator telescope has been constructed to provide rough measurements of the positions and angles of incident cosmic rays, as well as serving as a trigger for the stand-alone data acquisition system. The photon yield and timing characteristics of a variety of geometries and radiator/detector combinations will be investigated. Alternatives to the prototype design discussed above will be explored, including the use of UVT plastic rather than quartz to reduce the thickness of the device, and the possibility of increasing the quantum efficiency of the photon detector by replacing the phototubes by UV-enhanced APD’s[104].

Figure 43: Geometry of a reference CCT detector.

4.4.4 High pressure Cherenkov tubes

The use of pressurized gas Cherenkov radiators is being investigated. To achieve π/K separation in the region from 0.8–2.8 GeV/c with simple threshold counters requires one radiator. A refractive index of n = 1.0154 gives thresholds of 0.79 and 2.8 GeV/c for pions and kaons respectively. Since the photon yield vanishes at the Cherenkov threshold, one might choose to run at a slightly higher index, somewhere between n = 1.0154 and n = 1.02, near the upper limit of easily achievable indices.

The simplest configuration uses thin-walled cylindrical pressure vessels with reflective inner surfaces to direct the light to photodetectors at the ends. Deviation from this simple shape requires additional material and complicates the mechanical design. For the barrel region the detectors would be arranged in a cylinder with the tubes parallel to the beam axis and two layers would be used to insure continuous azimuthal coverage. The 20 cm space allotted to particle identification in CLEO III allows tube diameters up to 10.4 cm and lengths of 2.5 − −3 m. To achieve n = 1.0154 requires a pressure of about ∼230 psi for SF_{6}, a highindex non-explosive gas with good UV transmission down to 150 nm; n = 1.02 is achievable at about 300 psi, beyond which point the room temperature liquefaction point of SF_{6} is

Table 24: Number of σ’s π/K separation vs θ and p for a CCT detector.

θ
0^{o} 5^{o} 10^{o} 15^{o} 20^{o} 25^{o} 30^{o} 35^{o} 40^{o} 45^{o}
1.0 38.3 31.6 26.1 21.6 18.1 15.7 13.5 11.6 10.7 9.6
1.5 15.6 13.0 10.6 9.0 7.3 6.4 5.7 4.7 4.4 3.8
P  2.0 8.5 7.3 5.8 4.8 4.0 3.6 3.2 2.5 2.4 1.9
(GeV/c) 2.5 5.4 4.4 3.8 3.1 2.6 2.1 1.7 1.5 1.5 1.8
3.0 3.9 3.2 2.4 2.2 1.6 1.7 1.2 1.2 1.1 0.9

reached. We are discussing the feasibility of measuring the UV transmission characteristics of pressurized SF_{6} with various companies.

There are several potential advantages to a high pressure Cherenkov detector. These include:

• Intrinsic speed – signal production is instantaneous, transport to detectors defined by the speed of light, and signal development by the response time of the photodetector. This is particularly relevant for low-level triggering.

• Simplicity – of the order of 1000 channels of ADC readout.

Disadvantages are:

• Material thickness

• Safety concerns

• Inhomogeneous placement of material in front of the calorimeter

• Variation of the light collection with polar angle

It appears possible to build a system which contains a minimal amount of material. An infinitely long, 7.5 cm diameter cylindrical aluminum vessel with 0.8 mm wall thickness should hold a pressure of 600 psi. Tests on 1 m long, 7.5 cm diameter aluminum tubing (6061-T6 alloy/temper) of 0.8 mm wall thickness showed some failure in the glue joint of the end plugs at pressures which ranged between 450 psi and 600 psi but no damage to the body of the Al tube. With a 1 mm wall thickness and 10 cm diameter, two layers of tubes would present (at normal incidence) a maximum of 25% and an average of 5% of a radiation length of material. For beryllium the corresponding numbers are 6% maximum and < 2% average radiation lengths. Carbon fiber/epoxy laminate tubes with 2.5 mm wall thickness represent about 25% maximum and 5% average radiation lengths. It should also be possible to reduce the wall thickness by using stronger carbon fibers, at some increase in price. To these estimates, we must add the contribution of SF_{6} which is about 6% of a radiation length at 300 psi.

The tubes are lined with a mirrored surface to route the light to the photodetectors. We are manufacturing aluminized mylar with a magnesium fluoride over-coating and expect the reflectivity between 200 nm and 550 nm to average close to 90%. Kapton and PEEK substrates are also being investigated. For a straight cylinder, there is a decrease in the transmission efficiency between 70 and 110 degrees. A design with a double-sided mirror halfway along the length of the tube slanted at 5 degrees to the tube axis is being modeled. Preliminary Monte Carlo studies indicate that sufficient numbers of photons can be detected if the reflectivities are high enough. In another design the tubes will be lined with a reflecting grating. Another possible variation in light collection would have the light directed to the photodetectors by total internal reflection down a coaxial wave-shifter bar (or rod). Grande and Moss[105] have shown that conversion efficiencies exceeding 90% can be achieved with a mixture of p-terphenyl and bis-MSB in an acrylic film for wavelengths between 200 and 400 nm, along with attenuation lengths of > 1 meter. A more venturesome scheme eliminates the transport of the photons altogether by proposing long photosensitive “strips” either on part of the tube surface or located coaxially within the tube. With the proper optical design, most photons would be incident on the photodetector after only one or two reflections.

For photodetectors we are investigating UV-PMT’s (which must operate in a 1.5 Tesla field), APD’s, and CsI thin-film photocathodes in a parallel-plate gas avalanche chamber. The latter may be applicable to the coaxial photodetector scheme mentioned above. If the photodetectors can resolve single photon pulse heights, it is possible to obtain some particle separation above the K threshold of 2.8 GeV/c. The resolution will depend on uniformity of collection and signal/noise in the photodetector.

Monte Carlo studies have been used to compare the numbers of collected photons per path length of pions and kaons for the “ideal” configuration and n = 1.02 above the kaon threshold. From photon statistics alone, the π/K separation for a 10 cm radiator is approximately 3.8σ, 3.4σ, and 2.6σ in the momentum intervals 2.5–3.0 GeV/c, 3.0–3.5 GeV/c, and 3.5–4.0 GeV/c.

4.4.5 R&D

In addition to the technical details already discussed for each of the techniques, a critical part of selecting the particle identification system for CLEO III will be its ability to do the physics of interest. This requires both R&D and detailed Monte Carlo simulation.

A general Monte Carlo framework is being established for the following:

• Evaluate each technology for detection efficiency and rejection ratio. This includes the performance over all angles and with realistic mechanical structures.

• Simulate the performance in the presence of backgrounds from interactions in nearby material including splash back from the CsI calorimeter.

• Simulate the impact of the particle identification material on the CsI calorimeter performance.

• Evaluate the physics impact of eliminating the endcap detectors and extending the barrel detector from 77% to 83% of 4π.

Below we identify the R&D program for each specific technology. RICH

• Continue the CERN beam tests.

• Investigate the quantum efficiency of CsI photocathodes.

• Optimize the front-end electronics—a digital system has simpler electronics but a larger number of channels, while an analog readout of the pads with interpolation could reduce the channel count dramatically.

Aerogel

• Quantify the photon statistics that can be reliably achieved in a large system. Bulk material and sandwiches of aerogel/wavelength shifter will be studied.

• Develop photodetectors with the requisite light sensitivity and immunity to magnet fields. Candidates are mesh UV-PMT’s and APD’s.

CCT

• Investigate the use of materials with longer radiation lengths than quartz.

• Develop photodetectors with the required timing, noise, and magnetic immunity requirements.

• Develop a device that works in the endcap region.

High pressure tubes

• The viability and safety of a large number of high pressure, low mass containment vessels has to be demonstrated.

• Different pressure vessel materials and thicknesses need to be evaluated to minimize the material.

• While SF_{6} has been discussed, there may be other suitable gases. For SF_{6} and any other gases, determination of the UV transmission properties is essential.

• As discussed above, schemes to make the light collection more uniform over the whole angular acceptance of CLEO III must be developed.

• Evaluate or develop a suitable photodetector—PMT, APD, or CsI deposited cathode.

• Develop a device that works in the endcap region.

Figure 44: A side view of the endcap region of CLEO III showing a possible rearrangement of the endcap crystals to achieve a more projective geometry.

4.5 CsI endcap calorimeter

4.5.1 Mechanical modifications

The CLEO II endcap calorimeter is incompatible with the CESR upgrade because there will be a fixed vertical support column in the middle of the endcap region in order to provide a stable support for the accelerator and detector components near the IR region. On the positive side, the CLEO III detector design minimizes the gap between the endcap and the barrel calorimeters which exists in the CLEO II detector by creating room near the beam pipe for the routing of the cables from the inner detectors. This means that the solid angle coverage of the endcap and the barrel calorimeters can be made to overlap. Additionally, employment of a projective geometry in the endcap calorimeter will improve its performance, particularly at the low energies, since a reduced number of crystals will be hit by photons. This in turn will reduce the effects of electronics noise on the energy and position resolutions and also reduce the probability that two adjacent showers overlap. These considerable improvements will be obtained by redesigning the endcap calorimeter for CLEO III.

Figure 44 is a sideview of the CLEO III detector with the proposed endcap design. Figure 45 shows a quarter of the endcap viewed from the interaction point looking down the beam pipe. The squares in the figure represent the front faces of the endcap crystals. The x and y coordinates correspond to tan θ cos ϕ and tan θ sin ϕ. The two quarter circles represent the front face of the end barrel crystals. Though the overlap between the barrel and endcap crystals is not perfect, this design shows a substantial improvement over CLEO II.

Figure 45: Interaction point view of a quarter of one the CsI endcaps.

4.5.2 R&D

Some of the issues which require R&D efforts are the following:

• We are investigating reducing the amount of material needed to support the entire endcap, particularly in front of the crystals and between the endcaps and the barrel. Accomplishing this objective while maintaining the mechanical integrity is a major design challenge. A structural analysis of each design needs to be performed.

• Remanufacturing the endcap into a projective geometry requires substantial machining of the existing rectangular crystals. Monte Carlo simulation is needed to understand the probable performance gains and their physics implications. This effort has already started.

• In order to achieve a projective geometry, there are two possibilities. Either we use a small number of highly tapered crystals, leaving others rectangular or we modestly taper most crystals. The effect of tapering on the longitudinal light-output uniformity must be understood before we can make a selection between these two options.

• To achieve a projective geometry requires the purchase of additional crystals. Arrangements with industry need to be explored.

4.6 IR mounting and assembly sequence

4.6.1 Mounting system

A primary concern in the design of the tracking system has been to eliminate, as far as possible, any structural material in the active tracking region. This is accomplished by holding the quadrupoles from sturdy pedestals in the magnet poles that are fastened to the fixed magnet yoke as shown in Figure 46. The beampipe, silicon detectors, and drift chamber are supported from this yoke without a permanent support tube crossing the active tracking volume. Nevertheless, during assembly it is necessary to have such a support tube to hold the beampipe, rare-earth cobalt (REC) quadrupoles, and silicon detectors as this subassembly is slipped into the drift chamber. Providing this space for a temporary support structure is a defining feature of the tracking system design, but the design does not preclude the use of a thin permanent support tube if the need arises.

The tracking volume is divided into two regions: the space from the beampipe to a radius of about 15.7 cm, which is occupied by the silicon detector, and the space from about 17.5 cm to 82 cm, which is occupied by the drift chamber. The annular region from 15.7 cm to 17.5 cm is reserved for the temporary support structure. This space is sufficient to provide a rigid support for the REC quadrupoles and clearance to extract the structure without damage to the detectors. To allow for extraction, the mounting structure must not only clear the outer radius of the silicon detector but also the cables and cooling tubes of the silicon detector and the REC quadrupoles.

The separation of the interaction quadrupoles into REC and superconducting parts is a crucial part of the design. The two REC quadrupoles are captured on the vacuum chamber and all three must be installed as a unit through the drift chamber. These quads are relatively small in radius and simple—they do not require the complicated electrical and cryogenic connections, nor the vacuum jacket of a superconducting magnet. This significantly reduces the clearance required through the drift chamber and allows for a larger, unobstructed, momentum measuring volume. On the other hand, the superconducting quadrupoles are not captured on the vacuum chamber and may be installed separately from each side. When installation is complete, the REC quadrupoles will be fastened to the superconducting quadrupoles, which are fastened to “keyhole” steel supports in the poles, which are in turn fastened to the fixed portion of the magnet yoke.

After installation, the silicon detector will be supported independently by the drift chamber. The beryllium beampipe, masks, shielding, and vacuum chamber extensions on each end will be supported by the quadrupoles with clearance to the silicon detector to provide the best isolation of the silicon detector from the storage ring. The option of supporting the beampipe assembly through the support of the silicon, leaving clearance to the quadrupoles, is left as an option.

Figure 46: The CLEO III interaction region after installation of the CESR components including the beam pipe, REC and superconducting quadrupoles, and vacuum pumps.

4.6.2 Assembly sequence

The subassembly consisting of the beampipe, vacuum chamber extensions, REC quadrupoles and silicon detector will be assembled in a fixture away from the magnet yoke. Since the REC quadrupoles and shielding material weigh a few hundred kilograms and cannot be allowed to place much stress on the thin beryllium beampipe or the silicon detector, the assembly fixture must provide independent support and alignment capability for the REC quadrupoles. The REC quadrupoles will be slipped over individual vacuum chambers before the flanges on these chambers are welded in place. After appropriate vacuum procedures and testing, these flanges will be sealed to the beryllium beampipe flanges. This results in an approximately 3 m long subassembly. There will be shielding material, cooling connections and other operations to complete at this stage (Fig. 47a).

The silicon detector will next be assembled around the beampipe while the subassembly is supported in the fixture. Attachment will be made through temporary brackets to the REC quadrupoles as shown in Fig. 47b.

With the silicon detector completely mounted, wired, and plumbed, temporary longitudinal braces will be connected. These occupy the radius from 15.7 to 17.5 cm radius and attach through temporary brackets to the REC quadrupoles as shown in Fig. 47c. The longitudinal braces also attach to a stiffening ring on each end to connect the braces independently from the attachment to the quadrupoles and to provide a means of lifting the completed unit from the ends. Thus, the two quadrupoles are rigidly attached to each other, independent of the initial assembly fixture, and the entire subassembly may be picked up for installation in the drift chamber.

At this stage the new drift chamber will be installed in the magnet (Fig. 48a) and supported at its outside edge by its final “crown”-type support similar to that used for DR 2. (Alternatively, the new drift chamber may be held by temporary supports which will be replaced later by the final supports to the “keyhole” section of the magnet.) The installation of the new DR 3 will be similar to the technique used for DR 2, namely, a long beam is inserted through the axis of the whole detector on which the drift chamber rides on wheels.

The subassembly of beampipe, silicon detector, and REC quadrupoles will then be slipped through the drift chamber (Fig. 48b). The brackets holding the silicon detector to the longitudinal braces will be removed and replaced by brackets fastening to the drift chamber (Fig. 48c). Brackets that temporarily support the REC quadrupoles from the tapered section of the drift chamber end plate will replace those that attach the quadrupoles to the longitudinal braces. By completing this sequence at each of eight azimuthal positions in series, the fragile components are never stressed. The temporary longitudinal brace/stiffening ring assembly is then free and can be removed.

The superconducting quadrupoles will be introduced from each end (Fig. 49a) and fasa)

b)

c)

Figure 47: Assembly of the beampipe, silicon detector, and REC quads. In a), one REC quad is shown mounted around the beam pipe along with the bracing connecting it to the opposite quad. In b), the silicon dectector is mounted and temporarily attached to the REC quads. In c), the longitudinal braces are attached and the assembly can be lifted.

a)

b) c)

Figure 48: Insertion of the beampipe, silicon detector, and REC quad into the drift chamber. In a), the drift chamber outline is shown before insertion. In b), the beampipe, silicon detector, and REC quad have been inserted into the drift chamber bore. In c), the silicon detector and REC quad support has be transferred to the drift chamber endplates and the longitudinal bracing removed.

tened to the final “keyhole” supports (Fig. 49b). The REC magnets will be fastened to the superconducting quadrupoles and released from the drift chamber. This completes the transfer of support of the REC quadrupoles from the longitudinal braces to the fixed magnet yoke and the transfer of the support of the silicon detector to the drift chamber. The beampipe, quadrupoles, and detectors can then be surveyed and aligned. The drift chamber electronics, cabling and utilities will be installed to complete the installation.

4.7 Detector backgrounds

Beam induced detector backgrounds have been studied in the tracking system proposed in this report. The simulation and its predictive ability have been described elsewhere[106]. The CESR Phase 3 lattice with 0.5 A per beam, a crossing angle of 2 mrad, and a residual gas pressure in the ring of 3.5 × 10^{−9} Torr are used to generate a distribution of beam particles and photons scattered from the gas and a distribution of synchrotron photons generated by the transverse acceleration of beam particles in the magnetic structure of CESR. These distributions are input to the GEANT-based CLEO simulation where the tracking system has been altered to conform to the CLEO III design and input to a detailed simulation of x-ray scattering for the synchrotron radiation component. The GEANT geometry used is shown in Fig. 50. Occupancies in the first layer of the CLEO III DR are compared with the present performance of DR 2. Occupancies in the first layer of the CLEO III silicon detector are compared with those expected from the CLEO II/SVX upgrade.

In the CESR Phase 3 interaction region most of the masking of the detector is accomplished by the superconducting quadrupoles that are placed close to the interaction point. The only additional masking that is needed are small tungsten inserts in the regions 20.0 cm < z < 29.0 cm. These inserts provide a tip at z = 26.0 cm that shields the beryllium beam pipe from direct synchrotron radiation hits. The inside of the beryllium pipe is coated with 10 μm of gold that provides a further factor of 2 reduction in the flux of x-rays that strike the first layer of silicon.

Table 25 shows the breakdown of detector backgrounds for the CLEO III upgrade. The backgrounds are low and they compare favorably with the CLEO II/SVX upgrade. For CESR Phase 2 with a total beam current of 600 mA, the occupancy in the first silicon layer is expected to be 0.04% and the present DR 2 has an occupancy of 0.27% for each microsecond of detector exposure time. While the predicted occupancies are ten times better than what is required for good track reconstruction efficiency—1%/μsec in the silicon and 10%/μsec in the drift chamber—another iteration of mask optimization could improve the backgrounds even further.

Under the masking scheme proposed, the silicon detector and on-board electronics will be exposed to 30 krad/year. The electronics must be able to tolerate a minimum of 150 krad total exposure which will require their fabrication in a rad-hard process. a)

b)

Figure 49: Installation of the superconducting quads. In a), the superconducting quad is inserted, bolted to the REC quads, attached to the support pedestal, and the REC quad support disconnected from the drift chamber. In b), the installation is complete. The silicon detector is supported from the drift chamber and the quads, beampipe, beam position monitors, and vacuum pumps are supported from the keyhole pedestal in the CLEO magnet pole piece.

Table 25: Predicted occupancies in the CLEO III tracking system. Occupancies are in % per microsecond of detector live time.

Source  Silicon Layer 1 DR Layer 1
Beam Gas  0.055  0.82
Synchrotron  0.017  0.00
Total  0.072  0.82

Figure 50: Model of CLEO III and CESR Phase 3 used in particle and synchrotron radiation background studies.

5 Front-end electronics, triggering, and data acquisition

In this section we examine the ramifications of luminosities (L) between 10^{33} cm^{−2}s^{−1} and 10^{34} cm^{−2}s^{−1} on the CLEO front-end electronics, trigger, and data acquisition systems. We briefly review the present status of these systems and the implications of their unmodified use at L = 10^{33} in view of the expected backgrounds. Trigger and front-end architectures applicable to high luminosity operation are discussed, and a data-acquisition system having sufficient bandwidth and flexibility is presented. Since we are designing a new electronics architecture that will be in use beyond the initial CESR Phase 3 upgrade, it must be extensible to L = 10^{34}, the upper limit that CESR can achieve.

5.1 CLEO II system

CLEO II data recording is controlled by a three level hardware trigger system and a software trigger level that follows data acquisition readout. The design is matched to the single-hit nature of the front-end electronics systems. The performance of the overall system as parameterized by dead-time is limited by this single-hit architecture and by the inability to perform simultaneous front-end acquisition and data readout.

5.1.1 Trigger

A level 0 trigger operates synchronously with the storage ring, the decision being made in the 360 ns between crossings of the seven beam bunches. In this way, failed level 0 triggers incur no dead-time. At 5.28 GeV and L = 2×10^{32}, the level 0 trigger reduces the 2.8 MHz beam crossing frequency by a factor of ∼300 to 10 kHz. Examples of existing level 0 triggers are:

• Any two separated barrel time of flight counters

• Opposite endcap time of flight counters

• One time of flight counter and a track segment in the vertex detector

• One high threshold barrel crystal region

• Opposite high threshold endcap crystal regions

A limitation of the present system is that mixed tracking and calorimeter triggers can not occur at this level because of their disparate signal development times. CsI energy trigger information is available after ∼2 μs. The self-resetting feature of the drift chamber electronics causes the wire hit information to be reset 800 ns after the discriminator fires if no trigger inhibits the reset.

Successful level 0 triggers incur experimental dead-time to freeze the contents of the front-end electronics and initiate the level 1 trigger hardware which has a decision time of 1.2 μs. Failed level 1 triggers entail another 1 μs dead-time to reset the front-end electronics. At 5.28 GeV and L = 2×10^{32}, the level 1 trigger reduces the 10 kHz level 0 trigger output by a factor of ∼300 to 30 Hz. This includes pre-scaling of QED processes in the endcap regions. Some examples of level 1 triggers in use today are:

• Back-to-back high threshold crystal regions in the barrel or in the endcap

• One time of flight counter, two track segments in the drift chamber, and one minimum ionizing crystal region

• One each high threshold barrel and endcap crystal regions

Successful level 1 triggers initiate, if appropriate, the level 2 trigger which consists of two hardware tracking processors. One uses the inner 10 layer drift chamber (VD) and the other uses the axial layers of the main drift chamber. One works with digital hit information formed from the OR of 4 drift cells. The other uses wire-level hit information to find tracks in predefined road-widths. The total processing time is 40 μs. Under 1994 Υ(4S) running conditions, this trigger rejected 50% of the level 1 triggers.

Accepted level 2 triggers initiate front-end readout by the data acquisition system where the data is buffered at the input to the VME-based event builder system. The trigger is then re-enabled and the data sparsification and event building proceed asynchronously. After event building, data are transferred to an on-line workstation where they are analyzed by the level 3 software trigger. Access to the full data allows this code to calculate such quantities as neutral energy balance and to use the inner tracking detector hits to eliminate beam-wall events. Level 3 reduces the trigger rate by about a factor of two.

The overall performance of the trigger chain can be simply determined by examining the fraction of events that are qualified as “good” by the off-line reconstruction code. Fifty percent of the events processed by the level 3 trigger are rejected. They are generally categorized as not containing charged tracks or not being interesting pure neutral events. The efficiency of the trigger system has been determined to be > 91% for qq̄ events and nearly 100% for Υ(4S) hadronic events.

5.1.2 Gating and event flow control

The CLEO II gating system uses a subset of the timing modules developed to operate the storage ring. Different gates are generated for the various detector types and fanned-out to the crates. The gate widths and placement relative to beam crossing time are software controllable. An event trigger signal is also fanned-out to the crates to initiate the transfer of data to the raw data buffers. The data acquisition system (DAQ90) receives a copy of this signal so that it can check for the completion of the event read-in. On completion, DAQ90 sends a signal to the trigger system to re-enable it. The CsI electronics have an additional fast clear signal to reset the input integrators after a successful level 1 but a failed level 2.

5.1.3 Front-end electronics

There are three categories of front-end electronics in the present CLEO II detector: digital systems associated with the trigger systems, LeCroy FASTBUS ADC’s for the CsI, and CLEO-specific time-to-voltage and charge-to-voltage sample/holds (S/H) for the various drift chamber, muon, and time-of-flight systems. The front-ends are susceptible to three sources of background: cosmic rays, stray beam particles, and synchrotron radiation. The first one can be dealt with at the trigger level and the rate is low enough that overlap with physics events is negligible. This is not necessarily true in the cases of stray beam particles and synchrotron radiation, however. For example, due to the CLEO II single hit electronics, which have a memory time of a couple of microseconds, ∼10% of the VD cells read for an event appear to be the result of synchrotron photons and stray beam particles.

5.1.4 Data acquisition

The data acquisition (DAQ) system was designed to work within the limiting features of the front-end electronics such that it introduces no additional dead-time. The analogto-digital conversion time of the S/H’s and the restriction that data transmission cannot occur while the front-end electronics are active determines the dead-time. This has led to a scheme where the data conversion occurs rapidly (1 μs) and the datum from each conversion is sent immediately to a remote buffer. Each front-end crate has a 16-bit wide, 12 MByte/s point-to-point link to a dedicated VSB dual-port memory. A typical crate has 2000 front-end channels and the total conversion and transfer time is 2 ms. This contributes 0.2% dead-time per Hz of level 2 trigger rate. The dual-port memories are serviced by a number of VME processors that do sparsification and event fragment assembly in a distributed fashion. The finished events are transferred via Ethernet to a workstation running the level 3 trigger code. The average CLEO hadronic event is presently about 8 kByte and the average readout event is 5 kByte.

5.1.5 System performance

The trigger/DAQ performance is summarized in Table 26. As can be seen, the level 1 latency time and the DAQ conversion/readout dominate the deadtime. They will contribute

Table 26: CLEO II trigger performance summary at L = 2×10^{32} cm^{−2}s^{−1}.

Input Output Rejection Latency Dead-time
rate  rate  ratio  time  %
Level 0 2.8 MHz 10 kHz  300:1  0  0.0
Level 1 10 kHz 30 Hz  300:1 2.2 μs  2.2
Level 2  30 Hz 15 Hz  2:1  40 μs  0.1
DAQ  15 Hz 15 Hz  2 ms  3.0
Level 3  15 Hz  7 Hz  2:1 variable  0.0
Total  5.6

to massive dead-times at L = 10^{33} if architectural changes are not made.

5.2 CLEO III

We present an overview for proposed general architecture of the CLEO III front-end electronics, trigger, and data acquisition systems. This has been driven by several conditions:

• Luminosity

• Backgrounds

• Live time

• Scalability

• Funding

The primary goal of the detector upgrade is to operate with CESR Phase 3 at luminosities up to L = 3×10^{33} cm^{−2}s^{−1} and to easily accommodate L = 10^{34}. The exact accelerator bunch structure that will achieve these luminosities is uncertain at this time, ranging from nine bunch trains with 5 sub-bunches to almost continuous bunches 14 ns apart. Clearly, any CLEO III design choices must be compatible with any machine bunch structure.

Particle and synchrotron radiation background rates are uncertain. Different CESR topologies make their prediction even more uncertain. Furthermore, extrapolations to luminosities fifty times greater than today cannot be viewed as reliable. From discussions below, we will see that background rates in detectors near the beam pipe can be as high as 100 kHz.

Some CLEO III electronics cannot easily be upgraded to to accommodate the full reach of CESR luminosity beyond Phase 3. Among these systems are the front-end electronics (preamps, data boards, and the trigger information derived from the front-end boards), the board and crate level event buffering, the crate back-plane technology, and the fast timing system including event flow control. The electronics architecture must accommodate operation beyond L = 2^{33} just as the CLEO I and II architecture will have accommodated 15 years of luminosity improvements.

Items that are required to operate CLEO and that can be enhanced or upgraded include the hardware trigger processors and decision logic, the event builder data collection and assembling, the level 2 trigger, the level 3 trigger implementation, and the data delivery mechanism to the event builder/level 3 trigger processor(s).

We briefly state the modifications required for the CLEO III detector. In the trigger system we eliminate the lowest level trigger which is driven by the beam crossing rate; track and energy information can be used in coincidence in the lowest level trigger; the deadtime is determined by the level 2 trigger latency, ∼10 μs. In the drift chamber we preserve sub-nanosecond timing precision with memory times up to 2.5 μs; both charge and time measurements are performed; and multi-hit capability is provided to deal with track congestion and pile-up at small radii. The silicon detector requires an on-detector front-end IC optimized for use in CLEO which digitizes and buffers data in 10 μs and may provide trigger information. The particle ID system requires development of a preamp and readout electronics. The CsI and muon systems have their readout electronics upgraded for 10 μs conversion and data buffering. A new DAQ architecture is used to readout the detector at up to 1 kHz with little or no dead-time impact by using several levels of buffering to allow overlapped readout and acquisition.

5.2.1 Trigger

We will see that if we want to contemplate higher background and physics rates, we must decrease the product of the rate times latency of the lowest level trigger as well as the digitization/readout time of the front-ends. We discuss the trigger solution first.

Backgrounds

A good understanding of the expected low-level trigger rate for L = 10^{33} cm^{−2}s^{−1} and greater is desirable when designing the architecture for the trigger and front-end electronics. Unfortunately, this is a difficult number to predict. To set the scale for the trigger rate, the cross sections of interesting processes at the Υ(4S) are given in Table 27. This does not include two photon induced processes. At the Υ(1S) the resonance contribution to the hadronic rate is five times larger than at the Υ(4S). At L = 10^{33}, 1 nb of cross section gives 1 Hz of trigger. In round terms, the physics trigger rate will be ∼100 Hz at L = 10^{33}. Assuming that any new trigger system will provide the same rejection and efficiency as the present one, we would expect a hardware trigger rate of about 200 Hz to 2 kHz without any pre-scaling of QED events.

While the interesting physics event-rate can be scaled from present CLEO experience,

Table 27: Cross sections at 5.28 GeV for an active detector region extending down to 330 mrad from the beam line.

σ_{T otal} σ_{Barrel} σ_{Endcap}
(nb) (nb)  (nb)
e^{+}e^{−} → e^{+}e^{−}  72  19  53
e^{+}e^{−} → γγ  6.2  3.7  2.5
e^{+}e^{−} → μμ  0.72 0.60  0.12
e^{+}e^{−} → ττ  0.72 0.60  0.12
e^{+}e^{−} → hadrons 4
total  84

any fast (level 0) hardware trigger rate is largely driven by the IR background conditions. There are several ways to estimate this background rate for the L = 10^{33} configuration, although none is ideal. The first of these is to accept the rough estimates of the CESR machine designers, namely that changes to the IR and machine lattice will no more than double the background rates seen by CLEO II for each of the next two machine modifications (the completion of phase 2 and the introduction of phase 3). This means that at L = 10^{33}, the level 0 trigger rate should be 40 kHz or less.

One can also use an extrapolation of the CLEO II experience operating over a factor of ten in luminosity. Figure 51 shows the level 0, 1, and 2 output rates over a factor of 3 in luminosity. All rates show a nearly linear dependence. Extrapolating to a factor of 3.3 increase in luminosity to L = 10^{33} gives a level 0 rate of 50 kHz.

Finally, lost beam particles and synchrotron radiation can be carefully simulated. The advantage of this technique is that the algorithm used can be tested by simulating the present machine/detector configuration and comparing with CLEO II observables, such as level 0 trigger rates and wire chamber currents. Comparisons of this sort have shown good agreement between measured and simulated quantities, giving us some confidence that the simulation will be correct for the 10^{33} and future CESR configurations as well. This method predicts that the level 0 trigger rate will remain below 25 kHz at L = 10^{33}.

Since a definitive prediction of the level 0 trigger rate is not possible, the conservative approach we adopt is to take the largest of the above estimates (50 kHz). Furthermore, to allow for the initial period of machine commissioning when the backgrounds will inevitably be large, any new trigger/front-end designs should accommodate a level 0 rate of 100 kHz.

Summarizing, at a luminosity of 10^{33} cm^{−2}s^{−1} we should expect a level 0 (i.e.,prompt) trigger rate of up to 100 kHz, and a final hardware trigger event rate of about 200 Hz.

Figure 51: CLEO II measured trigger rates as a function of luminosity. Open squares are from December 1992, closed squares are from Summer 1993, and the triangles are from December 1993.

Level 1

A trigger architecture similar to that in CLEO II will lead to large dead-time introduced by the level 1 latency if it is driven at the level 0 rates described above. We propose to do without a trigger level that attempts to track the beam crossings. While technologically challenging, it is not impossible to build a trigger that operates at 72 MHz, but there is no need for this since drift chamber signals used in the trigger are not available for up to 500 ns after a beam crossing and since the interesting event rate is less than 1 kHz compared to the many megahertz of beam crossings. Event overlap is negligible but events can have about 1% occupancy due to background.

Level 1 will be the lowest level trigger combining the features of both the present level 0 and level 1 while maintaining the overall 10^{5} rejection. The total processing time for this level can be as long as 2.5 μs, an upper limit set by the need to start gating the CsI ADC’s. The rate at which the level 1 trigger decisions can be generated is naturally set by the longest drift times used in the tracker trigger processor. This will be between 200 and 1200 ns depending on the drift chamber gas and cell design choice.

Modest levels of pipelining will be required to implement this trigger. The pipeline also makes it possible to time-align the data from the CsI and trackers which generate trigger information on very different time scales relative to the event beam crossing. Since some detector components will have an uncertainty in the arrival time of their trigger information that is larger than the pipeline bucket width, their information must be stretched across several buckets if the logic interrogating the trigger pipelines is to obtain information from a unique set of pipeline positions. Equivalently, the trigger information could be stored only once and the interrogating hardware be made smart enough to examine several buckets in a single pipeline.

Note that we are not designing a dead-timeless system. Level 1 is dead-timeless until a successful trigger occurs. A small dead-time, 10–100 μs, is then incurred for data conversion and the level 2 decision time. This greatly simplifies the design. Furthermore, dead-time after a successful level 1 may be required for silicon readout systems that cannot perform simultaneous read and write.

The CLEO II trigger has been generally successful in providing the physics of interest to the collaboration. For L = 10^{33} we can maintain the general trigger categories used today, including low multiplicity final states of τ and γγ physics. Improvement is possible in tracking by using finer granularity hit information and in energy triggering by providing physics-oriented quantities such as total energy and energy flow.

Tracking triggers will continue to rely on linking track segments found in adjacent clusters of axial layers. The loosest pure track trigger would comprise one track spanning the entire tracking radius. If this is too loose, an additional track that exits the main drift chamber at its half radius can be demanded. One possible improvement to the present track segment approach where four or more wires are OR-ed to form a hit bit is to use each axial wire in the segment finding. This is a simple brute force extension of the present track segment processor design. It will provide tighter x-y vertexing by reducing the segmentation at small radii. Also, segments can be momenta binned and this additional information used in the linking of segments.

The CLEO II energy trigger is essentially topological in nature, examining with very coarse granularity whether regions of the calorimeter are above either or both of two thresholds (100 MeV and 500 MeV). The primitives used are two level-discriminator bits available for each 4×4 crystal array. The present trigger OR’s these into 16 regions in the barrel and 16 regions in the endcap. This is adequate for triggering on a few large showers or on constrained geometry of lower energy showers. An upgraded trigger could use these bits at their full granularity. In addition, there is another set of unused discriminator bits where the leading edge is determined by a level-threshold and the trailing edge is determined by the zero-crossing of a fast shaped analog sum of the 4×4 crystals. These may be important to “bunch-stamp” neutral energy triggers or to correlate with track triggers.

While the digital approach should be maintained, it should be augmented with an analog energy trigger. There are unused analog current sums from each 4×4 crystal array which could easily be made available at the front-end crate level. This trigger can be done in analog form with current sums or digitally with sums from FADC’s. In an analog implementation, one would conditionally combine the currents by digital thresholding to reduce noise contributions from quiet crystal arrays, and then threshold the combined analog sum. By appropriate weighting of the currents, total, transverse, and longitudinal energies can be formed. A level 1 trigger of this type could serve as the seed for a more sophisticated global energy trigger at level 2. These changes are attractive but their efficacy should be studied by Monte Carlo first.

It is conceivable that the particle ID system would be fast enough to contribute to the first level trigger. The major benefit could be to identify the event’s beam crossing number which would give a start time for the drift chamber and CsI gate.

A desirable feature of the level 1 trigger is that it be in continual operation even when the DAQ system is incapable of handling more events. The delivery of the level 1 triggers to the crates and level 2 can be vetoed by the buffer overflow signals from the front-end boards and crates and from the event builder. In this way, the level 1 trigger can be used to measure actual experimental dead-time since we will not be measuring the gated or ungated luminosity of each beam individual beam bunch crossing.

As for any trigger, the input data and hardware algorithms must be easily modeled by the detector simulation codes.

Level 2

It should be noted that a level 2 trigger may not be required at all if level 1 can reduce the event rate to less than 1 kHz. In this scenario, the on-line software trigger—level 3— will be required to play a bigger role on the overall rejection factor of the system. Any decision to proceed along this route would have to take into account the increased bandwidth requirements within a front-end crate and from the crate to the level 3 node(s). Advantages would include a simpler hardware system with great flexibility that is easier to understand and simulate.

If it is needed, the level 2 trigger is initiated by a level 1 trigger. Since front-end data conversion and storage are also initiated by level 1, level 2 can take as long as the slowest converting detector device without increasing the dead-time. If the front-end systems support being interrupted during conversion, a failed level 2 trigger can abort the conversion and reenable data taking. The present level 2 track processors require 50 μs and a new design should be able to cycle in ∼10 μs.

The level 2 tracking trigger would make use of the same wire hit information used at level 1. It would perform a wire-by-wire track linking algorithm and report a histogram of found momenta. The track finding would be in r − ϕ only since z track information is encoded in the stereo layer drift times and in the silicon. The former is hard to use in hardware and the latter may not be available to the trigger. After digitizing the ∼500 crystal energy currents, the data can be processed to provide refined total and projective energies. It is also possible to provide a list of cluster energies and photon pair masses.

Figure 52 shows a block diagram of the trigger.

5.2.2 Gating and event flow control

The gating logic interfaces to the CESR timing systems using components from the accelerator master timing system. The various gates and clocks used by the front-end and trigger systems and their synchronization with the beam bunch crossings are implemented with these CESR components. The timing system will be packaged in VME and be remotely software programmable.

The event flow control system and its network coordinate the DAQ, trigger, and gating systems. Gates, level 1 and level 2 trigger accepts, level 2 trigger aborts, and event numbers are distributed to the crates. Control of programmable calibration pulsers is also done here. Data collected via this network are the event buffer availability in front-end boards, front-end crates, and event builder memories. A grand OR of the status of these systems is used to inhibit—throttle—further event triggers to the front-ends.

It may be impossible to measure the individual luminosity of each bunch crossing in CESR. This is the motivation for having the level 1 trigger in continuous operation even when the DAQ system cannot handle more events. The throttle signals from DAQ convert successful level 1 triggers into failures. The gating and event flow control system provides scalers to record this activity allowing a reliable dead-time correction to be made to the off-line calculated luminosity.

Figure 52: Block diagram of CLEO III trigger system showing level 1 and 2 triggers and their interaction with the gating and event flow logic. The silicon and particle ID systems are shown as contributing to both levels of the trigger but this is yet to be determined.

Table 28: Estimated performance of present CLEO II trigger at L = 10^{33} cm^{−2}s^{−1} with single-hit, ADC per channel front-end electronics. The level 2 decision time is assumed to be 10 μs.

Input Output Rejection Latency Dead-time
rate  rate  ratio  time  %
Level 1 18 MHz 200 Hz  10^{5}:1  0 μs  0
Level 2 200 Hz 100 Hz  2:1  10 μs  0.2
DAQ  100 Hz 100 Hz  0  0
Level 3 100 Hz 50 Hz  2:1 variable  0.0
Total  0.2

An additional feature required of the event flow control network is a mechanism to perform a back-plane reset in the VME and FASTBUS crates. This is usual to recover from processor “hangs” that require system resets.

5.2.3 Front-end electronics

New front-end electronics will be required for several reasons. First and foremost, we must significantly reduce the readout dead-time. Although the present hardware supports a rather cumbersome way of halving this time, the improvement of a factor of twenty that we seek can not be realized without a complete redesign of the front-end digitizing and buffering hardware. Secondly, waiting ∼2.5 μs after a good event’s crossing for the trigger to assert dead-time means that single-hit electronics will effectively integrate over this time. This may not be acceptable in view of synchrotron-radiation background levels, especially in the inner regions of the tracking volume. The solution is to implement multi-hit front-end electronics, particularly in the regions of high occupancy, but preferably throughout the entire tracking system in order to maintain uniformity of architecture.

The design goal of the front-ends will be to digitize and buffer the entire detector in less than 100 μs, preferably less than 10 μs. This will give a readout induced dead-time of 1–10% at a 1000 Hz event-rate limit. This is acceptable for any foreseeable luminosity upgrade. Dead-time from a system having the trigger and front-end characteristics discussed in the above sections is shown in Table 28. Block diagrams of the front-end system components are shown in Fig. 53.

Drift chamber

A preamp will be mounted at the drift chamber endplates. This will provide some shaping and line driving ability. The signal is not discriminated here because we are interested in the wire charge for dE/dx measurements. To help reduce the mass at the endplates, monolithic preamps will be investigated. Smaller connectors and thinner cables to the crates boards

123

Figure 53: Block diagrams of CLEO front-end electronic systems.

will help to free up valuable space.

There are possible commercial solutions available for implementing a multi-hit front-end system. LeCroy has introduced an interesting timing circuit, the MTD133, that records the value of a 16-bit counter for up to 16 input transitions. A common stop signal records a final clock value and prepares the circuit for readout at which time the final clock value is subtracted from each of the up to 16 time recordings. The time resolution is 0.75 ns and the maximum depth is 40 μs.

Set up to record times on both rising and falling edges, the MTD133 can also be used as the counter in a Wilkinson run-down ADC. An experiment at CEBAF will multiplex timing and charge information into the same MTD133 channel using pulse width as markers. In future versions of the MTD133, LeCroy promises several additional features, including 0.5 ns resolution over a 32 μs range. Since Lecroy has not yet produced an MTD133 with the required performance, we must carefully evaluate other commercial solutions and the system used in other experiments.

The use of commercial TDC/ADC’s will require a separate development of a system to split the preamp signals into timing and charge chains. The timing chain requires a discriminator whose output is compatible with the TDC. The charge chain may need to integrate the signal depending on the charge measuring scheme—TDC or ADC. The timing chain may need to provide the wire hit information to the trigger system. For some LeCroy FASTBUS modules this is available on the auxiliary connector.

Calorimeter

The present ADC system provides 12-bit resolution on two scales to give 15-bit dynamic range. It requires 750 μs to convert. It is not self-sparsifying, does not have multi-event storage, nor does it allow simultaneous read and write. Desirable features of any new ADC system are increased resolution, 14 bits with a 17-bit dynamic range, and lower coherent noise. The present ADC resolution is comparable to the noise and it would be nice to better resolve the pedestal to study coherent versus incoherent noise contributions. The preamps and mixer-shaper system will be retained with some re-work of the trigger section. To use the analog trigger sums at the mixer-shaper card level requires a new upper back-plane.

Multi-hit electronics for the calorimeter is not a requirement. The design of the level 1 trigger has been constrained to make a decision before 2.5 μs specifically for the purpose of gating the calorimeter conversion hardware, and we will use this to greatly simplify these front-ends. A single-hit system with sufficient dynamic range can be constructed by viewing each crystal with a pair of 14-bit ADC’s in a dual-range configuration. The ADC outputs are stored in moderate depth FIFO’s that are read out over a digital back-plane asynchronous to the trigger. The requirement that this system convert and buffer its data at the 10 μs time scale (for a reasonable cost) is not particularly onerous given recent developments in commercial ADC technology. While two ADC’s per channel may appear excessive, it avoids the expense and bulk of fast, precision analog multiplexors. In addition, massively parallel digitizing and buffering at the front-end opens up the possibility of doing intelligent sparsification in the front-end crates, removing any bandwidth bottleneck presently associated with the “slightly parallel” DAQ90 sparsification scheme.

The use of the MTD133 for precision calorimetry is unclear at this time. With a dual range system where each charge is measured to 12-bit precision, a given crystal signal can occupy up to 4096 of the 65535 counts of the MTD133 time span. This essentially introduces an additional 3 μs crystal dead-time for each hit that starts the integrate/ramp-down process. This is probably not a problem since the occupancy of the crystals is small. There are other fast converting ADC systems commercially available.

Muons

A single range version of the gated integrator ADC’s used for the CsI will work in the muon system. A preamplifier-shaper specific to the muon system will need to be developed.

Silicon

On-detector silicon readout electronics is always multiplexed due to space, density, and power dissipation limits. It can easily define the performance of the DAQ system. We require an analog-storage silicon front-end (for best spatial resolution) that converts and buffers its data in 10–50 μs and allows overlapped readout and acquisition. On-chip sparsification can help to meet this goal. Serial-bus interfaces to the IC’s can reduce the bulk of the cable plant.

It is probably beyond the means of the CLEO collaboration to develop an entirely new silicon readout integrated circuit. Happily, this is an area of much activity in the HEP community. D0 and CDF are building new detectors based on the SVX-II circuit that provides analog pipelining, on-board digitization, and sub-10 μs readout. Two LHC R&D projects, RD2 and RD20, are developing analog pipelined, simultaneous write/read silicon front-ends. The time scale of these developments and others may be compatible with CESR phase 3 and it behooves us to leverage off these efforts. Modification of another experiment’s electronics to CLEO specifications is definitely within the abilities of the collaboration. One area of customization is to provide strip OR’s for triggering purposes. The selection of an existing front-end IC will depend on its compatability with the detector strip pitch, noise performance, and tolerance to an integrated radiation dose of ∼1 Mrad.

The crate-based data boards that will control and receive data from the detectors will probably be specific to CLEO. These entail sequencers, FIFO’s, sparsification logic, and multi-event buffers.

Particle ID

Since a particle ID scheme has not been chosen, the design of the relevant front-ends is not possible at this time. We will simply decree that the parameters of such a system will have to fit within the constraints described above, namely that it must be able to tolerate a 2.5 μs delay between the actual event and the time it is “strobed” and that converting and buffering the data should take ∼10 μs.

Trigger systems

The trigger system has another role as a source of event data. Like the other front-end systems, it must buffer multiple events at the individual board level and at the crate level.

5.2.4 Data acquisition

The DAQ architecture must accommodate operation up to 1 kHz readout rates. Some aspects of DAQ will support this initially since this choice affects the design of the front-end systems and the data collection network. Other aspects of DAQ are scalable and upgradable. This is especially true of the event builder and level 3 trigger nodes.

Front-end issues

To minimize dead-time from the DAQ system, the front-end boards will need to provide a small amount of event buffering. The readout of these events must be able to occur in parallel with either recording data between triggers or with digitizing data from a later trigger. Only the conversion time of ∼10 μs drives the dead-time. Data transfer to the crate controller can take 1 ms on average and occurs asynchronously.

Local buffers can overflow if they are not emptied fast enough by the crate controller. A trigger throttling signal needs to be generated from each crate to indicate that the board-level buffers are approaching overflow. This feeds into the event flow control network.

If data conversion time is longer than a level 2 trigger decision time, the front-end board conversion must be capable of being aborted. Similarly, a mechanism for flushing level 2 rejected events from local buffers must be available. It is desirable that data conversion and event buffering all occur in less than 10 μs to give a dead-time contribution of 1% at 1 kHz. Conversion times as long as 50 μs would not be disastrous.

Crate level issues

The packaging of the front-end electronics will probably not be homogeneous. The trigger system and other CLEO-specific electronics will probably be implemented in VME or VXI. At the moment, the most cost effective commercial electronics for time and charge measurements are packaged in FASTBUS.

In either case, a processor will be resident in each crate. Note that most commercial front-end boards are self-sparsifying and that the crate processor will be relieved of this chore. Primarily, it will read an event’s worth of data from the front-end boards, assemble them into an event fragment tagged with an event number, buffer it, and transfer it to the event builder/level 3 complex. Data is pulled from the front-end boards by the controller, which works from a list of outstanding events to process. Data is pushed from the crate controller into the data collection network as long as there is buffer space on the receiving end. Our present thinking is that the current overhead for pulling the data from the crates at the request of the event building is too high. This view may change.

As with the front-end boards, there is a finite amount of event buffer space available in the crate controller. A throttling signal needs to be generated from each crate to indicate that the buffer is approaching overflow. This feeds into the event flow control network.

For VME, there is an abundance of commercial processors and network interfaces. For FASTBUS, the choices are few. The most modern and flexible FASTBUS based processor is the MIPS-based FNAL FRC. In both cases, a uni-directional, high-speed data transfer port is required for the event data and a local area network (LAN) connection is required for configuration and control. Whether event flow control can be incorporated into the LAN is up for discussion. Additionally, VxWorks will become the real-time operating system of choice since it is required to support the FRC and more processors and I/O devices are supported under it than under our present pSOS^{+m}.

We observe that VME does not provide a mechanism for sparse scan readout, i.e.,slave boards cannot signal that all data has been transferred. Of course, this is not a problem in FASTBUS. The impact on the effective back-plane bandwidth has to be carefully understood.

Bandwidth

The present CLEO II readout event size is about 5 kByte. The size of events for an upgraded detector will depend on the kind of multi-hit electronics used, the front-end solutions adopted for the silicon and particle-ID components, and the overall detector occupancy. A conservative estimate is 25 kByte per event, spread over tens of front-end crates. The aggregate bandwidth at 1 kHz readout rate is about 25 MByte/s. The multiplicity of crates means that the required bandwidths for the crate back-plane and data link from the crate to the event builder/level 3 complex is modest. The impact of the full bandwidth is on the design of the event builder. Note that this bandwidth is estimated to be needed at luminosities approaching L = 10^{34} and need not be supported in an event builder at L = 10^{33}.

Data collection network

A variety of data collection schemes have been considered: extensions to the CLEO II system, packet and channel oriented switches, and central shared memory. For reasons of cost, timeliness, and implementability, the shared memory approach is preferred. It is believed that the CLEO II system can not scale much beyond 100 Hz. ATM and Fibre Channel switch-based systems are not yet well enough established for us to commit to them. Ring-oriented SCI is also early in its life cycle. The useful bandwidth of FDDI may not be adequate. They, along with HIPPI, are also relatively expensive, especially when implemented in each crate. A further restriction is that the use of FASTBUS greatly limits the commercial interfaces available. In-house engineering of ATM or other complex network interfaces is beyond the collaboration’s capabilities.

An alternative is to implement a point-to-point link from each crate to a shared memory system. Each point-to-point link terminates in a individual dual-port memory as in CLEO II. The link technology could be a version of the present CLEO II data link, LeCroy ECL-line, or a TAXI-like serial link on copper or fiber at a few hundred megabits per second. The FNAL FRC allows personalization of its data output port. The existence of what is already available for this port or what can be easily implemented will determine the link structure of CLEO III.

The centralized memory system is the terminus of that part of the DAQ system which must support 1 kHz operation with 25 kByte events at the time of commissioning. The rest of the system can be scaled with luminosity and event rate.

Unlike the problem of off-line computing, DAQ on-line computing is primarily an I/O problem requiring modest computing performance—1 Mips/event—but capable of absorbing data at 25 MByte/s and generating 6 to 12 MByte/s of output. Note that these are effective rates which will require at least a factor of two greater peak rates. A distributed level 3 system comprised of workstations requires the development of a full bandwidth event builder, an event distribution scheme, and interface to the workstation. The merging of the output data streams over conventional LAN technology is problematic even with CLEO II’s modest output data rates. A single box solution providing good I/O and CPU performance is appropriate here. Multi-processor computers with 1000 Mips and easy-to-interface I/O are available today. The merging of output data is easy since there is no data movement from the box over a network.

We observe that powerful computers such as the SGI Power Challenge series and the DEC AXP processors with FUTUREBUS+ will allow the attachment of multiple VME crates. A single node terminus for the links can serve as event builder and as level 3 processor. Essentially, we merge the present DAQ90 hardware with the level 3 workstation. Initially, processor CPU cycles can move data from the dual-port memories. As data bandwidths exceed 10 MByte/s, it may become necessary to add VME-resident data movement engines to DMA the data into the host memory. A possible configuration of a memory-based system is shown in Fig. 54 along with a switch-based system.

The cost of the above processor is not included as part of the CLEO III upgrade. The interfaces to the system are included.

This approach limits the custom hardware requirements to the link hardware, possibly the dual-port memories, and the host software. We leverage the computer vendors’ ability to interface to VME and to provide the CPU performance in an easily maintainable and programmable environment.

5.2.5 Slow controls

Less challenging than the event data collection system, but critical for detector operation,

Figure 54: Block diagram of possible CLEO III DAQ systems: a) memory based and b) switch based.

is the slow control system for controlling and monitoring the myriad of support systems necessary for data taking. These include high and low voltage systems, gas mixing and monitoring, environmental conditions in and around the experiment, the cryogenics plant, and CESR conditions. A detailed inventory is too large to document here. We lump these together under the heading of slow controls. An important component of this is the shift operators interface. CLEO III will be moving away from VAX-based systems and their Qbus. Many of the interfaces to the above slow control items will need to be re-invented. This presents an opportunity to take a new approach using distributed intelligence. For example, the present gas mixing system and the high voltage program share the same computer as the operator interface. These systems and others could be operated as autonomous systems using a distributed compute environment. This could be based on VME processors running a real-time operating system networked with Ethernet. Another approach would use PC’s running Lab Windows or PC Spuds where the hardware is inexpensive, the development environment is friendly, and the graphical tools make complex, informative displays easy to code.

5.2.6 R&D

Silicon

• \underline{Front-end.} Design and develop the on-detector integrated circuits which include the functions of preamp, analog memory unit, ADC, sparsifier, control/data links, calibration, and possibly trigger primitives.

• \underline{Data boards.} This off-detector system is responsible for sequencing the on-detector electronics, receiving the sparsified data, buffering multi-events, and possibly delivering data to the level 2 trigger.

• \underline{Slow controls.} Design the system for monitoring cooling, power, other environment quantities, and position.

Drift chamber

• \underline{Front-end.} Design and develop a preamp/shaper/line-driver for mounting on the drift chamber endplate. Design and develop the calibration pulsing and distribution system. Design the cable plant between the on-detector electronics and the crate-based electronics.

• \underline{Data boards.} Design and develop a line-receiver/discriminator/integrator and a triggerprimitive generator. Design or procure the TDC/ADC system.

• \underline{Slow controls.} Design the low and high voltage distribution and monitoring systems, the gas mixing and monitoring system, and the environment monitoring system.

Particle identification

• \underline{Front-end.} Determine if an analog or digital system is required and then develop the appropriate preamp, shaper, discriminator, and line-driver. If trigger information is desired, supply the requisite signal drivers.

• \underline{Data boards.} Develop a line-receiver/discriminator and trigger-primitive signals if required. Procure an ADC/TDC system.

• \underline{Slow controls.} Design the low and high voltage distribution and monitoring systems, the gas mixing and monitoring system, and the environment monitoring system.

CsI calorimeter

• \underline{Data boards.} Procure the ADC system. Modify the digital trigger bits if required and implement physical trigger quantities based on available analog trigger information.

• \underline{Slow controls.} Interface the mixer/shaper crates to the control system and provide monitoring for low voltages, crate and preamp cooling systems, and other environment monitors.

Muon

• \underline{Front-end.} Design and develop a preamp/shaper/line-driver— this may be identical to that of the drift chamber but mounted near the data boards. Design and develop the calibration pulsing and distribution system.

• \underline{Data boards.} Procure the ADC system.

• \underline{Slow controls.} Design the low and high voltage distribution and monitoring systems, the gas mixing and monitoring system, and the environment monitoring system.

Trigger

• \underline{Level 1 Processors.} This level of the trigger will require development of processors for tracking and calorimetry and the decision logic. Simulation of the trigger logic, development of codes for inclusion in the physics Monte Carlo, and testing on simulated physics and background events will be done.

• \underline{Level 2 Processors.} This level of the trigger will require development of processors for tracking, calorimetry, possibly silicon, and possibly particle identification, and the decision logic. Simulation of the trigger logic, development of codes for inclusion in the physics Monte Carlo, and testing on simulated physics and background events will be done.

• \underline{Level 3.} This software trigger requires the development and simulation of algorithms.

Flow control

• \underline{Clock/gating system.} Develop a uniform clock and gating distribution system and interface to the CESR timing system.

• \underline{Event and ID distribution.} Develop a uniform mechanism for distributing event triggers, event tags, and event aborts to the data crates.

• \underline{DAQ throttles.} Develop a uniform mechanism for collecting the status of the various event buffers and inhibiting further triggering as indicated.

• \underline{Dead-time monitors.} Provide a mechanism for measuring the deadtime induced by the trigger and data acquisition systems.

• \underline{Configuration software.} Develop code to configure and manipulate the gating and throttling system.

• \underline{Simulation.} Simulate the system to establish its viability and integrity.

Data acquisition

• \underline{Crates.} Establish a small set of crate formats for the crate-based electronics. Provide an environment monitoring port to the slow control system. Establish the location of the crate based electronics and any special housing requirements.

• \underline{CPU and OS.} It is assumed that each data crate will have a local processor. Select a small number of processors and a common operating system, develop coding standards and code. Select a LAN technology for non-data communications with the processors.

• \underline{Data links.} Develop a common data link protocol and hardware for moving event data from the crates to the event builder memories. Develop the event builder buffer memories.

• \underline{Online computer.} Establish coding standards and generate event building and level 3 trigger codes. Allow partitioned readout of the experiment for simultaneous maintenance of the sub-detector systems.

• \underline{Trigger throttles.} Generate signals for the event flow control throttling network.

Slow controls

• \underline{Development environment.} Provide a uniform coding environment for all slow control code development.

• \underline{Standard alarms/limits package.} Provide uniform and rugged alarms and limits utilities.

• \underline{Cryo interface.} Interface to the He cryogenics plant.

• \underline{CESR interface.} Interface to the CESR control system to share information on quantities of mutual interest.

• \underline{Operator interface.} Provide the interface for the shift operators including run control, alarms, and displays of time trends of detector systems.

6 IR design

6.1 Overview of CESR luminosity upgrade

A five-year program to increase CESR luminosity by a factor of 5 to 10 was begun in Fall 1993. The program is staged in two parts which are referred to as “Phase 2” and “Phase 3.”

Phase 2 operation is planned for the period from mid-1994 until installation of the CLEO III detector components in 1997. In this phase we will use bunch trains with a small horizontal crossing angle to store beam currents up to the limit of the 5-cell RF cavities (approximately 300 mA/beam). Nine trains of 3 bunches each will be used. The IR quad configuration will be modified to provide the larger horizontal aperture required for the separate electron and positron orbits. The peak luminosity for this operation will be around 6×10^{32} cm^{−2}sec^{−1}.

Phase 3 equipment will be installed in CESR at the same time as the installation of the CLEO III detector described in this report. Replacement of the four 5-cell RF cavities with four superconducting 1-cell cavities and modification of the IR to allow more closely spaced bunches will permit increasing the beam currents. This configuration will allow us to reach the limits of the ring vacuum chamber (gas desorption and available pumping rate) using 9 trains of 5 bunches each. The proposed IR layout will accommodate values of β_{V}^{∗} to 1 cm, making further increases in luminosity possible if the bunch length can be shortened comparably. The expected peak luminosity in the Phase 3 configuration is 1 to 2×10^{33} cm^{−2}sec^{−1}.

The IR quad arrangement must be modified from its present configuration to provide greater effective horizontal aperture for crossing angle orbits. The crossing angle operation requires additional aperture in the horizontally focussing quads. This may be achieved by exchanging the two electromagnetic quads on each side (Q1 and Q2) and lengthening the permanent magnet quad using spare sections from the former North area permanent magnet quads. Energy flexibility (4.7 to 5.6 GeV) is possible by moving the Q1 quadrupole longitudinally.

As the bunch spacing is decreased to 14 ns in Phase 3 operation, the first parasitic crossing point moves from 4.2 to 2.1 m away from the interaction point. The combination of small separation and large β_{V}^{∗} with the Phase 2 optics results in an excessive long range beam-beam tune shift. A high gradient (40 Tesla/m) superconducting quad focuses the beam vertically in the Phase 3 configuration to minimize the tune shift at the parasitic crossing. A 30 cm long stepped permanent magnet quad provides additional focussing as well as shielding of background.

6.2 Accelerator components

Components associated with two major accelerator systems, vacuum and magnets, must share space with the detector. This section describes the requirements and design considerations which led to the present design concept.

6.2.1 IR magnets

In order to reach the high charge densities required for high luminosity, a very short final focal length is needed, leading to placement of the quadrupole focussing magnets as close as possible to the interaction point. Several accelerator physics requirements are considered in choosing the focussing configuration.

6.2.2 Parasitic beam crossings

The minimum bunch-to-bunch spacing is a critical parameter in determining the attainable luminosity. As this spacing decreases, the first parasitic crossing moves closer to the interaction point. The effects of one beam on the other at the parasitic crossings must be minimized. It is known that separation distance, bunch current, β function, and beam size all influence the beam-beam interaction at the parasitic crossing points. In this design study, the long range beam-beam tune shift (calculated for the first parasitic crossing with bunches spaced by 14 ns) was used as the criterion for choice of optics. With 14 ns spacing, the first parasitic crossing point occurs 2.1 m from the interaction point.

6.2.3 Optics

The focal length of the quads must be small enough to achieve the desired optics and must decrease roughly proportionally to the distance from the interaction point to the quads. The effect of chromatic errors of the quads is minimized by keeping the quads close to the interaction point. Operating energy range must be considered in options using permanent magnets. The required energy range was assumed to be 4.7 to 5.6 GeV.

6.2.4 Aperture

The aperture required to avoid particles hitting the vacuum chamber in the interaction region increases as the quads are moved farther from the interaction point. Previously the aperture requirements were most difficult to meet in the vertical plane. The displacement of the horizontal closed orbit resulting from the crossing angle makes the horizontal aperture a critical consideration. The orbits must minimize the horizontal as well as the vertical aperture requirements of the beam.

6.2.5 Mechanical requirements

The inner radius of the quads is determined by aperture requirements and the vacuum chamber thickness. The outer radius is determined by detector clearance requirements, both in operation and during installation. The mechanical stability and reproducibility (e.g.,after opening of the poles) must be adequate. A 1 micron transverse motion of an IR quad typically produces an 80 micron peak-to-peak distortion in the closed orbit of the beam.

There are several other magnet functions required in this region. Adjustable dipole steering magnets are needed to correct the closed orbits and also aid in creating localized orbit distortions, or “bumps,” to help reduce experiment background and injection losses. The effect of the CLEO solenoid field on the motion of the electrons must be adequately compensated. This will be done as now by rotating normal IR quadrupoles and by additional skew quads.

6.2.6 Magnet layout options

Several options for layout and magnet technology were considered. Only the (vertically focussing) quad nearest the interaction point was studied, since the next (horizontally focussing) quad was located outside of the CLEO magnet pole tip and the present iron/copper quads are adequate for most configurations. For each option, optics were computed at 5.3 GeV which provided β_{V}^{∗} = 1.0 cm and β_{H}^{∗} = 1 m. Apertures were designed to accommodate a 12 σ beam stay clear and a ±2.5 mrad horizontal crossing angle. A double-walled, cooled beam pipe was assumed.

For each configuration three parameters were calculated: 1) the long range tune shift for 10 mA bunches and fixed crossing angle at a point 2.1 m from the interaction point (corresponding to bunches spaced by 14 ns); 2) the chromaticity from the IR quads (on one side only); and 3) the total horizontal aperture required for ±2.5 mrad crossing angle. The results are summarized in Table 29.

NdFeB and SmCo refer to high B_{rem} permanent magnet material. The NdFeB material has B_{rem}≃1.2 Tesla, giving a significant advantage in gradient over the SmCo, which was assumed to have B_{rem}≃1.0 Tesla. (The present CESR permanent magnet quads are made from an older type of SmCo which has B_{rem}≃0.9 Tesla.) The NdFeB material, however has a lower Curie temperature (∼310 C compared to ∼480 C for SmCo) which causes the temperature coefficient to be 4 times larger, and requires special care to avoid exposing the material to excessive temperatures.

The first four rows in Table 29 are design exercises that do not easily accommodate the 4.7 to 5.6 GeV energy range. The SmCo options which can accommodate this energy range (rows 5 to 7) require unrealistically large gradients (k=0.8 m^{−1}) in the associated Fe/Cu quads. The superconducting quad options (rows 8 and 9) require placing the cryostat far

Table 29: Summary of IR quad options.

Beam-beam Chromaticity Horizontal
Quadrupole  Tune Shift  Apperature
Configurations  (/1000)  (for 1/2 IR)  (mm)
dQ_{H} dQ_{V} Q′_{H}  Q′_{V}  H ap
1. NdFeB start at z = 25 cm  3.71 2.45 -3.0 -5.9
2. NdFeB start at z = 30 cm  3.51 2.78 -2.7 -6.0  73
3. SmCo start at z = 30 cm  3.30 3.94 -2.4 -6.7  73
4. Above with 2 section REQ  3.26 5.09 -2.4 -7.0  73
5. SmCo 4.7–5.6 config: 4.7 GeV  3.88 2.95 -2.7 -6.4  69
6. SmCo 4.7–5.6 config: 5.3 GeV  3.30 4.71 -4.0 -4.8  68
7. SmCo 4.7–5.6 config: 5.6 GeV  3.07 5.61 -4.3 -3.9  70
8. SCQ 70 cm long, start at 60 cm 2.69 4.46 -1.7 -8.6  59
9. SCQ 54 cm long, start at 60 cm 2.82 3.66 -1.8 -7.8  58
10. SmCo 30–60 cm, SCQ 75–145 cm 3.18 4.09 -2.3 -7.4
11. SmCo 30–60 cm, SCQ 75–125 cm 3.27 3.35 -2.4 -7.0  66
12. SmCo 30–60 cm, SCQ 70–122 cm 3.32 2.93 -2.5 -6.7
13. 1992 HEP layout, 1.8 cm β^{∗}  2.46 15.6 -2.7 -6.8  51/86

Figure 55: Proposed CESR Phase 3 interaction region quadrupole layout.

inside the drift chamber which places undesirable constraints on the cryostat and the drift chamber design.

Hybrid options (rows 10 to 12) using SmCo and superconducting quads satisfy all of the requirements discussed earlier. The SmCo is most effective when high fields are required in small volumes. It also is a very effective mask against high energy particle background. The superconducting quads provide a controllable high-gradient field and offer the capability to include additional magnet coils in the same cryostat. We plan to install in the cryostat horizontal and vertical dipole correction coils and skew quad coils for detector solenoid compensation. The option in row 11 of Table 29 was chosen because it provides the best set of machine parameters while providing sufficient space for mounting detector and accelerator hardware. A drawing of the proposed quads is given in Fig. 55.

Current densities for the quadrupole windings are modest—less than 300 A/mm^{2}. Peak magnetic fields at conductors are approximately 3.5 Tesla, not including the solenoid field.

6.2.7 Mounting and support

We have chosen to support the superconducting and permanent quads from a double pedestal fastened to the CLEO magnet return yoke. This removes any support structures from the space between the drift chamber endplate and the endcap particle ID and calorimeter. It allows the quads to maintain accurate alignment even after opening the pole for access to the endplate. A detailed analysis of the mechanical properties remains to be done; however, preliminary results show the static properties to be satisfactory.

Since the permanent magnet quadrupoles have a smaller inner radius than the beam pipe at the end of the drift chamber, they must be assembled on the pipe with the vertex detector before installation in the drift chamber. They have been designed to fit well within the drift chamber, allowing room for a temporary support structure during installation.

6.2.8 Vacuum system

The vacuum system must achieve sufficiently low pressures for acceptable beam lifetime and low CLEO background rates. It must be compatible with masks to control synchrotron radiation and shield against particle background. The inner surface should have sufficient conductivity and smoothness to minimize resistive and higher order mode RF losses. Adequate cooling must be provided to remove the heat deposited by electro-magnetic and synchrotron radiation processes. The chamber radius must be large enough to accommodate the full beam-stay-clear requirements. The central chamber should present a minimum of scattering material to the decay products from the interaction point.

The vacuum system is similar in concept to that planned for Phase 2 upgrade operation. The double-walled beryllium central pipe will handle the heat load expected in Phase 3 operation. The pipe in the permanent magnet and superconducting quads will be double walled with coolant circulating between the inner and outer skins. Pumping would be provided by a titanium sublimation or NEG pump located just beyond the superconducting quad.

Beam diagnostic instrumentation will be installed as part of the vacuum system (e.g.,beam position monitors, luminosity monitors, beam loss detectors). These are not considered in detail in this study.

7 Overall summary

In the previous sections we have outlined in detail both the design and physics goals of the proposed CLEO III detector. A separate document summarizes this proposal and provides the detailed budgets, schedules, and construction time lines. Although this document has contributions from a large number of physicists in the CLEO collaboration, the formulation, direction, and overall coordination was carried out by the Inner Detector Task Force of C. Bebek, H. Kagan (Chairperson), R. Kass, H. Nelson, E. Nordberg, M. Ogg, D. Peterson, D. Rice, S. Stone, and M. Witherell.

References

[1] R. Ammar et al., Phys. Rev. Lett. 71, 674 (1993).

[2] M. Battle et al., “Observation of B^{0} Decay to Two Charmless Mesons,” CLNS 93/1235 (1993).

[3] M. Gronau and D. Wyler, Phys. Lett. 265B, 172 (1991).

[4] L. Wolfenstein, Phys. Rev. D31, 2381 (1985)

[5] B. Stech, Mod. Phys. Lett. A6, 3113 (1991); M. Neubert and B. Stech, Phys. Rev. D44, 775 (1991); K. Maltman et al., Nucl. Phys. A508, 371c (1990); W. A. Bardeen, A. J. Buras and J. M. Gerard, Phys. Lett. 192B, 138 (1987); M. B. Gavela et al., Phys. Lett. 154B, 425 (1985); and references contained therein.

[6] N. Isgur and M. Wise, Phys. Lett. 237B, 527 (1990).

[7] N. Isgur, D. Scora, B. Grinstein and M. Wise, Phys. Rev. D39, 799 (1989).

[8] J. G. Korner and G. A. Stech, Zeit. Phys. C38, 511 (1988) and Zeit. Phys. C41, 690 (1988).

[9] M. Bauer, B. Stech and M. Wirbel, Zeit. Phys. C29, 637 (1985).

[10] M. Neubert, “Model-Independent Extraction of V_{cb} from Semi-Leptonic Decays,” Heidelberg preprint HD-THEP-91-13 (1991).

[11] J. Bartelt et al., CLNS-93-1240, Sep. 1993.

[12] G. Altarelli et al., Nucl. Phys. B208, 365 (1982).

[13] R. Fulton et al., Phys. Rev. Lett. 64, 16 (1990).

[14] H. Albrecht et al., Phys. Lett. 234B, 409 (1990).

[15] M. Pauli(ARGUS), “Proceedings of the Joint International Symposium and Europhysics Conference on High Energy Physics,” ed. by S. Hegarty, K. Porter and E. Quercigh (World Scientific, Singapore, 1992), p. 592.

[16] J. D. Richman(CLEO), “Search for Exclusive bℓ →u Semileptonic Decays of B Mesons,” Proc. of the 7th Meeting of the APS Division of Particles and Fields, FNAL, Nov. 10-14, 1992.

[17] For a review see N. G. Desphande, “Theory of Penguins in B Decays, in B Decays, ed. by S. Stone, World Scientific, Singapore, (1992), p430.

[18] J. Bartelt et al., Phys. Rev. Lett. 71, 4111 (1993).

[19] R. Grigjanis et al., Phys. Lett. 237B, 252 (1990); B. Grinstein, R. Springer and M. Wise, Phys. Lett. 202B, 138 (1988).

[20] A. Ali and C. Greub, Phys. Lett. 259B, 182 (1991).

[21] A. Ali, T. Ohl and T. Mannel “Radiative Rare B Decays into Higher K-Resonances,” DESY 92-113 (1992); P. J. O’Donnell et al., “Exclusive Rare Decay B→K^{∗}γ,” UTPT-92-12 (1992); A. Ali and T. Mannel, Phys. Lett. 264B, 447 (1991), erratum Phys. Lett. 274B, 526 (1992); P. Colangelo et al., Zeit. Phys. C45, 575 (1990); C. A. Dominguez, N. Paver and Riazuddin, Phys. Lett. 214B, 459 (1988).

[22] N.G. Deshpande and P. O’Donnell, “Proceedings of the Workshop towards Establishing a b Factory,” Syracuse, Sept., 1989, p. 1.35 and 1.66.

[23] D. Du and Z. Xing, Phys. Rev. D48, 4155 (1993).

[24] S. Weinberg, Phys. Rev. Lett. 37, 657 (1976); G. Bélanger, C.Q. Geng, and P. Turcotte, Phys. Rev. D48, 5419 (1993); T. Hayashi, M. Matsuda, and M. Tanimoto, paper 305 submitted to the International Symposium on Lepton and Photon Interactions at High Energies, Cornell, 1993.

[25] E. H. Thorndike, Bull. Am. Phys. Soc. 38, 922(1993); D. Besson, invited talk at “1993 International Symposium on Lepton and Photon Interactions at High Energies,” Cornell.

[26] F. Borzumati and A. Masiero, Phys. Lett. 192B, 437 (1987) and erratum Phys. Lett. 198B, 590 (1987); R. Barbieri and G.F. Giudice, quoted by A. Sanda, International Symposium on Lepton and Photon Interactions at High Energies, Cornell, 1993.

[27] M. S. Alam et al., Phys. Rev. Lett. 51, 1143 (1983); M. S. Alam et al., Phys. Rev. Lett. 59, 22 (1987); H. Albrecht et al., Zeit. Phys. C42, 519 (1989); G. Crawford et al., Phys. Rev. D45, 752 (1992).

[28] C. Dominguez et al., Phys. Lett. 197B, 423 (1987); L. Reinders, Phys. Rev. D38, 947 (1988); estimates by Shifman and by Narison quoted by R. Schindler (SLAC-PUB-4995).

[29] C. Bernard et al., Phys. Rev. D38, 3540 (1988).

[30] E. Eichten et al., Phys. Rev. D34, 1547 (1986).

[31] I. Hinchliffe, preprint LBL-18742 (1984).

[32] R. Ammar et al., “Search for B^{0} decays to two charged leptons,” CLNS 93/1258.

[33] Report of the Working Group on CP violation and rare decays, Proceedings of the 1984 Snowmass Summer Study on the Design and Utilization of the Superconducting Super Collider.

[34] S. Stone, “Possible Ways of Observing CP Violation at a Symmetric e^{+}e^{−} B Factory,” Syracuse University preprint HEPSY 92-2 (1992); S. Stone, Nucl. Inst. Meth. A333, 15 (1993).

[35] We ignore the small contribution from color suppressed decays for diagram (a).

[36] For example, this would be true if only one amplitude dominates each of the decays to D^{0} or D̄^{0}.

[37] The phase δ in this discussion is the difference δ − δ̄ between the hadronic phases δ and δ̄ of the two amplitudes A and Ā, respectively, that are defined by Gronau and Wyler.

[38] We use λ ≈ 0.22 and a for the Wolfenstein parameter A to avoid confusion with our amplitude A.

[39] We use |V_{ub}/V_{cb}| = 0.08 ± 0.03 and λ is taken as 0.22.

[40] P. H. Cox, S. Hovater and S. T. Jones, Phys. Rev. D33, 1503 (1986); J. H. K´’uhn, S. Nussinov and R. R´’uckl, Zeit. Phys. C5, 117 (1980).

[41] M. S. Alan et al., CLNS 94-1270, submitted to Phys. Rev. D.

[42] K. Hikasa et al., Phys. Rev D45, Part II(1992).

[43] This number is in agreement with that observed in the CLEO II data. It is possible to increase the efficiency by adding more D^{0} decay modes and improving the detector solid angle for photons and the charged particle tracking efficiency. In all estimates we use 86% for the efficiency of reconstructing a high momentum track.

[44] G. Valencia, Phys. Rev. D39, 3339 (1992).

[45] A. Deandrea et al., Phys. Lett. 318B, 549 (1993); M. Bauer, B. Stech and M. Wirbel, Zeit. Phys. C34, 103 (1987).

[46] G. Kramer and R. Palmer, Phys. Rev. D45, 193 (1992); Phys. Lett. 279B, 181 (1992).

[47] G. Burdman and J. F. Donoghue Phys. Rev. D45, 187 (1992).

[48] A. Bean(CLEO), “Measurement of B(D^{+}→π^{0}ℓ^{+}ν)/B(D^{+}→K̄^{0}ℓ^{+}ν),” in proc. of the 7th Meeting of the APS Division of Particles and Fields, FNAL Nov. 10-14, 1992; private communication from Arne Freyberger.

[49] D. Bortoletto et al., Phys. Rev. Lett. 69, 2046 (1992).

[50] J. F. Amundson et al., Phys. Rev. D47, 3059 (1003).

[51] D. Acosta et al., CLNS 93-14 (1993).

[52] F. Muheim and S. Stone, “Three Model Dependent Estimates of B(D_{s}^{+} → ϕπ^{+}),” HEPSY 93-3 (1993), to be published in Phys. Rev. D.

[53] I. Bigi, Proceedings of the Charm Physics Workshop, Peking(1987); I. Bigi and A. I. Sanda, Phys. Lett. 171B, 320 (1986); J. Donoghue, E. Golowich, B. R. Holstein and J. Trampetic, Phys. Rev. D33, 179 (1986); L. Wolfenstein, Phys. Lett. 164B, 170 (1985); P. Colangelo and G. Nardulli, Phys. Lett. 242B, 71 (1990); H. Georgi, Phys. Lett. 297B, 353 (1992); T. Ohi, G. Ricciardi and E. H. Simmons, Nucl. Phys. B403, 605 (1993).

[54] D. S. Akerib et al.(CLEO), “Observation of D^{0} → K^{+}π^{−},” presented at Photon Leption Conference, Ithaca NY(1993). Singapore (1992).

[55] J. Haissinski, Orsay Preprint No. LAL-90-64, 1990, to be published in Proceedings of the 8th EPS General Conference: Trends in Physics (Amsterdam, The Netherlands, 1990).

[56] S. L. Glashow, A. Salam and S. Weinberg, Rev. Mod. Phys. 52, 515 (1980).

[57] For a recent review, see for example, K. K. Gan and M. L. Perl, Int. J. Mod. Phys. A3, 531 (1988); B. C. Barish and R. Stroynowski, Phys. Rep. 157, 1 (1988).

[58] See for example, P. Langacker, B. Sathiapalan and G. Steigman, Nucl. Phys. B266, 669 (1986) and references therein.

[59] H. Albrecht et al., Phys. Lett. 202B, 149 (1988).

[60] Y. S. Tsai, Phys. Rev. D4, 2821 (1971).

[61] F. J. Gilman and D. H. Miller, Phys. Rev. D17, 1846 (1978).

[62] H. B. Thacker and J. J. Sakurai, Phys. Lett. 36B, 103 (1971).

[63] D. Akerib et al., Phys. Rev. Lett. 69, 3610 (1992), ibid. Phys. Rev. Lett. 71, 3395 (1993).

[64] T. Das, V. S. Mathur and S. Okubo, Phys. Rev. Lett. 18, 761 (1967).

[65] M. Artuso et al., Phys. Rev. Lett. 69, 3278 (1992).

[66] F. J. Gilman, Phys. Rev. D35, 3541 (1987).

[67] A. Pich, Phys. Lett. 196B, 561 (1987).

[68] C. Leroy and J. Pestieau, Phys. Lett. 72B, 398 (1978); S. N. Biswas et al., Phys. Lett. 80B, 393 (1979); N. Paver and D. Treleani, Lett. Nuovo Cimento 31, 364 (1981); V. P. Barannik, A. P. Korzh and M. P. Rekalo, Acta Phys. Pol. B13, 835 (1982).

[69] See for example, H. Harari, Fundamental Forces, edited by D. Frame and K. J. Peach (St. Andrews, 1984), p.357; R. D. Pecci, New and Exotic Phenomena, edited by O. Fackler and J. Tran Thanh Van (Les Arcs, Savoie, France, 1987), p.431.

[70] A. Bean et al., CLEO preprint, CLNS 92/1172, submitted to Phys. Rev. Lett.

[71] H. Albrecht et al., Phys. Lett. 246B, 278 (1990).

[72] W. Fetscher, H.-J. Gerber and K. F. Johnson, Phys. Lett. 173B, 102 (1986).

[73] W. Fetscher, in Proposal for an Electron Positron Collider for Heavy Flavour Particle Physics and Synchrotron Radiation, PSI PR-88-09 (1988).

[74] M. Artuso et al., “Measurement of the Cross-Section for γγ → pp̄,” CLNS93/1245, submitted to Phys. Rev. D.

[75] D. Acosta et al., “Two-Photon Production of Charged Pion and Kaon Pairs,” in preparation for Phys. Rev. D.

[76] V. Shelkov et al., ”Measurement of the Two-Photon Production of the χ_{c2},” CLNS93/1250, accepted by Phys. Rev. D.

[77] R. Sonnenblick et al., Nucl. Inst. Meth. A310, 189 (1991).

[78] G. Hall et al., Nucl. Inst. Meth. A326, 228 (1993).

[79] J. F. Bak et al., Nucl. Phys. B288, 681 (1987).

[80] Private communication, members of the L3 collaboration.

[81] Note that Radeka (Low-Noise Techniques in Detectors, Ann. Rev. Nucl. Part. Sci, 38 (1988) 217) discusses series noise in a different limit, namely where RC is large; in this limit, ENC^{2} ∼ 1/τ. In our case, however, RC is small, and the resulting ENC is independent of the integration time.

[82] W. Trischuk, private communication. See also comments in O. Toker et al., “Viking, a CMOS low noise monolithic 128 channel frontend for Si-strip detector readout,” CERN-PPE/93-141.

[83] J. P. Alexander et al., Studies of Double-Sided, Double Metal Silicon Strip Detectors, contribution to the International Symposium on the Development and Application of Semiconductor Tracking Detectors, to appear in Nucl. Instr. Meth..

[84] Current versions of the VIKING chip yield a noise slope as low as 12.5/pF, with a 2.0 μs shaping time. A new version, now being tested, appears to offer 5.4/pF (private communication, P. Weilhammer.) We are presently exploring the possibilities of using such a chip, suitably tailored to CLEO needs.

[85] Such kapton circuits have been fabricated for gas microstrip detectors by CLEO collaborators at Ohio State University.

[86] W. Snoeys et al., Nucl. Inst. Meth. A326, 144 (1993); See also, S. Parker, Nucl. Inst. Meth. A275, 494 (1989).

[87] cf. for instance J. C. Jernigan, SLAC-PUB-5925.

[88] D. G. Cassel et al., Nucl. Inst. Meth. A52, 325 (1986).

[89] W. O. Miller et al., Los Alamos Nat. Lab. report LA-12029.

[90] D. Hale, Univ. of California, Santa Barbara, private communication.

[91] W. Adam et al., “The Ring Imaging Cherenkov Detector of DELPHI,” presented at First Int. Conf. on RICH Detectors, Bari, Italy, (1993); K. Abe et al., “Performance of the CRID at SLD,” ibid.; U. M´’uller et al., “Particle Identification with the RICH Detector in Experiment WA89 at CERN,” ibid.; H.-W. Siebert et al., “The OMEGA-RICH,” ibid.; R. Baur et al., “The CERES RICH Detector System,” ibid.; A. Braem et al., “Fast Rich Detector with a CsI Photocathode at Atmospheric Pressure,” ibid.; M. Staric et al., “Measurement of Cherenkov Angles with a CsI Photocathode in a Multiwire Chamber,” ibid.

[92] M. French et al., Nucl. Inst. Meth. A324, 511 (1990).

[93] R. Arnold et al., Nucl. Inst. Meth. A314, 465 (1992).

[94] J. Seguinot et al., Nucl. Inst. Meth. A297, 133 (1990).

[95] For more details see M. Artuso and Y. Mukhin, “Comparison of Expected Performance of Fast Ring Imaging Cherenkov Counter Systems,” CBX 93-45 (1993); M. Artuso and Y. Mukhin, “A Fast Ring Imaging Detector for the CLEO Upgrade,” presented at First Int. Conf.on RICH Detectors, Bari, Italy, (1993), to appear in Proceedings.

[96] D. Anderson et al., Nucl. Inst. Meth. A323, 626 (1992); J. Seguinot and T. Ypsilantis, Ring Imaging Cherenkov Counters, CERN Preprint CERN–LAA/PI/91-004 (1991); D. F. Anderson, S. Kwan and V. Peskov, “CsI and Some New Photocathodes,” FERMILAB-Conf-93/128 (1993); R. Aleksan et al., “Measurements of CsI Photocathode quantum efficiency in Methane,” presented at First Int. Conf. on RICH Detectors, Bari, Italy, (1993).

[97] S. Ilie and G. Lenzen, “Perfluorocarbon Liquid, Specific Chemical Aspects For Use in the DELPHI RICH,” DELPHI 93-33 (1993).

[98] R. Brandelik et al., Phys. Lett. 83B, 261 (1979); A. Onuchin et al., Nucl. Inst. Meth. A315, 517 (1992).

[99] H. Burkhardt et al., Nucl. Inst. Meth. 184, 319 (1981); M. Bockhorst, “Measurement of the optical properties of the components of an aerogel Cerenkov counter with wavelength shifter light guide,” BONN-IR-92-35, Sep 1992.

[100] G. Eigen, D. G. Hitlin, M. Lakata, J. Y. Oyang, and W. J. Wisniewski, “Silica aerogel Cerenkov detectors for particle identification,” SLAC-BABAR-NOTE-100(1992); G. Eigen, D.G. Hitlin and J.Y. Oyang, “Studies of silica aerogel as a low mass particle identification system for the SLAC B factory detector,” SLAC-BABAR-NOTE-101(1993).

[101] M. Selen and K. Honscheid, Particle ID via Cherenkov Correlated Timing, CLEO Preprint CBX 92-116; see also paper of same name by K. Honscheid, M. Selen and M. Sivertz, Proceedings of the First Int. Conf. on RICH Detectors, Bari, Italy, 1993.

[102] The idea of using total internal reflection to guide the Cherenkov light and to measure the position and shape of the Cherenkov rings as projected from the end of the radiator was first discussed by Blair Ratcliff. For a review of his ideas, refer to his talk “Status of the B Factory Detector for PEP-II: A Progress Report” at the XXVI. International Conference on High Energy Physics, Dallas (1992) (SLAC-PUB 5946 (92)). The first usage of total-internal reflection in Cherenkov counters was suggested by Fitch and Motley (Phys. Rev. 101, 496 (1956)).

[103] This is a conservative assumption. The fastest photo-tube available from Hammamatsu has a single electron transit time jitter of 160 ps FWHM.

[104] Advanced Photonics Inc., Camarillo, California 93012.

[105] M. Grande and G. R. Moss, Nucl. Inst. Meth. 215, 539 (1983).

[106] T. Browder et al., “CESR-B R&D Report, June 1993,” Cornell University, Ithaca, N.Y. page 4-1.